### v1

# Bridging Earth and Space: A Survey on HAPS for Non-Terrestrial Networks

Abstract
[…]
HAPS are emerging as key enablers in the evolution of sixth-generation (6G) wireless networks, bridging terrestrial and non-terrestrial infrastructures. Operating in the stratosphere, high-altitude platform station (HAPS) can provide wide-area coverage, low-latency, energy-efficient broadband communications with flexible deployment options for diverse applications. This survey delivers a comprehensive overview of HAPS use cases, technologies, and integration strategies within the 6G ecosystem. The roles of HAPS in extending connectivity to underserved regions, supporting dynamic backhauling, enabling massive Internet of things (IoT), and delivering reliable low-latency communications for autonomous and immersive services are discussed. The paper reviews state-of-the-art architectures for terrestrial and non-terrestrial network integration, highlights recent field trials. Furthermore, key enabling technologies—such as channel modeling, artificial intelligence (AI)-driven resource allocation, interference control, mobility management, and energy-efficient communications—are examined. The paper also outlines open research challenges. By addressing existing gaps in the literature, this survey positions HAPS as a foundational component of globally integrated, resilient, and sustainable 6G networks.
Index Terms — HAPS, 6G networks, non-terrestrial network (NTN), terrestrial network (TN), ultra-reliable low-latency communication (URLLC), network integration, energy-efficient communications.
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
I Introduction
The integration of TNs and NTNs is widely regarded as a key enabler for achieving the ambitious goals of 6G networks — to “connect the unconnected” and “hyper-connect existing infrastructures.” While traditional TNs are foundational, they cannot solely meet the growing demands for coverage, capacity, and low latency required by emerging applications [1]. To overcome these limitations, NTN technologies have emerged as transformative solutions. NTNs leverage space-borne (satellites), airborne (HAPS), and other non-ground-based platforms (unmanned aerial vehicles) to extend connectivity beyond terrestrial infrastructure, enabling coverage in remote and underserved regions, as well as in mobile platforms, e.g., oceans, rural areas, or in-flight scenarios, respectively. A key distinction between conventional TNs and NTNs is the latter’s ability to decouple service coverage from the physical location of backbone infrastructure through backhauling via satellite or HAPS links, achieved using long-distance gateway connections or multi-hop inter-platform links.
HAPS operate in the stratosphere, complementing both TNs and other NTN nodes such as low Earth orbit (LEO) satellites and UAVs. They combine the advantages of these systems while mitigating their limitations. HAPS offer wide-area coverage with higher capacity and lower latency than LEO satellites, and lower energy consumption with longer endurance than UAVs. Their quasi-stationary positioning reduces dependence on ground infrastructure and ensures reliable line-of-sight (LoS) or near-LoS connectivity. This capability allows HAPS to enhance network performance in diverse environments, including urban, rural, remote, and disaster-affected areas [2]. Beyond bridging connectivity gaps, HAPS enable advanced services such as massive IoT, real-time autonomous systems, edge computing, and immersive technologies like augmented and virtual reality [3, 2]. These features establish HAPS as a key element in shaping next-generation global connectivity.
The main objective of this survey is to present a comprehensive overview of the state-of-the-art in HAPS networks, examining key applications, enabling technologies, and future research directions. We first outline the diverse roles of HAPS across several domains, including connectivity enhancement, network infrastructure and management, IoT and emerging services, advanced control and special applications, as well as energy and sustainability. The survey further examines TN-NTN integration, emphasizing the role of HAPS as a bridge between terrestrial and non-terrestrial systems. It outlines platform and terminal types within this framework, reviews TN-NTN architectures, and explores emerging frequency bands such as millimetre wave (mmWave) and centimeter wave (cmWave), underscoring their importance for enhanced connectivity. The survey then reviews key HAPS technical enablers, including channel modeling, user equipment (UE) association and resource management, mobility and multi-link connectivity, massive multiple-input multiple-output (MIMO) and machine learning (ML) applications, interference control, and energy efficiency. It also addresses deployment and operational challenges, highlighting areas needing further research to fully realize HAPS potential. By covering these topics, the survey offers an in-depth understanding of technological advancements, design considerations, and open challenges in HAPS-based fifth-generation (5G) and 6G solutions, serving as a valuable resource for researchers and engineers in TN-NTN integration.
I-A Review of Previous NTN-Based Surveys
Before diving into this survey, we first highlight how it differs from previous surveys on the NTN topic.
I-A1 General NTN Surveys
Several surveys analyze NTNs, but few provide an in-depth examination of HAPS. The study in [4] explores UAVs in future communication systems, covering use cases, architectures, and challenges, including sub-6 GHz and mmWave TN optimization and UAV-to-UAV performance. The work in [5] reviews aerial connectivity for airplanes, vertical take-off and landing (VTOL) vehicles, and UAVs, analyzing use cases, requirements, and providers such as satellites and HAPS, though its coverage of HAPS is limited and lacks TN-NTN integration. Study [1] examines TN-NTN integration in 5G and 6G, focusing on mmWave, IoT/ Mobile Edge Computing (MEC) integration, and ML, but emphasizes LEO satellites and UAVs over HAPS. Similarly, [6] and [7] discuss LEO-focused TN-NTN integration, spectrum, MIMO modeling, mobility management, third generation partnership project (3GPP) standardization, and NTN evolution toward 6G, but only briefly mention HAPS. Overall, these surveys provide limited insight into HAPS as a central component of NTN networks.
I-A2 AI-Focused Surveys
[…]
AI is essential for NTNs and TN-NTN integration, enabling intelligent resource management, optimization, and automation across network layers. Several surveys examine AI in NTNs, focusing on key technologies, ML-based algorithms, and performance improvements. Study [8] investigates ML for UAV applications, including image processing, ground sensor localization, real-time control, scheduling, and trajectory planning. Reference [9] reviews 6G aerospace networks, analyzing satellite configurations, architectures, protocols, and AI-driven resource management, while highlighting research gaps. Work [10] explores ML techniques for NTN challenges in 6G, addressing privacy risks and optimization via both, deep learning and deep reinforcement learning (DRL). Study [11] analyzes AI-based strategies for TN-NTN integration, convergence drivers, key applications, and architectures. Despite this coverage, these surveys primarily focus on satellites and UAVs, with limited discussion of HAPS.
I-A3 HAPS-Related Surveys
The survey in [3] provides a comprehensive tutorial on HAPS and low-altitude platform stations, highlighting their roles as aerial base stations and cellular-connected UE. It examines challenges such as three-dimensional (3D) deployment, channel modeling, and energy efficiency, and reviews analytical frameworks including optimization theory and ML for UAV-related issues, but does not address TN-NTN integration. Study [12] reviews LEO satellite and HAPS communications, covering system design, networking, medium access control, air interface, testbeds, and NTN integration, yet focuses primarily on satellites. Work [13] considers NTNs for 6G, simulating TN-LEO and TN-LEO-HAPS integration, but treats HAPS only as relays. The survey in [2] focuses on HAPS in mega-constellation NTNs, exploring super macro base stations, regulatory aspects, channel modeling, resource management, handoff, swarm control, AI applications, and emerging technologies such as reconfigurable intelligent surfaces and faster-than-Nyquist signaling, yet lacks insights on TN-NTN interoperability. Finally, [14] examines HAPS for coverage extension and capacity enhancement in TN-HAPS coexistence, discussing intelligent resource management, interference mitigation, spatial optimization, and past projects, but does not address advanced applications like IoT support, AI-driven optimization, or fully integrated TN-NTN architectures.
Among the most relevant surveys [3, 12, 13, 2, 14], none offer a comprehensive, up-to-date analysis of HAPS covering both technological advancements and TN-NTN integration. Studies [3, 12] provide general NTN overviews with limited HAPS discussion, focusing on use cases, architectures and capabilities. While [13] addresses HAPS in NTN-enabled 6G, it emphasizes relay-based integration rather the more powerful regenerative-based one based on independent network nodes in the sky. Survey [2] reviews HAPS mega-constellations, regulations, and enabling technologies, yet offers limited insight on NTN deployment and TN-NTN interoperability. Work [14] considers coverage extension and terrestrial coexistence, but does not cover broader TN-NTN integration, advanced use cases, or emerging AI-related technologies. To address these gaps, this survey provides a holistic, forward-looking review of HAPS, including TN-NTN integration, enabling technologies, real-world deployment challenges, and future research directions.
I-B Survey Structure
The rest of this survey is organized as follows Section II compares HAPS with other NTN nodes. Section III highlights the most promising use cases for HAPS. Section IV explores scenarios and architectures for HAPS deployment. Section V examines the key technical enablers for HAPS. Section VI discusses the current challenges related to HAPS communications. Section VII concludes the survey with findings and insights for future research directions.
List of Abbreviations
3D
three-dimensional
3G
third generation
3GPP
third generation partnership project
4G
fourth-generation
5G
fifth-generation
6G
sixth-generation
AI
artificial intelligence
AR
augmented reality
AV
aerial vehicle
BER
bit error rate
BS
base station
CNN
convolutional neural network
cmWave
centimeter wave
CoMP
coordinated multi-point
CPU
central processing unit
CSI
channel state information
D2D
Device-to-Device
dB
decibel
DDPG
deep deterministic policy gradient
DFT
discrete Fourier transform
DRL
deep reinforcement learning
EB
exabytes
eMBB
enhanced mobile broadband
FL
federated learning
FR
frequency range
FSO
free space optical
GEO
geostationary equatorial orbit
GUE
ground user equipment
HAPS
high-altitude platform station
HAPS-SMBS
high-altitude platform station mounted super macro base station
HARQ
hybrid automatic repeat request
IoRT
Internet of remote things
IoT
Internet of things
ISAC
integrated sensing and communication
ITU-R
International Telecommunication Union Radiocommunication Sector
LAPS
low-altitude platform station
LEO
low Earth orbit
LoS
line-of-sight
LoRa
long range
LTE
long term evolution
MBS
macrocell base station
MC
multi-connectivity
MDP
Markov decision process
MEC
Mobile Edge Computing
MEO
medium Earth orbit
MIMO
multiple-input multiple-output
ML
machine learning
mMIMO
massive multiple-input multiple-output
MMSE
minimum mean square error
mMTC
massive machine type communication
mmWave
millimetre wave
NB-IoT
narrow band IoT
NOMA
non-orthogonal multiple access
NR
new radio
NTN
non-terrestrial network
OFDM
orthogonal frequency division multiplexing
OFDMA
orthogonal frequency division multiple access
PDCP
packet data convergence protocol
PHY
physical
PPO
proximal policy optimization
QoS
quality of service
RA
resource allocation
RAN
radio access network
RAT
radio access technology
RB
resource block
REM
radio environment map
RF
radio frequency
RL
reinforcement learning
RSRP
reference signal received power
RIS
reconfigurable intelligent surface
SAGIN
space-air-ground integrated network
SatCom
satellite communication
SBS
small base station
SINR
signal to interference plus noise ratio
SLNR
signal to leakage interference and noise ratio
SMBS
super macro base station
SNR
signal to noise ratio
SVM
support vector machine
TBS
terrestrial base station
THz
terahertz
TN
terrestrial network
UAV
unmanned aerial vehicle
UE
user equipment
UMTS
universal mobile telecommunications system
URLLC
ultra-reliable low-latency communication
VEC
Vehicular Edge Computing
VHetNet
vertical heterogeneous network
VR
virtual reality
VSAT
very small aperture terminal
VTOL
vertical take-off and landing
II HAPS Comparison with Terrestrial Networks and Other NTN Nodes
Global connectivity, traditionally supported by TNs and satellites, still faces significant gaps. Despite mobile broadband traffic doubling from 419 exabytes (EB) in 2019 to 913 EB in 2022, about 33% of the world’s population remains offline [15]. This digital divide highlights the need for innovative solutions. HAPS, operating in the stratosphere at 20–50 km, offer a promising approach. Acting independently or as intermediaries between TNs and other NTN systems such as LEO satellites and UAVs, HAPS provide wide-area coverage, low latency, and operational flexibility. These features enable them to bridge connectivity gaps while complementing existing networks. This section examines HAPS capabilities compared with TNs (Section II-A), LEO satellites (Section II-B), and UAVs (Section II-C), highlighting their unique advantages. Table I summarizes the distinct characteristics of HAPS relative to TNs and other NTN nodes in terms of altitude, performance, operational flexibility, and cost.
Table I: Comparison of HAPS with TNs, LEO satellites, and UAVs.
Feature TNs LEO satellites UAVs HAPS
Altitude Ground-based 500–2,000 km Up to 1.5 km 20–50 km
Coverage area Local to regional Global (constellations) Localized Regional to wide-area
Latency Low Moderate to high Low Low
Operational duration Continuous Years Hours to days Weeks to months
Deployment flexibility Limited by infrastructure Rigid (requires launch) Highly flexible Moderate (fixed stratospheric path)
Cost High for remote areas Very high Moderate to low Moderate
Power source Grid or renewable energy Solar or batteries Batteries Solar-powered
Payload capacity High Moderate Low Moderate to high
Adaptability Limited Low High Moderate to high
Emergency response Slow (requires rebuilding) Limited Rapid deployment Rapid deployment
II-A Comparison with Terrestrial Networks
TNs serve as the backbone of modern communication systems, providing high-capacity, low-latency connectivity in densely populated areas. However, they are constrained by geographical and infrastructural limitations. Rugged terrains, remote regions, and disaster-stricken areas often lack adequate terrestrial infrastructure, leaving significant portions of the population unconnected. Furthermore, the deployment and maintenance of TNs in such regions involve high costs, making them impractical for universal connectivity. HAPS overcome these limitations by offering wide-area coverage without reliance on extensive ground infrastructure. They can provide reliable LoS connectivity across urban, rural, and remote areas, bridging gaps where terrestrial solutions are infeasible. Their rapid deployment capability also makes them ideal for emergency scenarios, while operational costs in remote areas are often lower than constructing terrestrial networks.
II-B Comparison with LEO Satellites
LEO satellites provide essential global coverage, particularly in remote and oceanic regions. However, their high operational altitudes (500 to 2,000 kilometers) result in increased signal attenuation and latency. Moreover, their large coverage areas imply that the bandwidth is shared among numerous UEs, leading to lower per-UE data rates. These constraints make LEO satellites less suited for high-bandwidth or latency-sensitive applications. In contrast, HAPS are positioned closer to Earth, which significantly reduces latency and enables a better spatial reuse, resulting in an enhanced bandwidth availability per UE and higher data rates. Additionally, HAPS are more operationally flexible than LEO satellites. While satellites require complex and costly launches, HAPS can be repositioned and upgraded with minimal disruption, making them adaptable to evolving connectivity demands.
II-C Comparison with UAVs
UAVs are highly versatile, providing localized and flexible connectivity for hot-spot tactical operations and disaster scenarios. However, their limited flight duration due to battery power and energy constraints, payload capacity, and frequent maintenance requirements hinder scalability in large or long-term deployments. In contrast, HAPS offer extended operational periods, lasting weeks or even months, powered by solar energy and autonomous systems. Their quasi-stationary positions enable them to serve as stable, high-capacity communication hubs, bridging the gap between the localized capabilities of UAVs and the global reach of satellites.
III HAPS Use Cases
The unique characteristics of HAPS enable a wide range of applications across telecommunications and beyond. In telecommunications, they primarily aim to extend coverage (Section III-A) and enhance network efficiency (Section III-B). Their rapid deployment and adaptability make them essential for bridging connectivity gaps in remote, underserved, or disaster-affected regions. Serving as mobile urban BSs or backhaul providers, HAPS improve capacity, reliability, and energy efficiency. Beyond communications, they support advanced applications (Section III-C), including real-time IoT sensing, immersive augmented reality (AR)/ virtual reality (VR), precision control via URLLC (Section III-D), and energy optimization (Section III-E). By bringing processing closer to UEs, HAPS reduce latency and enable innovations across sectors such as agriculture, logistics, and disaster management. This section explores their roles in connectivity enhancement and advanced use cases shaping next-generation networks.
III-A Connectivity Enhancement
III-A1 Remote and Rural Area Coverage
The COVID-19 pandemic highlighted global Internet access disparities, especially in remote and underserved regions, emphasizing the need for scalable connectivity solutions [16]. HAPS can address these gaps by extending mobile coverage to areas lacking terrestrial infrastructure. Operating similarly to TNs, they can bridge the digital divide and deliver reliable global connectivity [17].
Collaboration between HAPS and LEO satellites was explored in [18], which formulated a mixed-integer nonlinear programming problem to enhance large-scale access and backhaul. The proposed restricted trilateral matching algorithm improved coordination, connectivity, and efficiency in remote regions.
Study [19] analyzed HAPS-LAPS integration using stochastic geometry, introducing a framework to assess platform density and altitude effects under backhaul constraints, offering design guidance for robust aerial networks in rural areas.
In [14], HAPS were studied for extending rural coverage with low UE density through integration with TNs, emphasizing coexistence via dynamic resource management, interference mitigation, and techniques such as advanced antennas, radio environment maps, and device-to-device communication.
Case study [20] optimized resource allocation for heterogeneous services in rural HAPS networks using free space optical (FSO) backhaul. It proposed static and dynamic allocation schemes leveraging DRL, improving age of information, data rates, and overall network efficiency.
III-A2 Natural Disaster Region Coverage
In the aftermath of natural disasters, when TNs collapse, HAPS can provide real-time aerial surveillance and rapid, reliable communication. Covering vast and remote regions, they ensure effective disaster response and emergency connectivity for first responders [21]. Study [22] examined UAV-based HAPS for disaster scenarios, highlighting their rapid deployment, flexible coverage, and communication recovery, while identifying challenges such as backhaul limitations, platform constraints, and communication capacity that require targeted solutions. Real-time monitoring using HAPS was demonstrated in [23], showing persistent, high-resolution imaging from altitudes around 20 km, enabling stable, continuous coverage of affected areas. In [24], HAPS restored services in a disrupted universal mobile telecommunications system (UMTS) network, achieving 70% throughput recovery when terrestrial BSs were unavailable, underscoring their effectiveness in disaster recovery operations.
III-A3 Unexpected Traffic Events
The inherent mobility and on-demand deployment of HAPS make them well suited for managing dynamic communication demands, including large public events, sudden traffic surges, or routine daily traffic variations [2, 25]. By adapting to these fluctuations, HAPS alleviate congestion, enhance quality of service (QoS), and improve energy efficiency. Compared to small cell densification, which requires extensive planning, fixed infrastructure, and offers limited flexibility, HAPS provide a faster and more cost-effective solution that can be deployed in real time without physical constraints. A QoS-driven deployment framework was proposed in [25] to optimize HAPS operation based on coverage, capacity, and latency metrics, balancing efficiency and cost. In integrated networks where ground and aerial BSs share spectrum, interference remains a key issue. As noted in [26], advanced techniques such as dynamic resource allocation, spectrum sharing, and adaptive power control enable efficient HAPS operation and seamless coexistence with TNs, ensuring robust performance under variable traffic conditions.
III-B Network Infrastructure and Management
III-B1 Backhauling
Network backhaul is a fundamental component of communication infrastructure, responsible for high-capacity, low-latency data transmission between BSs and the core network. In dense urban areas, fiber optics remain the dominant solution, offering high throughput and minimal latency, whereas copper cables, limited in bandwidth, are largely obsolete for modern traffic demands. In contrast, rural and remote regions often rely on satellite or microwave links, as fiber deployment is economically and logistically prohibitive [27]. Non-geostationary satellite systems require costly ground gateways with precise placement to maintain reliable connectivity [28].
Although wireless backhauling has been studied extensively [29], HAPS are emerging as a promising enabler for next-generation networks. Positioned in the stratosphere, they provide near-LoS connectivity with lower latency compared to satellites, supporting applications such as small cell backhauling, wireless sensor networks, and Earth observation [30]. In-band mmWave backhaul has been demonstrated [31], while multi-band approaches for enhanced capacity are discussed in [32].
Recent studies propose advanced architectures leveraging HAPS. In [33], a cell-free scheme introduces HAPS as an aerial central processing unit (CPU) aggregating UAVs via terahertz (THz) links, where large antenna arrays mitigate path loss. Hybrid radio frequency (RF)/FSO solutions further enhance flexibility: [34] presents a cooperative hybrid ARQ protocol ensuring latency fairness under varying conditions, and [35] explores hybrid LEO-to-HAPS FSO and HAPS-to-ground RF links. Collectively, these advancements demonstrate the potential of HAPS-enabled backhaul to deliver scalable, low-latency connectivity across heterogeneous environments.
III-B2 Handover Management
Frequent handovers and associated failures represent a major challenge in satellite communication systems, particularly for fast-moving non-geostationary satellites that require continuous connectivity updates. HAPS mitigate these issues in two key roles: (i) as intermediaries between satellites and UEs, handling handover requests on behalf of satellites, and (ii) as standalone communication managers, directly managing UE connections. In both cases, HAPS improve network stability and reduce signaling overhead.
In [36], leveraging HAPS as intermediaries was shown to decouple satellite mobility from ground-based handover procedures, simplifying management and enhancing continuity. Dynamic LEO-HAPS handover strategies proposed in [37] achieved lower delays, signaling costs, and drop rates than traditional 5G methods. Similarly, [30] explored FSO-based LEO-HAPS links, demonstrating high-speed, low-latency handovers even under high satellite mobility.
When operating as quasi-stationary communication nodes, HAPS significantly reduce handover frequency for terrestrial UEs, providing seamless service in high-mobility scenarios (e.g., vehicles, aircraft). Their stable positioning eliminates inter-platform handovers common in LEO constellations. The airliner system in [38], combining beamforming for HAPS-to-ground and FSO backhaul to LEO/ medium Earth orbit (MEO) satellites, demonstrated efficient handover management even in dynamic environments. Moreover, optical HAPS-satellite links, as analyzed in [39], further enhance reliability by ensuring robust backhaul connectivity, thereby minimizing disruptions in HAPS-to-UE sessions.
III-C IoT and Emerging Applications
III-C1 IoT sensing
HAPS can significantly enhance IoT sensing and monitoring by providing wide, stable coverage and integrating edge computing for low-latency data processing close to the source. Their dynamic positioning supports optimized sensor deployment and efficient data acquisition, making them ideal for remote and critical environments.
A reliable IoT architecture leveraging HAPS was proposed in [40], integrating narrow band IoT (NB-IoT) and long range (LoRa) to deliver scalable, cost-efficient connectivity across underserved regions. Operating at 17–35 km, HAPS achieved coverage up to 600 km, surpassing terrestrial solutions in scalability and redundancy through overlapping cells.
In [41], a joint HAPS-UAV framework for IoT applications—especially disaster scenarios—was introduced. Combining umbrella coverage from HAPS and agile UAVs, the system ensured continuous service with sub-millisecond latency (0.07–0.4 ms for HAPS), while NB-IoT/LoRa access schemes minimized collisions under dense deployments.
Complementary LoRa-based monitoring was examined in [42], which optimized parameters such as altitude, transmission power, and device density under shadowing and collision constraints, confirming the viability of HAPS for large-scale environmental IoT.
To enhance HAPS network capacity, [43] introduced a 5G IoT architecture with a uniform rectangular array antenna and a Long Short-Term Memory (LSTM) model for direction-of-arrival prediction. The approach improved beamforming precision and link stability, indirectly benefiting IoT reliability. Joint LEO-HAPS IoT integration was explored in [44], where HAPS relayed data to ground centers via LEO satellites, optimizing throughput and power consumption under QoS and storage constraints. Similarly, [45] presented a smart rural framework combining 5G, cloud, IoT, and HAPS, enabling automated remote data access and processing.
III-C2 Mobile Edge Computing and Fast UE Movement
MEC enhances network performance by placing computing resources near UEs, reducing latency for real-time services. However, maintaining seamless connectivity for fast-moving UEs requires advanced handover and predictive mechanisms to sustain service continuity. In [46], the integration of HAPS and UAVs within Vehicular Edge Computing (VEC) systems was analyzed using queuing theory to optimize task offloading in dynamic vehicular environments. The study showed that HAPS, with quasi-stationary operation, broad coverage, and strong LoS links, outperform UAVs by supporting larger computing payloads and offering lower latency, making them ideal for large-scale, high-mobility VEC/MEC scenarios. Although prior works such as [47, 48] optimized UAV-enabled MEC for task completion and energy efficiency, these approaches can be extended to HAPS, which offer superior payload capacity and operational stability. Yet, research on HAPS-based MEC remains limited, representing a promising avenue for future exploration.
III-C3 Aerial Data Centers and AR/VR Support
Beyond edge computing, HAPS can function as aerial data centers, delivering centralized cloud services for data storage, processing, and distribution from the stratosphere. With wide coverage, high capacity, and low-latency connectivity, they support resource-intensive applications such as AR and VR [49], enabling immersive experiences even in remote or underserved regions [50]. In [51], a HAPS-enabled VEC system for rural areas was analyzed, showing performance gains through optimized trade-offs between computational load and latency. A cloud-integrated HAPS system proposed in [52] achieved up to 14% energy savings compared to terrestrial data centers, reducing offloading outages and delays in distributed server networks while supporting services such as IoT, vehicular networks, and social media. Furthermore, [44] introduced a collaborative LEO-HAPS framework in which HAPS acted as relays, forwarding IoT data to ground processing centers with optimized data rates, power efficiency, and QoS compliance.
III-D Advanced Control and Special Applications
III-D1 URLLC Support and Remote Driving Control
Operating at high altitudes but closer to the Earth than traditional satellites, HAPS offer a unique combination of wide-area coverage, low-latency communication, and substantial payload capacity. These features enable the integration of advanced antenna systems and onboard processing units, making HAPS well-suited for supporting URLLC services such as remote driving control. In [53], a heterogeneous network integrating HAPS, terrestrial BSs, and autonomous vehicles was analyzed, showing that multi-connectivity (MC) is essential to meet stringent URLLC requirements. HAPS acted as auxiliary links sustaining data rates above 700 kbps with high reliability under dynamic conditions. Building on this, [54] extended the model with LEO satellites, optimizing MC path selection to satisfy QoS constraints with minimal link usage. Results revealed that LoS interference is a key performance-limiting factor in radio access links, which was further addressed in [55] through refined MC strategies for enhanced reliability and efficiency. Beyond URLLC, HAPS can improve Global Navigation Satellite System (GNSS) performance due to their quasi-stationary position and strong signal power. In [56], a HAPS-aided GPS architecture demonstrated improved 3D positioning accuracy, enhancing navigation and control functions. This capability is particularly valuable for remote or underserved regions—such as mountainous, coastal, or oceanic areas—where HAPS can act as super macro base stations (HAPS-SMBS) to backhaul isolated nodes and extend coverage [2].
III-D2 UAV Control and Aerial Highways
The cellular community has made significant efforts to enable reliable UAV connectivity and control by leveraging existing terrestrial networks while minimizing interference with ground UEs [57, 58]. Current approaches, such as time- and frequency-domain resource separation, offer partial improvements but lack scalability to accommodate the projected growth of aerial traffic. To overcome these limitations, research has explored enhanced infrastructure densification, dedicated aerial communication systems [59, 60, 61, 62, 63], and LEO-based augmentation [6].
A promising solution involves establishing aerial highways—structured UAV corridors that enable regulated, safe, and efficient airspace management [64, 65]. To support seamless communication along these corridors, recent works have optimized cellular network configurations and resource allocation strategies tailored for aerial traffic [66, 67, 68, 69, 70], while theoretical frameworks have been proposed to design structured airspace operations [71, 72].
With their quasi-stationary position, large coverage footprint, and low-latency links, HAPS are well-suited to serve as backbone infrastructure for both open-air and corridor-based UAV networks. As discussed in Section III-B, HAPS significantly reduce handover frequency and ensure continuous connectivity for high-mobility UAVs, while avoiding the inter-platform coordination challenges inherent to LEO systems. This combination of coverage, stability, and operational efficiency positions HAPS as a key enabler for scalable and reliable UAV control.
III-E Energy and Sustainability
Energy efficiency is a critical consideration for future integrated TNs and NTNs. Traditional TNs rely on continuous operation of macro BSs to provide ubiquitous coverage, leading to substantial energy consumption even during low-traffic periods. This ”zero bit, non-zero watt” phenomenon reflects the inefficiency of maintaining coverage when demand is minimal.To address this issue, a promising solution involves offloading traffic from terrestrial macro BSs to NTN-powered hypercells, such as those supported by LEO satellites, HAPS, or tethered UAVs. During nighttime or periods of low demand, this approach allows power-intensive terrestrial macro BSs to enter low-power states or shut down entirely, significantly reducing energy consumption. In this context, LEO satellites are well-suited for offloading outdoor UEs due to their extensive coverage and ability to serve low-capacity traffic efficiently during nighttime. In contrast, HAPS, with their superior link budget and better signal penetration, are better suited for offloading indoor UEs, especially in urban environments where high spatial reuse is essential.
In [73], a heuristic algorithm was developed to optimize this offloading process. By prioritizing the least-utilized BS for offloading, and leveraging HAPS’s self-sufficient energy capabilities, the proposed approach minimized hourly energy consumption across the network. Simulation results indicated that this method could reduce energy consumption by up to 29% over a week, with savings reaching 41% during nighttime hours. These results highlight the complementary roles of HAPS in advancing sustainable and energy-efficient connectivity within integrated TNs and NTNs.
IV HAPS Architectures
The integration of TNs and NTNs is critical to addressing the diverse connectivity challenges identified in Section III, such as extending coverage to remote areas, enabling robust disaster communication, managing unexpected traffic surges, supporting IoT applications, and facilitating advanced services such as mobile edge computing and AR/VR. By serving as dynamic platforms and connectors, HAPS can play a key role in bridging these systems and creating a resilient, high-capacity, global communication infrastructure.
To this end, NTN platforms must efficiently transmit/receive communication signals between UEs, gateways, BS, and core networks. The NTN architecture, as defined by the 3GPP in [74], supports two fundamental payload designs: transparent (or bent-pipe) and regenerative. In the transparent approach, the NTN platform carries communication equipment that functions as a simple relay, forwarding signals without modification. In contrast, the regenerative architecture enables onboard signal processing at the NTN platform, allowing the digital processing of the —from packet data convergence protocol (PDCP) to physical (PHY) layer— to be performed directly within the platform. These architectural choices impact key factors such as spectral efficiency, rate, latency, and overall system complexity, shaping how NTNs integrate with TNs. Building on this foundation, we can distinguish three main network components within the NTN architecture:
  • Platforms, which carry communication equipment and implement either a transparent or regenerative payload to relay or generate communication signals.
  • Terminals, including direct UEs or specialized devices designed to interact with NTN platforms.
  • Gateways, which connect NTNs, e.g., satellites and HAPS, to core networks, facilitating data flow between TNs and NTNs.
At the same time, the NTN architecture incorporates the following types of links [75]:
  • Service links, which facilitate direct communication between NTN platforms and UEs, ensuring end-user connectivity across diverse environments.
  • Feeder links, which connect NTN platforms to gateways, operating at high frequencies to support high-capacity data transmission with error correction.
  • Backhaul links, which connect NTN gateways to core network facilities, typically relying on fiber-optic infrastructure to ensure high-speed data transport.
  • Inter-platform links, which enable communication between NTN platforms, supporting seamless handovers and reducing dependence on terrestrial infrastructure. They usually use FSO infrastructure.
In Section IV-A, we will describe the different NTN components, including platforms, terminals, gateways, and their interactions, focusing on HAPS. In Section IV-B, we will discuss promising frequency bands for NTNs based on the link types, with considerations related to range, throughput, latency, and interference sensitivity. In Section IV-C, we will compare 3GPP bent-pipe and regenerative architectures, focusing on the role of HAPS. Finally, in Section IV-D, we will present relevant examples of field trials and deployments involving HAPS.
IV-A NTN Components
Table II: Comparison of HAPS platforms.
Characteristic Aerostatic Aerodynamic Hybrid
Lift mechanism Buoyant force (helium/hot air) Aerodynamic lift (fixed/rotary wings) Combination of buoyant and aerodynamic forces
Mobility Low when operating stationary or drifting slowly; high with favorable wind patterns High, regardless of wind direction; can reposition dynamically Moderate (repositioning capability)
Payload capacity High Low Moderate (payload/mobility trade-off)
Energy efficiency High (minimal energy to stay aloft) Low (continuous energy for flight) Moderate (depends on lift mechanics)
Weather resilience Sensitive to wind and turbulence More resistant to adverse weather, better operational stability Balanced
Deployment flexibility Requires ground infrastructure (hangars, airfields) in close proximity to the serving region Can take off and land outside serving region (deployment flexibility) Moderate (limited by mobility)
Operational duration Long and persistent (weeks/months) Variable (days/months), and may be solar-powered to extend flight life Potentially long, depending on design and hybrid lift use
Development cost Low, due to simpler design and manufacturing High, requiring advanced materials and engineering Highest, due to the complexity of integrating multiple lift systems
Notable example Loon Balloons [76] SoftBank’s “Sunglider” [77] Tactical HAPS by CIRA [78]
Figure 1: General NTN representation.
In general, NTNs (see Fig. 1) include several key components, each playing a specific role in ensuring global, reliable communication. These components can be categorized into three main layers, namely space, aerial, and ground layers [79]. The space layer comprises geostationary equatorial orbit (GEO), MEO, and LEO satellites, while the aerial layer consists of HAPS and low-altitude UAVs. Satellites, HAPS, and UAVs in this case are considered as network platforms, which carry the node able to facilitate communications. However, in certain use cases, they may serve as network terminals, i.e., connectivity consumers, providing no communication functionality. The ground layer typically includes infrastructure that interacts with the space and aerial layers, such as gateway stations, core networks, and data centers. Considering TN-NTN integration, the different layers can be combined together to provide better coverage, capacity, resilience, and flexibility compared to standalone deployments [80]. Now, we shift our focus to the aerial layer, with a particular emphasis on HAPS, the subject of this survey. We distinguish between NTN platforms, terminals, and gateways.
IV-A1 Platforms
HAPS can be classified depending on the physical principle providing the lifting force:
  • Aerostatic platforms, which rely on buoyant force to stay aloft, and include airships and balloons.
  • Aerodynamic platforms, which generate lift through fixed or rotary wings to maintain flight.
  • Hybrid platforms, which rely on a combination of aerostatic and aerodynamic effects at different stages of flight, such as the tactical HAPS by CIRA [78].
All these types of HAPS have their own benefits and drawbacks, which we discuss in the following and are summarized in Table II.
Aerostatic platforms
Aerostatic platforms are less expensive to manufacture, launch, and operate compared to aerodynamic and hybrid systems. While prototyping and refining their products, the Loon project team was able to fill and launch a balloon into the stratosphere every 30 minutes with their designed and custom-built auto-launchers and keep it in the air for more than 300 days [76]. Thus, by hovering over a specific region, aerostatic platforms can provide consistent coverage and are well-suited for persistent surveillance and localized communication needs. Since this platforms do not require forward speed to maintain flight, they only need power to hold or adjust their position. This results in potentially better power efficiency, and allows for heavier payloads compared to aerodynamic platforms. Furthermore, their large volume capacity enables them to carry more complex communication payloads, making aerostatic platforms particularly suitable for both short-term and long-term missions over specific regions. However, these platforms are more sensitive to atmospheric conditions, which may impact their operational stability and reliability. Additionally, launching and landing aerostatic platforms require specific facilities such as hangars and airfields located near the target region, which may limit their flexibility in deployment. However, the Loon project team demonstrated a method to navigate the balloons using dominant winds at different altitudes. This approach requires advanced predictive models to identify favorable wind patterns and intelligent algorithms to determine the most efficient flight paths through varying wind layers [76]. This allows the launch facilities to be located away from the target region, but makes deployment reliant on wind patterns.
Table III: Comparison of NTN terminals.
Characteristic Portable consumer terminals Vehicular terminals Fixed UE terminals
Mobility Medium (handheld/wearable devices) High (cars, trains, aircraft, UAVs) Low (stationary gateways/sensors)
Power availability Limited (battery-powered) Moderate to high (power sources) High if fixed power supply available; limited for low-power IoT sensors
Antenna Small, integrated, limited gain Large, more powerful (e.g., VSAT, phased arrays) High-gain dish or directional antennas for long-range communication; limited gain for low-power IoT sensors
Coverage
Handover requirements
Seamless transition between TNs and NTNs for uninterrupted connectivity Multi-band, multi-standard support for high-speed mobility across regions Persistent connectivity with high-capacity backhaul support; limited data rates for low-power IoT sensors
Use cases Wearables, first responder devices, personal communication (phones) Connected vehicles, public transport, autonomous robots, UAVs Broadband Internet, industrial IoT, remote sensors
Network requirements Low latency for applications like gaming and video streaming Low latency for safety-critical applications, high throughput for real-time data High throughput for broadband applications, low power for IoT sensors
Deployment constraints Consumer-driven must be cost-effective and compact Requires vehicle integration and external antennas Fixed installation with higher infrastructure requirements
Example devices Smartphones, tablets, smartwatches, disaster relief radios Connected cars, aircraft communication systems, surveillance drones Satellite broadband dishes, industrial IoT gateways, environmental sensors
Aerodynamic platforms
Aerodynamic platforms offer significant advantages given their high flexibility, allowing them to move easily between regions in response to communication demands or environmental changes. This adaptability reduces the need for multiple airfields and enhances their capability for precise positioning through specific flight patterns. Modern aerodynamic designs are generally more robust against adverse weather conditions than aerostatic platforms, increasing their reliability in various operational scenarios. Additionally, many modern aerodynamic platforms, such as SoftBank’s “Sunglider” [77], can be entirely solar-powered, enabling them to remain operational for extended periods, ranging from weeks to months. However, these platforms come with some limitations. Their payload capacity is generally smaller compared to large aerostatic platforms, which can affect the scale and complexity of their communication systems. Moreover, the development and manufacturing costs of aerodynamic platforms are higher, reflecting their advanced engineering and specialized materials.
Hybrid platforms
Hybrid designs aim to combine the strengths of both aerostatic and aerodynamic platforms, offering a balance between payload capacity, stability, and mobility. They can carry significant payloads like aerostatic platforms, while also benefiting from the mobility and weather resilience typically associated with aerodynamic platforms. This combination makes hybrid solutions suitable for a wide range of applications, including scenarios requiring both high payload capacity and the ability to reposition dynamically. However, these advantages come with notable challenges. The complexity of hybrid designs makes them costly to develop and manufacture. Sophisticated control systems are required to manage the dual mechanisms for lift and propulsion, adding to the operational complexity. Moreover, while hybrid platforms mitigate some disadvantages of pure aerostatic or aerodynamic designs, they often inherit certain drawbacks, such as higher sensitivity to adverse weather compared to fully aerodynamic platforms and increased infrastructure requirements compared to simpler aerostatic systems.
IV-A2 Terminals
Platforms ultimately aim at delivering connectivity to network terminals, a.k.a, UEs. Given the broad spectrum of application scenarios outlined in Section III, it is evident that UEs extend far beyond conventional smartphones, and also include wearable devices, IoT sensors, robots, and UAVs. Terminals can be classified in terms of features like form factor, operational environment, and communication capabilities. In the following paragraphs, we will describe the main groups of terminals, namely portable consumer terminals, vehicular terminals, and fixed terminals, as well as their features, and specific requirements, as summarized in Table III.
Portable consumer terminals
Probably the largest group of UEs consists of portable consumer terminals, including smartphones, smartwatches, tablets, and laptops. These devices are designed for everyday use across urban, suburban, and rural areas, and require seamless handover between TNs and NTNs to ensure uninterrupted connectivity. The diverse range of capacity-demanding and latency-sensitive applications, such as gaming and video streaming, necessitates support for multiple frequency bands, compatibility with existing standards (e.g., fourth-generation (4G), 5G), and NTN-specific adaptations for satellite and HAPS links. Conventional mobile devices are further complemented by wearable and implantable terminals, such as fitness trackers, health monitors, and medical implants. These devices are typically ultra-portable, low-power, and designed for continuous operation. They often require only short-distance, low-data-rate communication with a UE’s hub device, such as a smartphone, relying on the ubiquitous coverage provided by the hub. Another significant group within portable terminals includes communication devices for first responders, disaster relief teams, and military personnel. These devices are specifically designed to operate in extreme conditions, providing reliable connectivity in areas where terrestrial infrastructure is damaged or nonexistent. They often require enhanced security and encryption features for sensitive communications, as well as high-capacity links to enable rapid data exchange in critical scenarios.
Vehicular terminals
Another distinct category of mobile devices comprises vehicular terminals, installed in cars, buses, trains, ships, and aircraft. These terminals operate under challenging conditions due to high mobility and diverse environments, including remote and maritime regions. Unlike handheld devices, they can accommodate more efficient power sources and advanced antennas, such as VSAT parabolic reflectors or phased array antennas.
very small aperture terminal (VSAT) systems typically feature parabolic dishes ranging from 0.75 m to 2.4 m in diameter, with some larger models, providing high-gain, narrow-beamwidth links ideal for point-to-point communications. The antenna size depends on required gain, operating frequency, and environmental factors. In contrast, phased arrays, such as those employed in Starlink or BGAN systems, are more compact and rely on electronic beamforming through controlled phase shifts across multiple radiating elements. This enables rapid, precise beam steering without mechanical movement, offering enhanced flexibility and robustness under vibration or harsh conditions. Modern designs integrate features like auto-acquisition, self-alignment, and automatic aiming, simplifying deployment and ensuring reliability during high-speed operations.
A related subset of vehicular devices includes unmanned vehicles, such as UAVs, self-driving cars, autonomous robots, and surveillance balloons. These systems combine high mobility with operational versatility across urban and remote areas. They require reliable, low-latency command-and-control links and high-capacity channels for real-time data transmission (e.g., video or telemetry). To maintain continuous connectivity beyond terrestrial coverage, NTN support—via platforms like LEO satellites or HAPS—is essential for ensuring global reach and service continuity.
Fixed UE terminals
Despite the widespread use of mobile UEs, fixed UEs remains an essential category of network terminals. These devices are primarily designed for stationary broadband access in rural and remote areas. Their antenna systems are similar to those used in vehicular VSAT terminals. Unlike mobile devices, however, they are not constrained by strict form-factor limitations, allowing for larger, more efficient antenna designs. Such equipment allows high throughput for broadband applications, the ability to support multiple simultaneous connections, and resilience to adverse weather conditions, particularly for satellite-based links.
A distinctly different subset of fixed terminals comprises IoT devices, including sensors, indoor consumer smart household appliances and industrial equipment. While industrial equipment and household appliances can be connected through industrial gateways or indoor access points, remote sensors rely on their own connectivity solutions. These sensors are typically compact, low-power, and deployed in large numbers. They require massive machine type communication (mMTC) connectivity, often in remote or hard-to-reach locations, and energy-efficient communication protocols to maximize battery life.
This diverse range of UE requirements in integrated TN-NTN systems highlights the critical role of platform adaptability and technology integration. Meeting these diverse needs necessitates a collaborative ecosystem of terrestrial, satellite platforms, and HAPS, each uniquely suited to specific terminal categories and operating conditions. Specifically, the TN provides the best capacity and latency for densely populated areas, but is constrained in coverage and mobility support in remote regions, while satellites are indispensable for remote and hard-to-reach areas, but suffer from higher latency and lower rates. At the same time, HAPS serves as a versatile middle ground, offering good coverage, capacity, latency, and mobility, especially in semi-remote areas or regions requiring temporary connectivity.
IV-A3 Gateways
Aside from NTN platforms, several essential components constitute the network infrastructure, including gateways and associated network elements that ensure connectivity, reliability, and efficient communication [81]. Gateways, often referred to as ground stations, serve as critical interfaces between TNs and NTNs. Strategically located to optimize coverage and capacity and minimize latency, gateways are typically unconstrained in terms of form factor and power consumption. They are equipped with a large antenna structure designed for higher frequency bands, and are capable of maintaining efficient uplink connections with NTN platforms. Depending on the overall architecture, gateways may perform functions such as signal processing, routing, and resource management. They also aggregate traffic, and facilitate integration of NTN components with the terrestrial core network.
IV-B NTN Links and Frequency Bands
Table IV: Comparison of frequency bands for NTN, in terms of range, throughput, latency, and interference sensitivity.
Band Range Throughput Latency Interference sensitivity
Sub-7 GHz 100s km Low (<<1 Gbps) Low Low
cm-Wave 10s–100 km Medium (1–10 Gbps) Medium Moderate (rain fade in Ku-band)
$m$-Wave 1–10 km High (10–100 Gbps) Medium High (severe rain fade)
Sub-THz <<5 km Ultra-high (100+ Gbps) Low Very high
THz <<1 km Tbps Very low Extreme
FSO 10s–100s km Ultra-high (Tbps) Very low Weather-dependent
To establish reliable NTN communication, various types of links facilitate data exchange between different network components, namely feeder, service, inter-platform, and backhaul links. These links differ in terms of function, propagation characteristics, and performance requirements, influencing the choice of frequency bands and transmission technologies.
  • Service links provide connectivity between NTN platforms and UEs. These links must accommodate a diverse types of devices, ranging from stationary to highly mobile terminals, as described in Section IV-A, posing challenges such as signal power constraints, Doppler effects, and interference management.
  • Feeder links connect NTN platforms directly to gateways, typically operating at high frequencies to enable high-capacity transmission. To ensure reliability, these links incorporate advanced error correction and signal processing to mitigate atmospheric disruptions.
  • Inter-platform links facilitate direct communication between NTN platforms, either within the same layer (e.g., satellite-to-satellite) or across different layers (e.g., satellite-to-HAPS or HAPS-to-UAV). By supporting direct routing and handover, these links reduce dependency on terrestrial infrastructure.
  • Backhaul links interconnect NTN gateways with core network facilities, typically utilizing fiber-optic infrastructure to ensure high-speed, low-latency data transport.
The performance of these links is heavily influenced by the choice of operating frequency bands. Each frequency range offers distinct propagation characteristics, affecting factors such as coverage, capacity, and resilience to atmospheric conditions. The introduction of 5G marked a significant shift in spectrum allocation, expanding the use of high-frequency bands such as in the mmWave spectrum to enhance network capacity. In parallel, NTNs operates in highly heterogeneous environments, requiring careful spectrum planning to address diverse communication needs. As 6G and future TN-NTN integration evolves, the selection of appropriate frequency bands will be critical in optimizing system performance.
Notably, in Release 17 the 3GPP formally integrated NTN into the 5G new radio (NR) framework, defining use cases within the sub-7 GHz spectrum, particularly for massive access and narrowband IoT applications. Release 18 further expanded NTN support, introducing enhanced 5G NR operations for handheld terminals in the 10 GHz and above range. These updates focused on improved spectrum management and coexistence strategies to ensure seamless integration between TNs and NTNs [75, 74, 82].
The following subsections examine the key frequency bands currently used in NTN communications and those with potential for future adoption depending on the type of link, as summarized in Table IV.
IV-B1 Sub-7 GHz
Also known as 3GPP frequency range (FR) 1, this well-established frequency range, spanning from 410 MHz to 7125 MHz [83], has been widely used in long term evolution (LTE) and NR networks. FR1 is valued for its excellent propagation characteristics, resilience to atmospheric conditions, and an extensive ecosystem of mature telecommunication equipment. An important advantage of this band is the possibility to establish direct satellite-to-UE links, as it was recently demonstrated by companies like AST SpaceMobile, Starlink and Lynk Global. However, its limited bandwidth and high congestion restrict the achievable data rates, making it less suitable for high-capacity applications.
IV-B2 mmWave
Referred to as 3GPP FR2, mmWave spans frequencies from 24.25 GHz to 71 GHz [83]. Given its large bandwidth and short orthogonal frequency division multiplexing (OFDM) symbols, this spectrum enables ultra-high data rates and extremely low latency, making it ideal for bandwidth-intensive applications such as enhanced mobile broadband (eMBB). mmWave also offers fiber-like speeds in fronthaul and backhaul communications, supporting network densification through small cells and high-directivity antennas.
Despite its advantages, mmWave has seen limited adoption in TNs due to its short range and high infrastructure costs. The severe free-space path loss, combined with high susceptibility to blockage from buildings, foliage, and human movement, requires a dense network of BSs, making large-scale deployment expensive. While small cell and advanced beamforming techniques based on MIMO helps to mitigate these limitations, deployment remains challenging, particularly in rural and remote areas [84].
The use of mmWave in NTNs for broadband connectivity has recently gained significant research attention. The studies in [13, 85] explored NTN applications in 6G, evaluating aerial- and space-based mmWave links to provide connectivity for ground terminals under various network configurations. However, expanding NTN into mmWave bands requires effective spectrum-sharing techniques to balance isolation, licensing costs, and coexistence with existing TNs. Moreover, large-scale propagation models must be re-evaluated, considering factors such as path loss, penetration loss (also through the atmosphere), and shadow fading, which significantly impact NTN protocol design [86, 87].
IV-B3 cmWave
Also known as 3GPP FR3, cmWave spans 7.125 GHz to 24 GHz, bridging the gap between sub-7 GHz and mmWave bands [88]. Unlike mmWave band, which suffer from severe propagation characteristics, FR3 offers improved coverage and efficiency, with channel properties closer to sub-7 GHz in the sub-12 GHz region than to mmWave [89, 90, 91, 92]. Propagation studies [93] indicate that FR3 communication at 10 GHz experiences approximately 9.5 dB higher path loss in LoS, and 9.7 dB in non-LoS, compared to 3.5 GHz; yet these losses remain manageable with advanced beamforming. This band also exhibits lower penetration loss than at mmWaves, making it a strong candidate for outdoor-to-indoor coverage [90]. Additionally, rain attenuation is negligible below 10 GHz, and reduced delay spread improves signal quality, while angular spreads remain consistent with sub-6 GHz deployments. A key challenge in FR3 networks is the coexistence with satellite services, particularly in the 12 GHz band, which is heavily utilized by LEO constellations such as Starlink [89]. Interference mitigation techniques, including advanced beamforming and ephemeris-based coordination, have been proposed to ensure harmonious spectrum sharing between terrestrial and satellite networks [90]. Dynamic spectrum access mechanisms and interference nulling strategies are also being explored to optimize spectral efficiency while maintaining reliable satellite communications [94].
IV-B4 Beyond-100 GHz: sub-THz and THz
The beyond-100 GHz region—covering sub-THz (92–174 GHz) and THz (0.1–10 THz)—offers extreme bandwidth for immersive communications and sensing [89, 95, 96]. Measurements at 140 GHz report 2.3 Gbps over 120 m, highlighting the rate–range trade-off and equalization needs [97]. Simulation studies also show that integrated sensing and communication (ISAC) in these bands can outperform 5G-spectrum baselines [98]. For NTN topologies, sub-THz has been proposed for backhaul between UAV-BSs and CPUs, with HAPS acting as the CPU to preserve LoS [99]. Feasibility analyses indicate that THz feeder/backhaul links around 10 Gbps are possible in light rainfall given careful link budgets and signal processing [100] Reconfigurable intelligent surfaces (reconfigurable intelligent surfaces) can further alleviate blockage [101].
Despite this potential, beyond-100 GHz faces severe constraints for NTNs: high free-space path loss, strong water-vapor absorption, weak diffraction, and hardware limits (low output power, thermal stability) restrict range and robustness [102, 103]. Consequently, sub-THz/THz should be viewed as complementary, beyond-5G options suited mainly to short-range, LoS inter-platform or backhaul links under controlled conditions—not for handheld service or long-range feeder links. Practical deployment will require multi-link coordination (e.g., fallbacks to FR2/FR3 or FSO), highly directive arrays with null steering, site/spatial diversity, and aggressive modulation/coding adaptation [95, 96].
IV-B5 FSO
FSO technology is a valuable approach for future wireless networks, offering cost-effectiveness, easy deployment, high bandwidth, and strong security. The study in [104] provides an in-depth analysis of FSO systems, highlighting their role in next-generation optical communication. However, FSO propagation is challenged by atmospheric turbulence and alignment issues, necessitating mitigation strategies to maintain link stability. The study in [105] evaluates a ground-to-HAPS-to-RIS-to-UE communication system, where FSO is used for the ground-to-HAPS link, while RF communication is established for HAPS-to-RIS-to-UE transmissions. Key performance metrics, such as ergodic capacity, average bit error rate (BER), and outage probability, are analyzed using the bivariate Fox-H function. Results indicate that pointing errors and shadowing significantly affect system performance, with larger zenith angles and severe shadowing leading to degradation. Heterodyne detection improves efficiency by enhancing capacity and reducing BER and outage probability.
Based on the above discussion, it is important to recognize that power limitations, operational altitude, and environmental factors play a crucial role in HAPS communication. Selecting the appropriate frequency band for each link requires balancing multiple considerations:
  • Power constraints: Higher frequency bands typically require more power for transmission, which can be challenging given the strict power budgets of HAPS.
  • Weather sensitivity: Atmospheric conditions, e.g., rain and turbulence, impact high-frequency bands more significantly than lower ones, affecting link reliability and QoS.
  • Coverage vs. capacity: Lower frequencies (e.g., sub-7 GHz and cmWave) offer wider coverage (with indoor penetration) with lower throughput, while higher frequencies (e.g., mmWaves, THz, FSO) provide higher capacity but require precise beamforming.
The design of efficient HAPS-based NTNs requires balancing trade-offs between coverage, capacity, and resilience. A hybrid multi-band approach can enhance adaptability by leveraging different frequency bands for specific link requirements. For example, they suggest the following configurations:
  • Service links may operate at lower frequencies (sub-7 GHz, cmWave), to provide broad coverage and reliable signal propagation [75, 74, 82].
  • Feeder links may operate at cmWave and mmWave bands, to enable high-capacity, long-range transmission [106].
  • Inter-platform and backhaul links require ultra-high-speed data transfer, and may operate at sub-THz, THz, and FSO frequencies for seamless network integration.
This multi-band, hybrid spectrum strategy ensures a balance between coverage, capacity, and system resilience, supporting the development of robust HAPS-assisted NTN architectures.
IV-C 3GPP Specified Architectures
As introduced earlier, in the 3GPP NTN framework [74], two primary platform payload configurations are defined: transparent (or bent-pipe) and regenerative. The choice of payload configuration determines the NTN architecture at a fundamental level, shaping its operational behavior. In the transparent architecture the platform acts as a transparent relay, forwarding signals directly without significant processing. A bent-pipe payload typically implements frequency conversion, analog filtering, and signal amplification. This straightforward design results in reduced payload complexity, lower power consumption on the platform, and generally lower development and deployment costs. However, the absence of onboard signal processing limits the system’s ability to handle dynamic traffic or execute advanced processing tasks. These functions are offloaded to the ground station, leading to a round-trip signal path (from the UE to the NTN platform and then to the ground station) which inherently increases latency.
The regenerative architecture involves onboard signal processing, including tasks such as demodulation, decoding, and routing, in addition to the functions performed in the bent-pipe configuration. This onboard processing capability enables dynamic traffic management, such as load balancing based on QoS requirements. Platforms with sufficient computational resources can execute some UE tasks independently of the ground station, significantly reducing latency for time-sensitive applications. However, this capability comes at the cost of higher payload complexity and power consumption. The advanced technology required for regenerative systems also increases development and operational expenses.
Both transparent and regenerative NTN configurations can be implemented in multiple ways, depending on the functionalities assigned to the platform. In a transparent (bent-pipe) architecture—used in 3GPP Rel-17 and Rel-18—the NTN platform acts as a radio repeater (with or without beam control) while the gNB remains on the ground; this limits payload complexity and enabled early deployment. Notably, the Rel-17 “Satellite Access Node” definition in 3GPP RAN4 still allowed some distribution of gNB functions between Earth and space while keeping RAN4 RF/performance specifications applicable. In contrast, Rel-19 introduces a regenerative (packet-processing) architecture in which a complete gNB is placed on the platform [107]. This enables native use of the $X_{n}$ interface over inter-satellite links (inter-gNB mobility, resource coordination, energy saving), on-payload packet switching and routing/handovers directly between satellites with minimal ground involvement, and reduced $U_{u}$ round-trip time (e.g., for random access and hybrid automatic repeat request (HARQ)). Regenerative payloads also provide the building block for Rel-19 “store-and-forward” (S&F) for IoT UEs, allowing continued operation if the feeder link is severed and even supporting UE-to-UE termination in space. These gains come at the expense of higher payload power and complexity. Earlier alternatives with only a gNB-distributed unit onboard (gNB-central unit on the ground) were found suboptimal due to heavy F1 signaling over the feeder link, lack of a standardized direct inter-satellite interface, and limited support for feeder-link switching and S&F. Consensus therefore converged on the full gNB-on-board approach.
Regardless of the chosen architecture, NTN platforms can be integrated into a multi-connectivity (MC) framework for uplink, downlink, or bidirectional communication [74]. In this context, a UE can connect simultaneously to both an NTN platform and a terrestrial access point, or to multiple NTN platforms. This approach improves throughput, enhances reliability, and ensures seamless service continuity in heterogeneous service areas.
The role of MC in integrated TN-NTN systems depends on the operational objective. Beyond conventional dual- or multi-node connectivity, secondary MC nodes may be activated on demand to boost throughput. For example, [108] proposed a secondary-node activation algorithm, demonstrating a 9.1% increase in maximum average per-UE throughput compared to systems without MC, and a 5.3% improvement over baseline algorithms.
In addition, MC enables dynamic traffic splitting between terrestrial and non-terrestrial nodes, allowing the network to balance load across layers and offload traffic from congested terrestrial nodes to NTN links. In [109], the authors introduced a framework for UE association, BS transmit power and activation, and bandwidth allocation between terrestrial and non-terrestrial tiers. Simulation results showed a 45% reduction in total power consumption and an approximately 250% increase in average throughput compared to standard 3GPP-based network models.
Within integrated NTN systems, HAPS can assume diverse roles: from serving as terrestrial network extensions or satellite signal repeaters to operating as full-scale network centers equipped with baseband units and edge data processing capabilities. These roles directly impact latency, energy efficiency, and coverage performance. For instance, [110] proposed a quasi-stationary HAPS to enhance inter-satellite connectivity via FSO links with LEO satellites. Simulations under extreme conditions, such as volcanic activity, demonstrated the feasibility of supporting optical inter-satellite communication.
Experimental evaluations further confirm the versatility of HAPS-based systems. In [111], an automatic gain control system enabled a HAPS wireless repeater compatible with 5G NR to maintain stable transmission at maximum power, ensuring wide-area coverage. Similarly, [106] analyzed different HAPS-based link architectures and showed that HAPS can reduce latency in UAV ad-hoc networks, enhance energy efficiency in UAV cell-free architectures, improve coverage probability in satellite cell-free networks, and expand the service area of macro base stations through integrated access and backhaul designs.
Nevertheless, real-world data highlight operational limitations. In [112], empirical analysis of Loon project flight tracking data revealed that while stratospheric balloons offer a cost-effective, delay-tolerant connectivity service, outages can last several hours, limiting the achievable service stability compared to planned terrestrial networks. Such constraints may explain the current trend toward fixed-wing HAPS designs in modern projects, which promise greater reliability and sustained performance.
IV-D HAPS-Based Network Trials and Deployments
Over the past decades, numerous HAPS projects have explored applications in communication, as well as surveillance, disaster management, and environmental monitoring. While some achieved promising results, many faced technological, financial, or regulatory challenges that led to their termination even at advanced stages.
A prominent example is the Loon project [76], initiated by X Development LLC (formerly part of Google), which aimed to extend Internet connectivity to underserved regions using stratospheric balloons. Loon demonstrated HAPS potential by providing emergency connectivity in Peru (2017) during floods and in Puerto Rico after a hurricane, and by launching a commercial Internet service with Telkom Kenya in 2020 [112]. The balloons, following a regenerative architecture, carried an e-NodeB and antennas to serve ground UEs and communicate with gateways. Despite these achievements, concerns over commercial viability led to its discontinuation in 2021 [76].
Zephyr, a fixed-wing solar-powered HAPS by Airbus [113], is a payload-agnostic platform providing persistent surveillance and communication. It supports both transparent and regenerative architectures, using either its in-house connectivity payload to extend 4G/5G networks or other compatible technologies, enabling platform-as-a-service. By combining communication, monitoring, and sensing functions, Zephyr offers opportunities for intelligent network optimization and QoS improvement.
Similarly, HAPSMobile, a collaboration between AeroVironment and SoftBank [77, 114], plans a network of Sungliders to provide connectivity to billions lacking wireless access and to support emerging 5G and IoT networks. SoftBank positions HAPS as “base stations in the sky” for wide-area coverage, recovery during large-scale disasters, and 3D connectivity for drones/UAVs in the 6G era, complementing its ongoing heavier-than-air development with lighter-than-air vehicles. No details found on the architecture. Pre-commercial stratospheric services are expected in Japan by 2026 [115].
V HAPS Technologies
Technology achievements enable HAPS to enhance connectivity, optimize performance, and ensure sustainable operations. As an example, the authors in [2] explored the application of advanced techniques such as massive MIMO and mmWave communication into HAPS systems, identifying these as promising directions for future research. Similarly, the authors in [116] discussed key technological developments, trends, and challenges associated with HAPS. Nonetheless, this work primarily adopts an aeronautical engineering perspective, and lacks a thorough analysis from a telecommunications point of view. In addition, the study was published in April 2016, and does not reflect the most recent developments in the field. In this section, we address these limitations, and present an updated and comprehensive overview of recent technological trends that support the large-scale adoption of HAPS in the context of 6G networks. Specifically, we focus on the design of channel modeling (Sec. V-A), resource management (Sec. V-B), user assignment and cell association (Sec. V-C), multi-link connectivity (Sec. V-D), mobility management and handover (Sec. V-E), antenna (Sec. V-F), channel estimation (Sec. V-G), interference management and beamforming (Sec. V-H), energy efficiency (Sec. V-I), and ML (Sec. V-J) techniques for HAPS.
V-A Channel Modeling
Channel modeling for HAPS is evolving with new technologies, improving accuracy in predicting signal propagation, especially based on advanced simulation techniques and real-time data. Network simulators often lack realistic path loss and fading modules. To address this, models for rain, cloud, gaseous absorption, and Ricean fading have been developed [117]. Using QualNet, the authors showed how different attenuation types affect the packet delivery rate of HAPSground links, improving simulation realism and accuracy. The researchers in [118] used ray tracing simulations in suburban and urban environments, including downtown Chicago, to study HAPS-to-ground signal propagation. They developed a basic route loss model that estimates total propagation loss by considering terrestrial clutter and free-space propagation. The model adapts to different aerial vehicle heights and elevation angles using two environment-specific factors, and the study also analyzes multipath characteristics to support system design and fast fading modeling.
The authors of [119] analyzed a channel model to determine the capacity of HAPS-to- HAPS links, considering polarization mismatches, atmospheric gases, rain, and clouds/fog. Their simulations revealed the path gain characteristics and channel capacity across different antenna types, platform heights, and carrier frequency scenarios.
The study in [120] modeled HAPS-to-satellite communication channels with a focus on ionospheric effects. It showed that plasma density variations influence the channel while magnetic field effects on the refractive index can be neglected. Using the Appleton-Hartree model and an adapted Chapman model, the authors analyzed how Total Electron Content impacts path loss, supported by extensive numerical results.
A wideband channel measurement campaign at Ku- and X-bands investigated propagation effects in satellite-to-mobile radio links in four settings: urban, suburban, rural, and woodland. Results showed low coherence bandwidth in suburban and urban settings, suggesting channel equalization for high-data-rate services [121].
A multi-state statistical channel model for HAPS communication in the Ka-band was proposed in [122] to account for tropospheric weather effects and large-area user mobility. The model uses three fading states—LoS, moderate shadowing, and deep shadowing—represented by a Markov chain. It also provides the probability density function of the received signal power, a method for generating time-series, and evaluates BER performance using the proposed model.
The authors in [123] showed Ka-band rain attenuation is about 10× higher than in C-band and strongly location-dependent, and presented an analytical model based on the International Telecommunication Union Radiocommunication Sector (ITU-R) empirical model to predict Earth–space attenuation for fully and partially rain-affected paths. For a 1-km cell, attenuation reaches 43 dB at 0.01% availability and roughly doubles as cell size increases. Site diversity was evaluated as a rain-fade mitigation technique for HAPS-Earth links using the ITU-R model with three Earth stations at different rainfall rates to predict rain attenuation statistics [124]. Results show that site diversity raises link availability from 97% to 99.9%, as indicated by improvements in carrier-to-noise ratio.
The study in [125] shows that HAPS mobile channels invalidate the terrestrial plane-wave assumption, requiring consideration of both elevation and azimuth spreads. Large delay differences, Doppler shifts, and strong spatial correlation further complicate analysis. By characterizing dispersion across time, frequency, and space, the authors develop a wideband tapped-delay-line HAPS channel model with environment-specific parameters to support simulation and verification of broadband HAPS systems.
The study in [121] examined coherence bandwidth and RMS delay spread of satellite channels at Ku- and X-bands across four propagation scenarios. It found lower delay spread in woodland areas, abrupt location-dependent variations in rural environments, and urban delay spread marginally lower than suburban due to fewer, stronger multipath components with smaller relative delays.
FSO communication provides several benefits, e.g., in terms of bandwidth, data rate, and licensing costs, compared to traditional technologies, as described in Sec. IV-B. However, it faces significant challenges due to its dependency on precise transmitter-receiver alignment and the physical properties of the medium. Along these lines, the authors in [126] focused on optimizing internal parameters for FSO systems such as wavelength, detector type, beam divergence, modulation, and power to maximize link range and availability.
V-B Resource Management
Given the recent relevance of HAPS in wireless communication, resource allocation (RA) has become one of the key areas of investigation, including the positioning of the aerial tier units within the network. While multiple parameters must be optimized to address the RA problem comprehensively, more practical solutions operate under specific circumstances and constraints, defining optimization objectives accordingly. In this subsection, we will consider various system models, from single-tier (based on HAPS) to multi-tier (also known as space-air-ground integrated networks) architectures.
V-B1 Single-tier architecture
In [127], the authors maximize the number of UEs receiving multicast in an orthogonal frequency division multiple access (OFDMA) frame under QoS constraints by decomposing the problem into two linear programs solved iteratively. Sub-problem 1 selects UEs and assigns frequency–time slots via three Lagrangian-relaxation methods. Sub-problem 2 allocates power, modeled as a continuous knapsack, solved optimally with a greedy algorithm. The scenario assumes a single HAPS serving ground user equipments in a defined area.
In a multi-platform setup with several HAPS serving an area, [128] exploits fixed user dishes with narrow beamwidths steered to their main HAPS to boost bandwidth efficiency. Two cases are studied—single-beam per HAPS and multi-beam per platform—and results show higher capacity when multiple HAPS share the same spectrum, with gains that scale as the number of users grows.
In multi-cell settings, both inter- and intra-beam interference are significant. In integrated networks, overlapping coverage between terrestrial base stations and HAPS-based BSs—especially with co- or adjacent-channel operation—induces inter-system interference, making effective interference coordination essential for future integration (see Sec. V-H).
In [129], a dynamic spatial–time–frequency coordination method between HAPS-mounted BSs and TBSs is proposed. Two service modes are considered: direct HAPSGUE links and indirect access where HAPS serve TBSs that, in turn, serve GUEs. UEs associate to HAPS (direct or indirect) or TBSs based on maximum received power. The scheme distributes traffic load and employs an ergodic search over allocation patterns across HAPS and terrestrial networks. Simulations show the best average per-cell throughput for user groups at any distance from HAPS.
Unlike the single-HAPS model in [129], [26] studies a HAPS network via stochastic geometry, modeling TBSs and HAPS as Poisson point processes and deriving a closed-form coverage probability for spectrum sharing. Results show HAPS cause minimal interference to TBSs, supporting spectrum sharing in integrated networks.
V-B2 Multi-tier architecture
In the space domain, LEO satellite systems like Starlink by SpaceX are already extending communication services to areas beyond the reach of terrestrial networks. In January 2024, Starlink launched its first six satellites equipped with direct-to-cell capabilities, enabling device-to-device connections with unmodified LTE phones, and providing text messaging services integrated with mobile operators’ core networks [130]. To achieve this, Starlink developed large, advanced phased arrays, measuring 2.7 m x 2.3 m, which feature highly sensitive receivers and powerful transmitters. These arrays allow the satellites to function as cell towers in space, capable of communicating with cellphones equipped with low-gain antennas and limited transmit power. The goal faces significant technological challenges, particularly in densely populated areas with many users. Addressing these requires advancements in both network architecture and RA. An integrated multi-tier SAGIN can provide major performance benefits through deployment across diverse nodes and links.
The study in [131] focused on RA optimization within a multi-tier architecture where LEO satellites provide wide-area coverage, while HAPS and full-duplex terrestrial relays cover hot-spot regions to enhance capacity. Additionally, GEO satellites were considered as routing nodes. The authors designed a hierarchical resource management framework for this system, and proposed an optimization method to maximize LEO-to-ground and HAPS-to-ground downlink throughput based on predictions derived from the regular orbit information of LEO satellites. In the proposed architecture, mobile UEs in hot-spot regions connect via HAPS or terrestrial relays, while other areas use direct LEO links. This approach boosts system capacity in hot-spots and lowers handover rates. Different frequency bands are allocated for space/air-to-ground, terrestrial, space-to-air, and inter-satellite links. Capacity analysis and simulations show that the multi-tier design greatly outperforms satellite-only systems. Additionally, dynamic radio resource optimization improves the average throughput of HAPS and LEO beams by over 10%. In this scenario, spectrum sharing is required if HAPS and satellites are designed to transmit on the same or adjacent frequency bands, or when deploying HAPS in a region already covered by a separate satellite system. A LEO-HAPS spectrum sharing technique under non-ideal spectrum sensing was studied in [132], where a multi-beam LEO served as the primary system and a single-beam HAPS supported GUEs without direct LEO access. A dynamic spectrum and power allocation strategy was proposed to maximize HAPS-GUE throughput with minimal QoS degradation for LEO-associated GUEs. Results showed that HAPS interference to LEO-associated GUEs is negligible and affects only a small fraction.
RA for a multi-tier cognitive satellite-HAPS network with non-orthogonal multiple access (NOMA) was studied in [133]. The model included a primary satellite and a secondary HAPS network sharing spectrum via underlay cognitive radio to serve terrestrial users. The sum-rate maximization problem for the secondary network was addressed under NOMA group connectivity limits, power constraints, and QoS requirements. The proposed joint RA algorithm (sub-channel and power allocation) outperformed orthogonal multiple access, especially as user or sub-channel numbers increased.
Introducing smaller aerial components, such as LAPSs like drones, into the network architecture adds an additional degree of freedom in deployment planning. While this approach can further improve coverage and system capacity, it also complicates RA due to the need for dynamic load balancing and 3D placement of mobile network nodes. An aerial heterogeneous cellular network with HAPS-based macrocells and LAPS-based femtocells was studied in [134]. Using stochastic geometry, the authors analyzed cell loads, outage probability, optimal LAPS placement, and downlink signal to interference plus noise ratio (SINR). Results showed that users associated with an LAPS BS experienced much lower outage probability, and the proposed LAPS positioning algorithm further reduced outages.
The study in [135] investigated joint RA and UAV 3D-location optimization in a cluster-NOMA network with a HAPS macro BS and multiple UAV small BSs. The goal was to maximize IoT uplink sum rate by decomposing the problem into spectrum/power allocation and UAV positioning (horizontal and altitude), solved iteratively. Results showed improved uplink sum rate and spectrum efficiency over benchmark RA schemes.
As previously mentioned, the SAGIN architecture is likely the most promising option for future communication networks, including 6G. Therefore, it is crucial to clarify the features that enable the coexistence and collaboration of HAPS and terrestrial systems within this integrated framework. In [136], a RA framework for an integrated satellite-HAPS-TBS network was proposed to improve GUE downlink throughput under power, access, and backhauling constraints. Two problems were formulated: a short-term problem optimizing access, backhauling, and power with a fixed HAPS location, and a long-term problem optimizing HAPS placement based on average user distribution.
The studied system model comprised three tiers: a satellite in the space tier, multiple HAPS in the air tier, and several TBSs in the ground tier, with gateway feeder stations for backhauling. Based on OFDMA, spectrum was divided into resource blocks, with each user served by one BS using one RB, ensuring no intra-cell interference but allowing inter-cell interference among TBSs. No inter-cell interference was assumed for satellites and HAPS due to distinct frequency sets. Simulations showed a notable average data rate gain compared to two baselines: (i) joint optimization with uniform power, and (ii) optimized HAPS placement with random access/backhaul under uniform power.
From the literature, it is clear that aerial network mobility boosts capacity and flexibility but complicates optimization due to 3D positioning, dynamic load balancing, and spectrum diversity. Effective interference management is vital for stability, while a well-designed RA strategy can mitigate inter-tier interference and enable smooth integration with terrestrial networks.
V-C User Assignment and Cell Association
Unlike stationary TBSs, HAPS offer flexible deployment that can be optimized to user distribution. The study in [137] examined HAPS deployment for maximizing network transmission rate while meeting service requirements, where each GUE connects to the HAPS offering the best rate. A game-theoretic, self-organized optimization using a restricted spatial adaptive play algorithm was proposed and shown to converge in simulations. Although simplified compared to full SAGIN architectures, the results highlighted the potential of HAPS for self-organized deployment optimization.
To meet 6G data rate demands, [138] studied user selection in a downlink NOMA-based multi-user HAPS system. A joint user selection and correlation-based pairing algorithm was proposed to maximize beamforming gain, along with a polar-cap codebook design for limited-feedback HAPS channels under Rician fading. Simulations showed superior data rates over existing NOMA schemes. However, the model did not consider integration with terrestrial/satellite networks or support for aerial users, which are essential for 6G.
The user association problem in the context of computational task offloading was studied in [139]. The authors considered a disaster scenario where TBSs have been destroyed, and two-tier UEs, including both GUEs and UAVs, with limited computational capabilities offload disaster information to HAPS and/or LEO satellite servers for further processing. The study aimed to jointly optimize uplink fronthaul capacity and computation performance by modeling the problem as a weighted 3-uniform hypergraph, where each hyperedge weight represents the combined fronthaul uplink and computing rates. A local search-based algorithm was developed to find a near-optimal 3D hypergraph matching with maximum total weight.
The model considered a LEO satellite with an enhanced MEC server providing backhaul to multiple HAPS, which cover disaster areas and offload computation, while UEs use orthogonal sub-channels within shared Ka-band RBs, avoiding uplink interference. In simulations with 6 HAPS and 20 UAVs over a 0.25 km² disaster area, the proposed hypergraph search algorithm achieved significant system sum rate gains over greedy search, showing strong practical potential.
In [140], HAPS-mounted BSs were studied as backup transmitters in areas with strong terrestrial coverage. The user association problem between a TBS and a HAPS was addressed under outdated channel state information (CSI), with a DRL-based algorithm designed to maximize downlink sum-rate without requiring HAPS-user CSI. The TBS-user CSI was updated only for users served in the previous slot, and scenarios with imperfect CSI were also considered. Simulations confirmed the algorithm’s near-optimality, stability, and scalability despite CSI limitations.
A vertical heterogeneous network (VHetNet) comprising one GEO satellite, one HAPS, and several TBSs was examined in [141]. In this setup, both the HAPS and TBSs are equipped with multiple antennas, enabling them to simultaneously serve multiple GUEs through RF links. The connection between the satellite and the HAPS is established via an FSO link. In this scenario, when a GUE is served by the HAPS, the required data is first transmitted from the satellite to the HAPS, and then relayed to the GUE. The system model employs a space division multiplexing scheme, where all RF links share the same central frequency. Consequently, the performance of the proposed architecture is highly dependent on the intra-cell and inter-cell interference among TBSs, intra-cell interference within the HAPS, and the interference between HAPS and TBSs. The study formulated a joint optimization problem for associating GUEs with either TBSs or the HAPS and designing corresponding beamforming vectors, aiming to maximize sum-rate under constraints of HAPS payload connectivity, FSO backhaul capacity, and power limits of HAPS/TBSs. An iterative algorithm optimized parameters step by step, and simulations demonstrated that integrated satellite-HAPS-terrestrial networks can significantly enhance the sum-rate.
As heterogeneous integrated networks face load imbalance from uneven user distribution and mobility, [142] proposed a priority-based load-balancing scheme for a two-layer downlink SAGIN using network slicing to mitigate underutilization and overload across cells. The proposed two-layer radio access network (RAN) includes TBSs in the terrestrial layer (priority for user association) and UAVs/LEO satellites in the non-terrestrial layer. Overloaded TBSs offload traffic to the non-terrestrial layer via RAN slicing into high-throughput, low-delay, and wide-coverage slices, with lower-priority slices preferentially offloaded. The optimization problem jointly considered throughput, delay, coverage, user association, resource allocation, and UAV deployment. Results showed improved slice throughput and average delay. While HAPS were not included, the study tackled the critical load-balancing challenge in SAGIN.
In [143], load balancing for 5G integrated satellite-terrestrial networks was studied using multiple radio access technologys. The approach handled both intra- and inter-RAT UE offloading from overloaded cells by introducing the radio resource utilization ratio as a common load metric, enabling efficient traffic redistribution without overloading target cells. Simulations showed improved throughput and QoS. While limited to a GEO satellite, the metric can be generalized to other NTN components, including HAPS.
In [144], load balancing was studied as a joint radio resource management problem—covering bandwidth allocation, UE association, and power control—in an integrated TN-NTN rural downlink network with multiple macro TBSs and LEO satellites. The proposed framework maximized the sum of log throughput across all UEs, improving overall network capacity.
The framework tackled the optimization sequentially—first user association and bandwidth allocation under fixed power, then transmit power with the first two fixed. Simulations showed notable throughput gains over the maximum reference signal received power (RSRP) baseline and confirmed that integrating TNs and NTNs in rural areas can greatly reduce out-of-coverage UEs.
In [109], a framework for bandwidth allocation, user association, and power control was proposed to balance capacity and energy consumption by adapting to traffic conditions. The approach improved mean throughput during high-traffic periods and reduced power consumption during low-traffic periods.
Based on the reviewed studies, user association in SAGINs with HAPS-mounted BSs integrated with TNs still requires further research. Future work should address scenarios with a realistically large number of mobile ground and aerial UEs, while accounting for practical constraints from interference management and load balancing.
V-D Multi-Link Connectivity
As mentioned earlier, some emerging 6G applications, such as aerial vehicle control and autonomous driving, impose stringent requirements in terms of both latency and connection reliability, a paradigm referred to as URLLC. In this scenario, multi-link connectivity (or MC) has shown the potential to meet the URLLC service requirements [53]. Specifically, MC architectures can be categorized at the BS level (i.e., the user connects to multiple BSs within the same network, usually managed by the same operator), at the network level (i.e., the user is connected to different networks, possibly from different operators), or at the technology level (i.e., the user connects via different RATs) [54]. In this paper, we mostly focus on the first category, which also includes techniques such as coordinated multi-point (CoMP), which gained significant importance in LTE-advanced and 5G, where multiple BSs within the same network simultaneously serve the same user.
V-D1 MC at the BS level
In [145], the authors illustrated the benefits of MC for latency-sensitive applications like video streaming over local wireless network. However, the system model is limited to a local network scale, and cannot support high-speed transportation systems, such as high-speed trains. To fill this gap, the study [146] focused on dynamic multi-link scheduling for cellular networks serving high-speed train passengers. The study aimed at optimizing the use of existing public cellular network resources to provide more reliable and stable connections by minimizing latency and packet loss rates. Their findings demonstrated improvements in transmission time and bandwidth compared to various baseline methods.
In [147], air-to-air communication was studied using a 3D model integrating aerial BSs, aerial UEs, TNs, and GUEs. Two UAV types were considered: aerial BSs connecting via HAPS/TBS, and aerial users relaying data to GUEs. A joint-transmission CoMP scheme was proposed where each aerial user is served by four aerial BSs forming a tetrahedral cell. The study derived data rates and coverage probabilities, and introduced a frequency allocation scheme to reduce inter-cell interference. Compared to the binomial-Voronoi dynamic cooperation set, the approach achieved similar performance with much lower complexity and signaling overhead.
In [148], aerial UE connectivity was studied using CoMP transmissions for UAVs served by TBSs. Two scenarios were analyzed—static hovering and mobile UAVs—with coverage probability bounds derived and handover rates evaluated via a custom mobility model. Results showed that CoMP transmission greatly improves achievable coverage probability.
In [149], an uplink CoMP scenario was considered where UAVs collect and forward signals from GUEs. A UAV placement and movement scheme was proposed to maximize average throughput for mobile GUEs, though UAV energy constraints remain a critical challenge.
In [54], an MC SAGIN algorithm was proposed for remote aerial vehicle (AV) piloting beyond LoS. The system model included LEOs, HAPS, and TBSs for AV communication, along with direct inter-AV links. However, it excluded LEO-HAPS and LEO/HAPS-TBS links, assuming both LEO and HAPS connect to the core network via their ground stations.
The transmission strategy in [54] used message cloning, sending multiple copies via all available links for independent decoding at the AV. The model incorporated practical antenna setups, unreliable backhaul, buffer queues, and fading effects. Reliability and availability of different MC configurations were evaluated through simulations, with brute-force optimization used to find the optimal MC path ensuring QoS. Results showed that LoS interference strongly impacts link performance, and MC is vital for safe operation of certain AVs.
A latency-sensitive scenario was explored in [150], which addressed the issue of MC computational task offloading in NTN. The system model featured a HAPS and several UAVs providing on-demand MEC services to randomly distributed GUEs. The simulation results showed that MC offloading strategies can significantly reduce task computation delay.
V-D2 MC at the network level
In [151], network-level MC was studied for smart agriculture, where users simultaneously connected to satellite and terrestrial networks. Latency was the key metric, evaluated via drive tests using Starlink and two public 5G non-standalone TBSs. Results showed that MC greatly reduced latency compared to single connectivity with either satellite or TNs.
In [152], latency in an MC scenario with a UAV-based client was measured across four flight heights, comparing a single-operator setup with a dual-operator hybrid access system, which improved latency and reliability toward 3GPP targets. Similarly, [153] evaluated MC across two public LTE networks in urban, suburban, and rural settings with a UAV-based UE, showing that operator-specific performance variations make network-level MC effective for enhancing reliability.
V-E Mobility Management and Handover
Mobility management faces major challenges in seamless handovers, especially for LEO and HAPS systems. In [154], the standard 5G NR handover procedure was evaluated for a high-speed train in a LEO-based deployment versus a 3GPP urban macro TN. Simulations showed radio connection failures during handovers were 10× higher and outages 5× longer in the NTN scenario.
The authors in [155] proposed a platform that emulates hybrid satellite–terrestrial networks, to simulate TN/NTN handover procedures. This approach overcomes the need to physically launching satellites, and provides a cost-effective time-efficient simulation environment prior to real-world implementation.
In [156], three inter-satellite handover methods were compared for space-to-ground performance: Closest Satellite, Max Visibility, and SINR-based. The first two improved spectral efficiency and reduced delays, while the SINR-based approach minimized the number of handovers.
In [157], handover performance in a 5G NR network over LEO satellites with Earth-moving cells was studied. An intra-satellite mobility solution was proposed that avoids UE radio measurements by exploiting predictable satellite trajectories and antenna gains. Simulations in rural and urban scenarios showed reduced service failures and unnecessary handovers, with improved downlink SINR.
In [158], Event A4 was proposed for mobility management in LEO satellite networks, since minimal RSRP variation in NTNs makes A3-based handovers ineffective. Unlike A3, which compares serving and neighboring cells, A4 triggers when RSRP exceeds an absolute threshold. Simulations showed A4-based handovers outperform A3 in reliability and robustness, though determining cell borders remains difficult in NTNs, especially for fast-moving UEs.
In [159], an angle-based handover method was proposed for LEO networks, where the satellite elevation angle is calculated from its separation with the UEs and cell center. This approach was shown to provide reliable, continuous connectivity for UAVs in remote and challenging areas.
In [160], a graph theory-based algorithm was proposed for UAV handovers between terrestrial and satellite networks. By adapting in real time to UAV mobility and network changes, it minimizes latency and disruptions. The algorithm considers RSRP, signal to noise ratio (SNR), and elevation angle to optimize handover timing and ensure smooth connectivity.
The authors in [37] developed various handover techniques for an integrated LEO-HAPS network, and evaluated the handover delay and signaling overhead. Specifically, they formulated the problem of determining the optimal handover triggering time, which was analytically solved using Lagrange duality.
In [161], a mobility management solution was proposed for integrated HAPS-ground networks, allowing aerial users to connect to either terrestrial BSs or a HAPS. The approach focused on minimizing unnecessary handovers while ensuring service continuity and quality of experience.
In [162], an adaptive handover scheme with cooperative transmissions was proposed for HAPS systems to mitigate outage probability and long service interruptions. The algorithm adaptively selects the platform with the highest channel gain based on user mobility and real-time channel conditions, improving reliability and service continuity.
In [163], the impact of HAPS vertical and swing movements on handover performance was analyzed. A refined ground coverage model and analytical expressions were developed, showing that path loss strongly affects coverage, with swing movement exerting a greater influence on handover probability than vertical movement.
In [164], an adaptive handover algorithm was proposed for quasi-stationary HAPS, addressing the limitations of fixed-threshold methods. Using a least mean square prediction-based power model, the scheme adjusts handover parameters according to terminal speed and the confidence levels of expected signal strength, improving reliability under HAPS mobility conditions.
Furthermore, a new handover approach based on convolutional neural networks was introduced in [165]. The authors modeled the handover process as a directed graph, where a user with (partial) knowledge of its future signal strength can identify the optimal handover sequence. The CNN was trained to learn near-optimal handover strategies from historical signal strength data.
V-F Antenna Technologies
Integrating advanced antenna technologies, such as massive multiple-input multiple-output (mMIMO), spatial multiplexing, digital beamforming, and RISs, into HAPS systems can improve network capacity and efficiency across a range of diverse and challenging environments.
V-F1 MIMO-like antennas
The authors in [166] proposed a cylindrical mMIMO system for HAPS. They showed that the cylindrical mMIMO shape provides up to 2.1 times higher capacity compared to traditional planar arrays, with an extended coverage of up to 100 km in radius, and improves the SINR by approximately 10 dB at the cell edge.
In [167], a vertical uniform linear array sector-based cylindrical antenna was proposed for HAPS, combined with a NOMA clustering method to mitigate strong geographic user correlation. An optimization algorithm for power distribution was introduced to maximize energy efficiency, spectral efficiency, and QoS using successive interference cancellation. Simulations showed that spatial correlation strongly affects spectrum and energy efficiency in multi-antenna HAPS systems.
In [168], a two-stage precoding system for TN-HAPS spectrum sharing was proposed using phased array antennas. The first stage forms nulls toward TN cells, and the second directs beams to HAPS users. With a 196-element cylindrical array, the approach improved SINR for TN users by up to 10 dB over a conventional minimum mean square error (MMSE) precoder.
In [169], RF precoder design was formulated as an average signal to leakage interference and noise ratio (SLNR) maximization problem using a reduced-column discrete Fourier transform (DFT) matrix. The derived lower bound was converted into a trace problem for initial precoder design, and a greedy algorithm was proposed to update the RF precoder, achieving superior performance over existing methods.
In [170], a hemispherical antenna array for HAPS was proposed to maximize gain by aligning users with specific antenna elements. Using analog beamforming, user data rates were estimated from steering vectors, while an antenna selection algorithm and Bisection-based power allocation optimized performance. Results showed cumulative data rates up to 14 Gbps.
The authors in [99] proposed an algorithm to mitigate inter-cell interference between GUEs and neighboring UAV-BSs. The scheme operates in two stages: GUEs send messages to UAVs in the sub-6 GHz band, and UAVs then allocate resources and beamform signals to a HAPS in the sub-THz band acting as a CPU. An optimization problem was formulated to maximize the minimum SINR by optimizing power allocation and UAV positions.
In [171], multi-beam antenna technology was used to enhance coverage by adaptively adjusting beam widths within elliptical cells, equalizing signal quality and reducing the number of cells while maintaining coverage and link quality. The study also showed that integrating multi-beamforming with space-time block coding can further increase capacity without extra bandwidth.
The authors in [172] evaluated the advantages of using a triangular (rather than square) lattice for phased arrays in HAPS systems operating in the 5G NR n257 and n258 frequency bands. By considering a 64-element planar array, the results demonstrated that the triangular lattice array offers superior performance in terms of array gain, average sidelobe level, and mutual coupling, making it a more effective choice for 5G mMIMO systems.
V-F2 RIS antennas
The use of RISs can enhance signal quality, coverage, and LoS connectivity at low cost, and when combined with mMIMO, improve data rates and reliability [173]. In [174], a HAPS-RIS system was studied where RISs serve as amplify-and-forward relays to bypass blockages. Evaluation via BER, capacity, and outage probability showed significant performance gains, confirming the potential of integrating RISs with HAPS.
Another example of a HAPS-RIS system was considered in [175]. The RIS was mounted on the HAPS side for reflecting signals from GUEs to an upstream satellite. The authors used DRL for the joint optimization of RIS phase shifts and transmit beamforming, which significantly improved system efficiency.
In [176], a RIS was installed on a HAPS to provide extensive LoS and full-area coverage for UAV-BSs deployed in response to sudden urban traffic spikes. The combined optimization of RIS phases, array-partitioning algorithms, and positioning enhanced the energy efficiency of each UAV-BS.
In [177], an energy-optimization algorithm was proposed for HAPS, enabling mode switching to reduce consumption and extend loitering duration. The modes include a low-power HAPS-RIS mode for passive communication and a HAPS-SMBS mode for enhanced computing, caching, and communication. The system prioritizes the passive mode, activating higher-energy modes only when additional capacity is required.
V-G Channel Estimation
As 6G moves to high-frequency bands (cmWave, mmWave, sub-THz), channel conditions worsen due to path loss, fading, and interference, making accurate estimation essential for reliable connectivity and low latency. Advanced techniques such as beamforming, mMIMO, and ISAC also depend on precise CSI. In [178], channel estimation errors in an integrated satellite-HAPS-terrestrial system were analyzed, and a NOMA-based power allocation scheme was proposed to meet QoS requirements.
In [179], a HAPS-LEO architecture was studied to address channel estimation and synchronization challenges. A CNN-based estimator was proposed to mitigate carrier frequency offset from residual Doppler effects, with results showing improved mean square error and BER performance over benchmark methods.
The study in [180] evaluated LTE downlink performance over HAPS, comparing least square and MMSE channel estimation under a Ricean fading model, and found that GUEs at low elevation angles suffer degradation from multipath and Doppler. Similarly, [181] analyzed a single-carrier LTE HAPS channel, showing that elevation angles above 40° yield optimal conditions, with performance influenced by bandwidth, modulation, and Doppler shifts. More recently, [182] compared least square, MMSE, and MMSE+DFT estimation methods for HAPS-based OFDM systems.
In [183], sparse HAPS channels were estimated using a superimposed pilot sequence method that leverages prior knowledge of UEs’ locations. The model featured a HAPS-based beamforming multi-antenna BS serving mobile GUEs, with backhaul via ground stations and satellites. Numerical results examined the influence of antenna configuration, channel sparsity, SNR, and interferer density.
In [184], a VHetNet combining GEO/LEO satellites, HAPS, and TN was studied with full-duplex RIS-assisted backhauling over HAPS channels. A graph attention network-based channel estimator was proposed to address channel variability and hardware impairments, requiring estimation at only one transmitting node without on-off control or half-duplex switching. Simulations confirmed strong performance across diverse channel and hardware conditions.
V-H Interference Management and Beamforming
In [185], stratosphere-to-Earth co-channel interference from HAPS was analyzed, with a methodology to assess its impact on TNs through performance degradation. The study accounted for HAPS mobility, realistic azimuth/elevation distributions of terrestrial microwave links, and progressive HAPS loading, providing a comprehensive framework for interference evaluation.
The study [186] examined a 3D mMIMO system where a HAPS mounting a horizontal planar antenna array hovers in the sky, serving multiple GUEs. The study modeled intra-cell uplink interference using Beta distributions, and demonstrated that adjusting the altitude of the array can significantly reduce interference, thereby improving key performance metrics such as signal to interference ratio, coverage, and throughput.
In [94], interference between terrestrial and satellite platforms was studied, and a beamforming method using ephemeris data was proposed to dynamically suppress interference. By creating spatial nulls toward visible satellites, the technique protected satellite uplinks while maintaining terrestrial downlink performance.
To expand the operational coverage area for UAVs in a segregated airspace, Ku-band links can be utilized. In [187], the authors proposed a new conformal phased array antenna, developed with ultra-lightweight aerogel material, to support this architecture while managing interference. This antenna design offered high performance at low size, weight, and power.
In [171], a space-division multiple-access scheme was proposed with a multi-beam antenna and beamforming algorithm that adapts array weight vectors to improve performance and reduce interference. An improved elliptical cell design balanced SNR by broadening inner beams and narrowing outer ones, expanding inner cell coverage while shrinking outer areas. This reduced the number of required cells while maintaining coverage quality. The integration of space-time block coding with multi-beamforming further enhanced coding and diversity gains without extra bandwidth.
In [188], a dual-polarized shared aperture antenna for Ku- and Ka-bands was proposed, developed through a four-step derivation process using slot aperture coupling to achieve broadside radiation in both bands. Based on array factor theory, the optimal inter-element spacing was determined, resolving the trade-off between beam steering angles and array spacing in dual-band systems.
In [189], RA for HAPS-to-terrestrial and HAPS-to-satellite uplinks was optimized by addressing interference from uplink back lobes affecting GUEs. Unlike traditional zero forcing, which allocates random nulls, the proposed method generates targeted nulls based on GUE locations using a codebook of antenna patterns tied to weight matrices, enabling dynamic adaptation to reduce interference and improve communication quality.
In [129], traffic-aware RA was proposed for integrated HAPS-terrestrial systems, addressing the inefficiency of static methods that ignore load variations. The scheme adapts RA through interference coordination based on traffic distribution and network topology, achieving higher throughput than fixed strategies.
In [190], uplink performance in a multi-HAPS system jointly covering the same cell area was studied. To address the uplink bottleneck, the work explored spectrum reuse and interference mitigation using space-time diversity techniques, including MIMO and SIMO receive diversity.
In [191], a multi-cell HAPS system was analyzed, where capacity is limited by feeder link bandwidth. To enhance spectrum efficiency, a multi-gateway approach was proposed, allowing gateways to share feeder link frequencies to reduce interference and increase capacity. Building on this, [192] introduced interference cancellation techniques for the multi-gateway feeder link, further improving transmission quality and overall communication performance.
To further reduce interference, the work in [193] introduced a null-sweeping technique where the HAPS dynamically adjusts null points, while terrestrial BSs schedule GUEs accordingly. This coordinated approach improved the SINR for GUEs without degrading the HAPS link quality, resulting in up to a 15% increase in ground system capacity.
V-I Energy Efficiency
In [194], an adaptive link-state advertisement routing algorithm was proposed for integrated HAPS/satellite networks to support energy-efficient emergency communications. Designed for slow-flat Rician fading channels, the method estimates link information and selects the most energy-efficient path, reducing terminal power consumption while preserving connectivity.
In order to improve energy efficiency, the authors in [195] proposed a low-power-consumption decoding and forward relay system via HAPS. The results demonstrated that, in certain channel conditions, a relay-based system can outperform a non-relay-based system in terms of energy efficiency.
In [73], energy efficiency of a HAPS-based RAN was evaluated against a terrestrial RAN using a numerical model based on real traffic trends. Results showed that leveraging HAPS features—self-sustainability, high altitude, and wide coverage—can cut energy consumption by up to nearly 30%.
In [107], HAPS were studied as repeaters and aerial BSs, with energy efficiency assessed via consumption factor theory in single- and multi-cell scenarios using Monte Carlo simulations. Results showed both architectures achieve good downlink spectral efficiency, though coverage is constrained by GUE transmit power and antenna gain. Performance in uplink and downlink can be further improved with directed GUE antennas on LTE band 1.
In [196], a LEO satellite network using a solar-powered HAPS relay was studied with the goal of minimizing energy consumption for downlink data transmission. In the non-causal case, an energy-minimization problem was formulated and solved using future channel/energy knowledge, while in the causal case, a greedy online scheduling method was proposed, significantly improving throughput and energy efficiency.
In [197], an energy-optimized trajectory planning problem was formulated for high-altitude, long-endurance solar-powered aircraft. Using a 6-DoF dynamic model, the optimization maximized energy storage by combining battery and gravitational potential energy, solved via a nonlinear simplex search and interior point methods. The resulting optimal trajectories trained an Adaptive Neuro-Fuzzy Inference System for real-time planning, achieving a 15% gain in energy storage.
The study in [198] proposed a cost-effective power control algorithm for cognitive satellite-terrestrial networks with outdated CSI, where the satellite network opportunistically reuses the TN spectrum. This scheme was formulated as an energy efficiency maximization problem, subject to interference power constraints at the TN, and rate, energy and latency constraints at the satellites. The average transmit power was also determined by solving a nonlinear concave optimization problem.
In [199], a vertical HetNet with a HAPS as an SMBS, alongside an macrocell base station (MBS) and multiple small base stations, was studied to improve energy efficiency. An algorithm was proposed to minimize consumption by leveraging the HAPS and dynamically managing SBS sleep modes based on real-time traffic load, reducing power use by 30% while maintaining QoS. In [200], energy efficiency for data collection in a UAV network supported by a HAPS aerial BS was optimized. The approach dynamically clustered the target area, selected sink sensing devices, adjusted transmit power, and managed UAV operations. A centralized reinforcement learning (RL) framework based on proximal policy optimization (PPO) jointly optimized UAV trajectories and power, while an Affinity Propagation clustering algorithm identified optimal sink devices for device-to-UAV communication.
The research in [201] focused on energy-efficient RA in SAGINs for Internet of remote things (IoRT) using UAV relays. The goal was to improve energy efficiency by optimizing sub-channels selection and the uplink transmission power using Lagrangian dual decomposition, and UAVs deployment via sequential convex approximation. The proposed approach achieved more than 21.9% improvement in terms of energy efficiency.
The authors in [202] proposed a wireless backhaul solution using a RIS mounted on a HAPS to support UAV-BSs during sudden traffic surges in urban areas. To overcome signal blockages from urban obstacles, the aerial RIS reflects backhaul signals and adjusts the phase of each RIS element, position, and array partitioning, ensuring reliable connectivity. The use and optimization of RIS maximize energy efficiency and guarantees stable backhaul links for each UAV.
The study in [203] explored the use of HAPS in SAGINs to support data transmission between LEO satellites and smart devices for IoRT applications. The proposed approach jointly optimized RA and HAPS deployment to minimize the power consumption. By breaking down the complex problem into two sub-problems, the authors found a near-optimal solution. Simulation results showed a significant reduction in power consumption compared to conventional RA strategies, while still satisfying smart device rate requirements.
While MEC enables GUEs to offload resource-intensive tasks, the stringent network requirements of 6G applications may exceed the capabilities of terrestrial MEC servers. To address this issue, the work in [204] proposed to use aerial MEC servers to support GUEs and emergency communications. The authors introduced a hierarchical deep federated learning RA scheme for intelligent aerial-enabled smart city cyber-physical system, aiming at optimizing energy consumption while preserving task offloading privacy and maintaining QoS in dynamic environments.
V-J Machine Learning
AI has been utilized to tackle several critical challenges facing current and future wireless networks [205]. Notably, ML-based algorithms can be applied to HAPS to optimize key functions such as topology management, RA, handover management, and improve interference mitigation, ensuring more efficient, reliable, and adaptive communication networks [2]. These algorithms can dynamically adjust to changing conditions, reducing disruptions and maximizing service quality.
[…]
RL is increasingly applied to NTNs to optimize resource allocation, platform positioning, and trajectory control. By leveraging Markov decision process (MDP) frameworks, RL enables aerial platforms like HAPS and UAVs to autonomously adapt and enhance network performance under real-world constraints. The survey [206] reviewed RL applications for NTNs—including LEO satellites, HAPS, and UAVs—focusing on resource allocation, platform positioning, and trajectory control under NTN-specific constraints. Using MDP frameworks, it provided a taxonomy of current research, assessed study realism, and identified open challenges to narrow the simulation-to-reality gap.
In this context, the authors in [207] considered an heterogeneous network with UAV LAPSs and HAPS, aiming at optimizing fairness and QoS for GUEs. The proposed approach combines DRL with fixed-point iteration to determine the optimal UAVs positions, 3D trajectories, and channel allocation strategies based on real-time traffic load. A RL-based approach is proposed to optimize UAV base station positioning in emergency scenarios where rapid deployment is critical [208]. By leveraging radio signal sensing and PPO, UAVs dynamically learn and generalize positioning strategies across diverse user mobility patterns. Their numerical evaluations demonstrate the algorithm’s adaptability, consistently achieving throughput above 7 Mbps in simpler scenarios — close to the maximum achievable rate of 8 Mbps — and maintaining robust performance even in more complex, dynamic environments such as multi-cluster and hotspot user movements.
The study in [209] highlighted the advantages of optimizing user scheduling in space-HAPS-ground networks by ensembling deep neural networks. In particular, it was demonstrated that, when the ensemble size increases, the suggested approach performs better than all existing offline optimization strategies, with a sum-rate improvement of up to 26%.
Considering RL techniques, the authors in [206] addressed key challenges such as limited resources, integration with TNs, and the need for intelligent control of aerial BSs. They also provided a detailed taxonomy of RL applications in NTN. The study [210] develops a deep RL-based offloading strategy for HAPS-assisted vehicular networks. The partial task offloading problem is formulated as a MDP and solved using the deep deterministic policy gradient (DDPG) algorithm. Among all evaluated schemes, the DDPG-based approach achieves the highest task success rate, with only a 0.7% failure rate when the task size is 1.2 Mbits. The authors in [175] focused on an integrated satellite-HAPS-TN architecture to enhance network performance by using RISs for signal reflection. Specifically, they optimized transmit beamforming and RIS phase shifts, a complex problem due to RIS constraints. The study employed a DRL algorithm for effective online optimization, demonstrating superior performance, speed, and feasibility in real-time decision-making through simulations.
In [140], a deep SARSA-based RL algorithm was proposed for user scheduling in a wireless network with a HAPS backup. Modeled as an MDP, it operated with outdated CSI to maximize sum-rate while minimizing active antennas at the HAPS for energy savings. Results showed near-optimal sum-rate performance, and a heuristic method was also introduced for solving the scheduling problem.
In [211], a dynamic coverage path planning algorithm for multi-airship formations was proposed to overcome the limitations of static approaches for mobile platforms such as fixed-wing aircraft and LEO satellites. Using RL, the method stored experiences in a pool for rapid updates via semi-centralized exploration and centralized playback, with task-specific rewards improving network coverage.
In [212], a multi-agent deep Q-learning algorithm was proposed for HAPS power control to reduce outage probability while satisfying interference constraints with TNs. A double deep Q-learning variant was also introduced to avoid action-value overestimation. Simulations showed that agents collaboratively learned a near-optimal power control policy through effective state, reward, and training design.
In [213], an RA method for uplink HAPS-LEO communication using multi-carrier NOMA was proposed. The problem was formulated as a mixed-integer nonlinear program under power and channel constraints, and solved with a multi-agent double deep Q-network employing prioritized experience replay and priority sampling to handle dynamic resources and short satellite visibility.
In [214], HAPS were studied as flying BSs serving GUEs with NOMA in a downlink THz system. The method optimized power allocation, bandwidth, antenna beamwidth, and HAPS altitude to improve transmission rates while meeting data requirements. A DRL-based DDPG algorithm was applied to manage network dynamics and complexity, maximizing long-term rewards with continuous decision variables.
Recent work shows that federated learning (FL) can strengthen HAPS-enabled networks by enhancing privacy and lowering communication overhead. Approaches like support vector machine (SVM)-based FL for task offloading and hierarchical FL with HAPS as distributed servers improve model accuracy, energy efficiency, and reduce training time for large-scale applications.
For example, the authors in [215] address the minimization of energy and time consumption in mobile-edge computing-enabled HAPS networks. It formulates an optimization problem for user association, service sequencing, and task allocation, and proposes a SVM-FL algorithm to predict user associations without sharing private data. Simulations demonstrate up to 15.4% improvement in energy and time efficiency compared to terrestrial centralized methods. To enable global-scale FL in NTNs, a distributed hierarchical FL framework is proposed [216], HAPS as intermediate servers. Integrating LEO satellites and ground clients while using GEO and MEO satellites for model relaying, the framework enhances privacy, improves model accuracy, manages latency, and boosts scalability. Numerical results confirm its efficiency in training loss reduction and network optimization. Addressing the limitations of traditional ML in satellite communications, the study [217] introduces a federated learning framework that employs HAPS as distributed parameter servers for LEO satellite constellations. Proposed framework integrates a hierarchical communication architecture along with model dissemination and aggregation algorithms, significantly reducing training time from days to hours while improving model accuracy.
VI Open Research Challenges
The integration of TNs and NTNs through HAPS-enabled connectivity represents a transformative opportunity for future 6G wireless communication systems. However, this paradigm presents several challenges, including balancing the trade-off between coverage and capacity, optimizing radio resource allocation, developing efficient beamforming techniques, addressing mobility and handover issues, and ensuring seamless interaction between TN and NTN components. Moreover, the deployment of HAPS systems in dense urban environments requires careful planning to avoid interference and ensure efficient spectrum utilization, particularly when multiple HAPS and ground stations are involved. In the following, we state the hottest research challenges:
VI-1 Channel Modeling
HAPS in NTNs introduce unique channel modeling challenges due to their stratospheric operation (17–25 km altitude) and highly dynamic environment. Unlike satellites or terrestrial infrastructure, HAPS experience mixed terrestrial and space-like propagation conditions with high variability. Key challenges include channel non-stationarity, caused by platform movement, altitude shifts, and atmospheric changes, which impact Doppler shift, delay spread, and shadowing; and channel consistency, ensuring validity across wide coverage areas with diverse terrain and weather. Environmental effects—rain, clouds, atmospheric gases—become significant at mmWave and higher frequencies, leading to frequency-selective fading and attenuation. Moreover, antenna array impairments from platform instability complicate beamforming and degrade link reliability.
VI-2 Spectrum
Since HAPS systems share frequencies with satellites and terrestrial systems, spectrum allocation and interference management are also important issues. To prevent interference, particularly in highly-occupied frequency bands, HAPS networks must effectively coordinate their spectrum utilization. Efficient spectrum sharing and interference mitigation techniques, such as advanced beamforming, interference cancellation, and cognitive radio methods, must be developed to ensure reliable communication without signal degradation.
VI-3 Mobility management
The high mobility of NTN components demands dynamic 3D positioning and load balancing to optimize resource distribution, handovers, and power use. Continuity and latency pose additional challenges when HAPS link both satellite and terrestrial networks, making seamless, low-latency handovers difficult for time-sensitive applications like AR, autonomous driving, and industrial automation. Addressing these issues requires advances in multi-connectivity protocols and intelligent routing to adapt paths based on network state, traffic, and UE needs. Improved synchronization techniques are also essential to reduce latency spikes, and ensure reliable, responsive performance in integrated TN-NTN systems.
VI-4 Interoperability
Seamless integration of TNs and NTNs into a unified 6G network architecture requires overcoming significant standardization and interoperability challenges. Currently, terrestrial and satellite systems operate via distinct communication standards and protocols. Developed 3GPP standards provide a good base for interoperability, but many features of the protocol stuck requires optimization—or a new design all together— to accommodate for the larger delays and massive number of connected UEs. In short, the inclusion of HAPS as a central component of TN-NTN networks will require new protocols and standardization efforts to ensure compatibility across diverse systems. This includes designing common interfaces and protocols for satellites, HAPS, and terrestrial nodes, as well as ensuring that HAPS can dynamically interact with both space- and ground-based infrastructure in real-time.
VI-5 Energy
Energy efficiency is a major challenge for HAPS in NTNs, especially for future 6G systems operating under strict energy constraints. Sustainable operation requires energy-aware communication protocols and power management strategies, as HAPS rely on limited resources like solar power to support communication, sensing, and navigation. Key advancements include improved energy harvesting, battery storage, dynamic power allocation, and resource optimization. While efficient daytime solar harvesting is vital, storing energy for nighttime or low-sunlight conditions remains difficult, making enhanced storage technologies and advanced scheduling essential to extend mission duration and ensure uninterrupted service.
VI-6 Legal regulations
HAPS will share airspace with manned and unmanned aircraft, as well as space-bound vehicles, which raises regulatory and safety concerns about air traffic control, collision avoidance, and airspace management. The expanding role of HAPS in global communication infrastructure will require the evolution of international aviation and telecommunications regulations to ensure the safe integration of HAPS with existing air traffic.
VI-7 Deployment costs
The large-scale deployment of HAPS hinges on economic feasibility, requiring business models that offset high initial costs through mechanisms such as public–private partnerships, rural connectivity subsidies, or integration into smart city projects. Demonstrating long-term benefits—enhanced coverage, rapid deployment, and reduced infrastructure costs—is essential to justify investment. Overcoming regulatory, safety, and cost-effectiveness challenges will be critical to realizing the full potential of HAPS in future networks.
VII Discussion and conclusion
The evolution toward 6G requires tight TN–NTN integration. Operating in the stratosphere, HAPS can bridge terrestrial cells, LEO satellites, and UAV networks to deliver wide-area, reliable, low-latency connectivity. This survey synthesized HAPS use cases, deployment architectures (transparent vs. regenerative), spectrum options from sub-7 GHz to FSO, and key enablers (channel modeling, UE association, resource allocation, multi-link connectivity, mobility management, antenna technology, channel estimation, beam and interference control, AI/ML, and energy efficiency), and outlined open challenges. Unlike prior surveys, we positioned HAPS as intelligent network nodes—not mere relays—within standardized NTN frameworks and multi-band designs. Looking ahead, priorities include: (i) efficient multi-band strategies (FR1/FR3/FR2 + optical) with robust coexistence; (ii) cross-layer TN–NTN interoperability aligned with recent 3GPP NTN releases; (iii) AI-based control with reliability/energy guarantees; (iv) sustainability-driven platform and gateway design; and (v) progress on regulation and cost models to enable scale. With continued innovation, HAPS can become a practical cornerstone of 6G and a lever to narrow the digital divide.