### v2

# Bridging Earth and Space: A Survey on HAPS for Non-Terrestrial Networks

Abstract
[…]
HAPS are emerging as key enablers in the evolution of sixth-generation (6G) wireless networks, bridging terrestrial and non-terrestrial infrastructures. Operating in the stratosphere, high-altitude platform station (HAPS) can provide wide-area coverage, low-latency, energy-efficient broadband communications with flexible deployment options for diverse applications. This survey delivers a comprehensive overview of HAPS use cases, technologies, and integration strategies within the 6G ecosystem. The roles of HAPS in extending connectivity to underserved regions, supporting dynamic backhauling, enabling massive Internet of things (IoT), and delivering reliable low-latency communications for autonomous and immersive services are discussed. The paper reviews state-of-the-art architectures for terrestrial and non-terrestrial network integration, highlights recent field trials. Furthermore, key enabling technologies—such as channel modeling, artificial intelligence (AI)-driven resource allocation, interference control, mobility management, and energy-efficient communications—are examined. The paper also outlines open research challenges. By addressing existing gaps in the literature, this survey positions HAPS as a foundational component of globally integrated, resilient, and sustainable 6G networks.
Index Terms — HAPS, 6G networks, non-terrestrial network (NTN), terrestrial network (TN), ultra-reliable low-latency communication (URLLC), network integration, energy-efficient communications.
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
I Introduction
The integration of TNs and NTNs is widely regarded as a key enabler for achieving the ambitious goals of 6G networks — to “connect the unconnected” and “hyper-connect existing infrastructures.” While traditional TNs are foundational, they cannot solely meet the growing demands for coverage, capacity, and low latency required by emerging applications [32]. To overcome these limitations, NTN technologies have emerged as transformative solutions. NTNs leverage space-borne (satellites), airborne (HAPS), and other non-ground-based platforms (unmanned aerial vehicles) to extend connectivity beyond terrestrial infrastructure, enabling coverage in remote and underserved regions, as well as in mobile platforms, e.g., oceans, rural areas, or in-flight scenarios, respectively. A key distinction between conventional TNs and NTNs is the latter’s ability to decouple service coverage from the physical location of backbone infrastructure through backhauling via satellite or HAPS links, achieved using long-distance gateway connections or multi-hop inter-platform links.
HAPS operate in the stratosphere, complementing both TNs and other NTN nodes such as low Earth orbit (LEO) satellites and UAVs. They combine the advantages of these systems while mitigating their limitations. HAPS offer wide-area coverage with higher capacity and lower latency than LEO satellites, and lower energy consumption with longer endurance than UAVs. Their quasi-stationary positioning reduces dependence on ground infrastructure and ensures reliable line-of-sight (LoS) or near-LoS connectivity. This capability allows HAPS to enhance network performance in diverse environments, including urban, rural, remote, and disaster-affected areas [116]. Beyond bridging connectivity gaps, HAPS enable advanced services such as massive IoT, real-time autonomous systems, edge computing, and immersive technologies like augmented and virtual reality [162, 116]. These features establish HAPS as a key element in shaping next-generation global connectivity.
The main objective of this survey is to present a comprehensive overview of the state-of-the-art in HAPS networks, examining key applications, enabling technologies, and future research directions. We first outline the diverse roles of HAPS across several domains, including connectivity enhancement, network infrastructure and management, IoT and emerging services, advanced control and special applications, as well as energy and sustainability. The survey further examines TN-NTN integration, emphasizing the role of HAPS as a bridge between terrestrial and non-terrestrial systems. It outlines platform and terminal types within this framework, reviews TN-NTN architectures, and explores emerging frequency bands such as millimetre wave (mmWave) and centimeter wave (cmWave), underscoring their importance for enhanced connectivity. The survey then reviews key HAPS technical enablers, including channel modeling, resource management, user equipment (UE) association, multi-link connectivity, mobility management and handover, antenna technologies, interference control, energy efficiency, machine learning (ML) applications, and cybersecurity. It also addresses deployment and operational challenges, highlighting areas needing further research to fully realize HAPS potential. By covering these topics, the survey offers an in-depth understanding of technological advancements, design considerations, and open challenges in HAPS-based fifth-generation (5G) and 6G solutions, serving as a valuable resource for researchers and engineers in TN-NTN integration.
The literature for this survey was sourced through a manual, systematic search across key academic databases, including IEEE Xplore, Scopus, and Google Scholar, resulting in the selection of over 200 peer-reviewed studies. The search prioritized publications from the last decade (2015–2025) to capture recent advancements in HAPS technologies, with keywords such as "HAPS," "stratospheric platforms," "6G networks," "non-terrestrial networks" etc., and related terms in telecommunications and radiophysics. An initial screening was conducted based on titles and abstracts to identify potentially relevant works, followed by a full-text review to apply inclusion criteria focused on topic relevance to HAPS/NTN-based system models, and TN-NTN integration; methodological rigor, including empirical validation or theoretical consistency with standards like International Telecommunication Union Radiocommunication Sector (ITU-R) recommendations; and comprehensiveness in addressing challenges. Exclusions were made for a few non-peer-reviewed materials (arXiv preprints, press-releases and reports), relevant studies predating 2015, or those lacking substantive contributions, but ensuring a comprehensive yet focused corpus that supports the survey’s objectives.
The rest of this survey is organized as follows (see Fig. 1): Section II reviews the most relevant prior surveys including general studies in the area of non-terrestrial networks, studies focused on AI and ML applications and those specifically considering HAPS. Section III compares HAPS with other NTN nodes. Section IV highlights the most promising use cases for HAPS. Section V explores scenarios and architectures for HAPS deployment. Section VI examines the key technical enablers for HAPS. Section VII discusses the current challenges related to HAPS communications. Section VIII concludes the survey with findings and insights for future research directions.
Figure 1: Survey structure
II Review of Previous NTN-Based Surveys
Before diving into this survey, we first highlight how it differs from previous surveys on the NTN topic.
II-A General NTN Surveys
Several surveys analyzed NTNs, but most provided no in-depth examination of HAPS. Baltaci et al. [36] reviewed aerial connectivity for airplanes, vertical take-off and landing vehicles, and UAVs, providing a detailed analysis of use cases, requirements, and network architectures. It also examined aerial connectivity providers, including satellites and HAPS, as well as communication challenges. However, its coverage of HAPS is limited, and lacks examination of TN-NTN integration. Geraci et al. [73] explored the role of UAVs in future communication systems, covering use cases, network architectures, and research challenges. It reviewed potential optimization opportunities in sub-6 GHz and mmWave TNs to provide service to UAVs. Additionally, it included a performance analysis of UAV-to-UAV communications.
Azari et al. [32] studied TN-NTN integration in 5G and 6G networks, focusing on enabling technologies such as mmWave, IoT/ Mobile Edge Computing (MEC) integration, and ML. It also highlighted outcomes from field trials. However, it mostly emphasized LEO satellites and UAVs, rather than HAPS. TN-NTN integration was also examined by Geraci et al. [74], mostly focusing on LEO networks, spectrum challenges, multiple-input multiple-output (MIMO) modeling, and simulations to evaluate the performance of UAVs when served from either a massive MIMO-based TN or a LEO-satellite-based operator. The study also highlighted relevant third generation partnership project (3GPP) standardization efforts and solutions, such as three-dimensional (3D) radio access, mobility management, and network orchestration. Similarly, Rinaldi et al. [180] explored the evolution of NTN from second-generation to fourth-generation (4G), and discussed TN-NTN integration, mobility management, and 3GPP support for 5G new radio (NR). The survey also highlighted the role of NTN toward 6G for achieving global coverage and increased capacity.
Jamshed et al. [107] presented a comprehensive overview of NTN as a pivotal enabler for 6G networks, emphasizing their role in achieving seamless, intelligent, and sustainable global connectivity by integrating satellites, aerial platforms (UAVs, HAPS), and TNs to address the digital divide and support applications like IoT, navigation, disaster recovery, and interplanetary communication. The paper delved into enabling technologies, real-world applications illustrated through UAV/HAPS-assisted fronthaul/backhaul, security considerations, and future research directions, positioning NTN as essential part of resilient 6G ecosystems. Despite their comprehensive coverage of NTNs, these surveys only briefly mentioned HAPS as part of a broader NTN framework, and did not explore them in depth.
II-B AI-Focused Surveys
AI played an essential role in NTNs and TN-NTN integration, enabling intelligent resource management, optimization, and automation across various network layers. Several surveys analyzed the application of AI in NTNs, focusing on key technologies, ML-based algorithms, and network performance improvements. AI-based optimization strategies for TN-NTN integration were analyzed by Wang et al. [221], who discussed the potential of TN and satellite network integration to enhance global connectivity. This study provided an in-depth analysis of convergence drivers, network architectures, and key application areas. Kurunathan et al. [127] explored the use of ML for UAV-related applications, including image processing, ground sensors localization, real-time UAV control, scheduling and trajectory planning. Zhou, Sheng, Li and Han [252] reviewed 6G aerospace-integrated networks, analyzing architecture, satellite configurations, protocols, and key components, while also identifying future research challenges. It examined advanced technologies, with a focus on AI-driven resource management, and NTN research gaps. Iqbal et al. [101] also explored ML and AI techniques for addressing NTN-related challenges in 6G, highlighting privacy risks and optimization techniques such as deep learning, and deep reinforcement learning (DRL). Fontanesi et al. [68] provided a comprehensive overview of AI and ML applications in satellite communication (SatCom) and NTN, focusing on geostationary equatorial orbit (GEO), medium Earth orbit (MEO), and LEO systems, alongside with an extensive classification of the most important use cases, its related challenges and the conventional tools used to address them. Onboard vs. on-ground AI architectures were compared, emphasizing training paradigms (offline/online). Hardware solutions for ML deployment were reviewed, including radiation-tolerant chipsets. Long-term challenges and insights for future research concluded the study. Mahboob and Liu [146] categorized AI applications across current research directions covering physical (PHY) layer, data link layer and upper layer aspects such as computational offloading, network routing and traffic prediction. Existing works, research efforts, including ML and software defined radio testbeds, and practical challenges were reviewed, concluding with insights for future AI-driven NTN deployments.
Despite such a broad coverage, these surveys primarily focused on satellites and UAVs, with limited discussion of HAPS.
II-C HAPS-Related Surveys
Mozaffari et al. [162] provided a comprehensive tutorial on both HAPS and low-altitude platform stations in wireless communications, highlighting their role as aerial base stations and cellular-connected UEs. It examined key challenges such as 3D deployment, channel modeling, and energy efficiency, while exploring potential research directions. Additionally, it reviewed analytical frameworks, including optimization theory and ML, for addressing UAV-related issues. However, it did not consider TN-NTN integration, comparing instead UAV- and ground-based networks.
Kodheli et al. [124] reviewed advancements in LEO satellite and HAPS communications, capturing the state of the art in SatCom, and highlighting key developments such as new constellation architectures, on-board processing, and NTN integration. It systematically analyzed five critical areas: system design, air interface, medium access control, networking, as well as testbeds and prototyping. While the study covered HAPS, its primary focus remained on satellites, offering only a high-level assessment of HAPS without detailing their unique technological challenges. Giordani and Zorzi in [77] considered NTNs as a crucial enabler for 6G, enhancing connectivity through technologies such as mmWave and flexible beam steering. It complemented previous works by providing simulations that compared different integration scenarios, including TN-LEO, as well as TN, LEO, and HAPS integration. However, HAPS was considered only as a relay, without evaluating their potential as independent network nodes or fully-integrated NTN components. Karabulut Kurt et al. [116] took a more focused approach on HAPS, exploring their role in NTN communications, particularly in the context of mega-constellation systems for next-generation wireless networks. It examined key use cases, and focused on the potential of HAPS as super macro base stations, while also addressing regulatory considerations, alongside critical technologies such as channel modeling, resource management, handoff mechanisms, and swarm control. The survey also discussed AI applications, particularly in network topology management, resource allocation, and mobility optimization. Additionally, it reviewed emerging technologies such as reconfigurable intelligent surfaces for cost-effective deployment and faster-than-Nyquist signaling for spectral efficiency. While this survey provided a strong technological perspective, it lacked insights into TN-NTN interoperability and real-world deployment challenges. The role of HAPS in NTN was further explored by Arum, Grace, and Mitchell [30], who examined HAPS potential for coverage extension and capacity enhancement in TN-HAPS coexistence scenarios. The study focused on leveraging HAPS as an alternative to terrestrial systems, particularly for wireless communication in rural and remote areas where broadband access remained limited. It discussed techniques such as intelligent radio resource management, interference mitigation, and spatial optimization using array antennas, radio environment map, and device-to-device communication to maximize coverage while ensuring coexistence with TN. Additionally, the study highlighted past HAPS projects but did not explore new applications beyond coverage expansion, such as AI-driven optimization, network infrastructure and management, support for IoT, or TN-NTN integrated architectures. Abbasi et al. [11] outlined potential HAPS use cases and identified particular challenges which may arise due to integration of HAPS with conventional TNs and because of platform’s physical characteristics. The authors also investigated the usefulness of HAPS-enabled vertical heterogeneous networks through the estimating of cell-edge users’ capacity enhancement and the benefits of the cell-free communication system of aerial BSs with HAPS backhauling. Among the most relevant surveys to this work, i.e., [162, 124, 77, 116, 30, 11], none provided a comprehensive, up-to-date analysis of HAPS, covering both technological advancements and TN-NTN integration. The surveys in [162, 124] were still generalist NTN studies, offering only limited exploration of HAPS, focusing on use cases, technological capabilities, and network architectures, without detailing dedicated HAPS advancements. Although [77] considered HAPS in the NTN-enabled 6G ecosystem, it primarily focuses on relay-based NTN integration, without assessing HAPS as independent network nodes or their role in multi-layered communication systems. The survey [116] provided an extensive review of HAPS mega-constellations, regulatory aspects, and key enabling technologies, but offered little insight into TN-NTN interoperability, the opportunities and challenges it presents, and real-world deployment considerations. The work in [30] examined HAPS for coverage extension and terrestrial coexistence, but did not explore their broader integration into NTN architectures, advanced use cases beyond connectivity, or emerging technologies, especially those related to AI. While the paper [11] covered particular use cases and challenges it is more concise and less exhaustive in referencing. To fill these gaps, this survey provided a holistic and forward-looking review of HAPS, including TN-NTN integration, enabling technologies, real-world deployment challenges, and future research directions. Table I provided a comparative summary of this survey against the previous relevant literature.
Table I: Comparison of our survey against other HAPS-related surveys.
Survey Scope HAPS
focus
HAPS
use cases
HAPS
technologies
TN-NTN
integration
Challenges
[162] Examines HAPS and LAPS in wireless communications, identifying challenges and research directions but focusing on UAV-ground comparisons rather than TN-NTN integration. Secondary
[124] Reviews advancements in satellites and HAPS communication, highlighting key developments such as new constellations, on-board processing, and NTN integration, but focusing on satellite communications. Secondary
[77] Considers NTN for 6G, with simulations of TN-LEO-HAPS integration, but focusing on HAPS as a relay rather than independent nodes. Secondary
[116] Focuses on HAPS mega-constellations, regulatory aspects, and key enabling technologies, including AI and swarm control, but lacks TN-NTN integration insights. Primary
[30] Examines HAPS for coverage extension and coexistence with terrestrial networks, but does not explore broader use cases, advanced technologies, or TN-NTN integration. Primary
[11] Outline particular HAPS use cases and challenges in 6G vHetNets with some simulation validations. Primary
Ours Comprehensive review of HAPS, covering use cases, enabling technologies, TN-NTN integration, current HAPS-related projects, and future research directions. Primary
List of Frequently Used Abbreviations
3D
three-dimensional
3G
third generation
3GPP
third generation partnership project
4G
fourth-generation
5G
fifth-generation
6G
sixth-generation
AI
artificial intelligence
AR
augmented reality
AV
aerial vehicle
BER
bit error rate
BS
base station
CFO
carrier frequency offset
CNN
convolutional neural network
cmWave
centimeter wave
CoMP
coordinated multi-point
CPU
central processing unit
CSI
channel state information
D2D
Device-to-Device
dB
decibel
DDPG
deep deterministic policy gradient
DFT
discrete Fourier transform
DRL
deep reinforcement learning
EB
exabytes
eMBB
enhanced mobile broadband
FL
federated learning
FR
frequency range
FSO
free space optical
GEO
geostationary equatorial orbit
GUE
ground user equipment
GNSS
global navigation satellite system
GPS
global positioning system
HAPS
high-altitude platform station
HAPS-SMBS
high-altitude platform station mounted super macro base station
HO
handover
IoT
Internet of things
ISAC
integrated sensing and communication
ISTN
integrated space–terrestrial network
ITU
International Telecommunication Union
ITU-R
International Telecommunication Union Radiocommunication Sector
LAPS
low-altitude platform station
LEO
low Earth orbit
LoS
line-of-sight
LoRa
long range
LTE
long term evolution
MBS
macrocell base station
MC
multi-connectivity
MDP
Markov decision process
MEC
Mobile Edge Computing
MEO
medium Earth orbit
MIMO
multiple-input multiple-output
ML
machine learning
mMIMO
massive multiple-input multiple-output
MMSE
minimum mean square error
mMTC
massive machine type communication
mmWave
millimetre wave
NB-IoT
narrow band IoT
NOMA
non-orthogonal multiple access
NR
new radio
NTN
non-terrestrial network
OFDM
orthogonal frequency division multiplexing
OFDMA
orthogonal frequency division multiple access
PHY
physical
PNT
Positioning, Navigation, and Timing
PPO
proximal policy optimization
QoS
quality of service
RA
resource allocation
RAN
radio access network
RAT
radio access technology
RB
resource block
RF
radio frequency
RIS
reconfigurable intelligent surface
RL
reinforcement learning
RRC
Radio Resource Control
RSRP
reference signal received power
RTT
round-trip time
SAGIN
space-air-ground integrated network
SatCom
satellite communication
SBS
small base station
SIC
Successive Interference Cancellation
SINR
signal to interference plus noise ratio
SLNR
signal to leakage interference and noise ratio
SMBS
super macro base station
SNR
signal to noise ratio
SVM
support vector machine
TBS
terrestrial base station
THz
terahertz
TN
terrestrial network
UAV
unmanned aerial vehicle
UE
user equipment
UMTS
universal mobile telecommunications system
URLLC
ultra-reliable low-latency communication
VEC
Vehicular Edge Computing
vHetNet
vertical heterogeneous network
VR
virtual reality
VSAT
very small aperture terminal
VTOL
vertical take-off and landing
III HAPS Comparison with Terrestrial Networks and Other NTN Nodes
Global connectivity, traditionally supported by TNs and satellites, still faces significant gaps. Despite mobile broadband traffic doubling from 419 exabytes (EB) in 2019 to 913 EB in 2022, about 33% of the world’s population remains offline [105]. This digital divide highlights the need for innovative solutions. HAPS, operating in the stratosphere at 20–50 km, offer a promising approach. Acting independently or as intermediaries between TNs and other NTN systems such as LEO satellites and UAVs, HAPS provide wide-area coverage, low latency, and operational flexibility. These features enable them to bridge connectivity gaps while complementing existing networks. This section examines HAPS capabilities compared with TNs (Section III-A), LEO satellites (Section III-B), and UAVs (Section III-C), highlighting their unique advantages. Table II summarizes the distinct characteristics of HAPS relative to TNs and other NTN nodes in terms of altitude, performance, operational flexibility, and cost.
Table II: Comparison of HAPS with TNs, LEO satellites, and UAVs.
Feature TNs LEO satellites UAVs HAPS
Altitude Ground-based 500–2,000 km Up to 1.5 km 20–50 km
Coverage area Local to regional Global (constellations) Localized Regional to wide-area
Latency Low Moderate to high Low Low
Operational duration Continuous Years Hours to days Weeks to months
Deployment flexibility Limited by infrastructure Rigid (requires launch) Highly flexible Moderate (fixed stratospheric path)
Cost High for remote areas Very high Moderate to low Moderate
Power source Grid or renewable energy Solar or batteries Batteries Solar-powered
Payload capacity High Moderate Low Moderate to high
Adaptability Limited Low High Moderate to high
Emergency response Slow (requires rebuilding) Limited Rapid deployment Rapid deployment
III-A Comparison with Terrestrial Networks
TNs serve as the backbone of modern communication systems, providing high-capacity, low-latency connectivity in densely populated areas. However, they are constrained by geographical and infrastructural limitations. Rugged terrains, remote regions, and disaster-stricken areas often lack adequate terrestrial infrastructure, leaving significant portions of the population unconnected. Furthermore, the deployment and maintenance of TNs in such regions involve high costs, making them impractical for universal connectivity.
Recent studies on networking have also emphasized these limits. For instance, previous research [59] illustrates the challenge of extending traditional terrestrial infrastructure into rural areas, while efforts on community cellular networks [88] and virtual coverage [92] reveal that operational complexity, cost, and power availability are the main obstacles to large-scale terrestrial deployment in underserved areas.
HAPS overcome these limitations by offering wide-area coverage without reliance on extensive ground infrastructure. They can provide reliable LoS connectivity across urban, rural, and remote areas, bridging gaps where terrestrial solutions are infeasible. Their rapid deployment capability also makes them ideal for emergency scenarios, while operational costs in remote areas are often lower than constructing terrestrial networks.
III-B Comparison with LEO Satellites
LEO satellites provide essential global coverage, particularly in remote and oceanic regions. However, their high operational altitudes (500 to 2,000 kilometers) result in increased signal attenuation and latency. Moreover, their large coverage areas imply that the available bandwidth is shared among a large number of UEs, leading to reduced per-UE data rates.
These system-level limitations are further supported by recent research. In LEO systems, downlink opportunities, contact windows, and ground-station coordination dominate end-to-end latency, according to work on distributed downlink architectures and constellation-scale traffic scheduling [217]. Direct-to-cell designs [137] highlight fundamental per-user capacity and resource-sharing limitations when large footprints serve massive user populations.These constraints make LEO satellites less suited for high-bandwidth or latency-sensitive applications. In contrast, HAPS are positioned significantly closer to Earth, which reduces latency and enables better spatial reuse, resulting in higher bandwidth availability per UE. Additionally, HAPS offer greater operational flexibility, as they can be repositioned and upgraded without the complexity and cost of satellite launches.
III-C Comparison with UAVs
UAVs are highly versatile, providing localized and flexible connectivity for hot-spot operations and disaster scenarios. However, their limited flight duration due to battery constraints, restricted payload capacity, and frequent maintenance requirements hinder scalability for large-area or long-term deployments. Recent system-level investigations also show these limitations. The efficiency of UAVs for short-term, localized coverage is demonstrated by research on drone-based communication relays [144] and integrated rural connectivity platforms [216], while also highlighting endurance and sustainability constraints that preclude persistent wide-area service.
In contrast, HAPS offer extended operational durations, lasting weeks or months, enabled by solar power and autonomous control. Their quasi-stationary positioning allows them to function as stable, high-capacity communication hubs, effectively bridging the gap between the localized flexibility of UAVs and the global reach of satellite systems.
III-D Quantitative and Structured Comparison Across NTN Architectures
III-D1 Latency
LEO satellites experience non-negligible propagation and routing delays, on average around 25 ms [155, 71] of a round-trip time (RTT). For instance, a satellite operating at an altitude of approximately 550 km has a theoretical RTT propagation delay of about 3.6 ms; however, practical RTTs are significantly higher, ranging between 30 and 50 ms, due to routing, ground station access, and network management overheads. Notably, in our experimental evaluation [67], we revealed periodic high-magnitude RTT spikes of up to 140 ms, at regular intervals of approximately 13 s. This periodic behavior is indicative of constellation-specific control and resource management mechanisms, such as inter-beam or inter-satellite handovers, which require additional signaling for path reconfiguration and session re-initialization. By contrast, HAPS platforms operating in the stratosphere at altitudes of 18–25 km have nearly negligible propagation delays, and the resulting RTTs is typically comparable to that of terrestrial 5G systems, consistently below 10 ms. This makes HAPS particularly attractive for latency-sensitive services. Finally, UAV swarms operate at very low altitudes (e.g., a few hundred meters), and achieve sub-millisecond propagation delays. However, in practice, the end-to-end latency is often dominated by multi-hop processing and coordination overhead within the swarm’s mesh network [152, 143].
III-D2 CAPEX/OPEX
Global LEO constellations require extreme upfront capital investment. Public estimates for Amazon’s Project Kuiper and SpaceX’s Starlink indicate a total CAPEX of approximately 10 billion USD to reach full operational capability [168, 132]. According to [236], the average CAPEX per Starlink satellite is approximately 548,000 USD, for both satellite manufacturing and launch costs. Inter-satellite links (ISLs) in second-generation satellites introduce an additional cost of several hundred thousand USD per link. Regarding OPEX, costs are primarily dominated by ground stations, which require an average annual cost of approximately 215,000 USD on average, vs. around 20,000 USD for each satellite. Furthermore, the relatively short operational lifetime of LEO satellites (approximately five years) necessitates frequent satellite replacement, thereby increasing long-term costs. In comparison, the CAPEX of HAPS is significantly lower. For example, the manufacturing cost of a Airbus Zephyr platform is estimated at approximately 4 million USD [212], a number that has decreased significantly as stratospheric technology have matured. Unlike LEO satellites, HAPS platforms are typically designed to operate for around of 20 years, so that initial investments can be amortized over a much longer period. In terms of OPEX, costs are mainly due to maintenance and logistics, reaching up to 15% of the manufacturing cost annually. However, energy costs remain minimal as HAPS platforms utilize high-efficiency solar panels for daytime flight and onboard batteries for nighttime operations, resulting in an annual energy cost of only 30,000 USD per platform. Finally, the individual CAPEX of large civilian heavy-lift UAVs is relatively low, typically on the order of a few thousand USD, while OPEX is dominated by energy consumption and maintenance, in particular battery replacement and charging logistics.
III-D3 Scalability
While large-scale LEO satellite constellations are designed to provide ubiquitous coverage on Earth, the network capacity must be shared among a large number of UEs, which can saturate the available resources and complicate channel access [181]. In contrast, HAPS platforms only operate at a regional level, but provide high-capacity quasi-stationary connectivity via aggressive frequency reuse. Finally, UAV swarms provide extreme localized (though high-speed) coverage in hot-spot areas (e.g., for temporary events or emergency scenarios), even though they do not scale efficiently over large geographic regions.
III-D4 Operational risk
LEO constellations face both short- and long-term operational risks, including launch failures, orbital congestion, and the increasing threat of space debris and Kessler Syndrome effects. HAPS operations are primarily challenged by stratospheric wind dynamics and seasonal power constraints, particularly during the winter solstice at latitudes above approximately 50, where the shorter daylight duration may limit battery recharging capability for overnight flight. UAV swarms face the most stringent regulatory and legal constraints on the maximum flight altitude, airspace access, and spectrum usage, especially for Beyond Visual Line-of-Sight (BVLOS) flight operations over populated areas, which typically require case-by-case authorization and continuous human supervision. Moreover, UAVs are vulnerable to adverse weather conditions, such as strong winds and precipitation.
III-E Cross-Platform Energy Modeling for HAPS
A single energy-model framework that separates (i) solar energy harvesting, (ii) storage restrictions, and (iii) platform/payload power budgets is necessary for a consistent comparison of HAPS types (such as fixed-wing solar aircraft and aerostatic/airship systems). The reported energy efficiency and endurance findings are not directly comparable across research due to the lack of established measures and assumptions.
Solar harvesting is commonly represented as a function of solar irradiance $G(t)$ (W/m2), conversion efficiency $\eta_{\mathrm{PV}}$ (%), and effective photovoltaic area $A_{\mathrm{PV}}$ (m2). The harvested power can be expressed as
$P_{\mathrm{harv}}(t)=\eta_{\mathrm{PV}}\,A_{\mathrm{PV}}\,G(t)$ (1)
using temporally integrated daily gathered energy. To present harvested energy in kWh/day and facilitate cross-platform normalization, fixed-wing solar HAPS experiments explicitly parameterize these values under latitude- and season-dependent irradiance circumstances [29]. Similar irradiance-driven formulations are used in system-level HAPS energy studies, confirming their applicability as a platform-independent reference model [15].
Storage limits are characterized using the storage capacity $E_{\mathrm{sto}}$ (kWh), the usable fraction (depth-of-discharge and round-trip constraints), and the equivalent sustainment time under a defined load. The storage sustainment time is given by
$T_{\mathrm{sto}}=\frac{E_{\mathrm{use}}}{P_{\mathrm{tot}}}$ (2)
where $P_{\mathrm{tot}}$ denotes the total platform power demand. Studies on the viability and power budget of high-altitude airships show how storage sizing relates to propulsion or station-keeping requirements and payload loads. These works also provide guidance for consistent reporting of storage capacity and endurance across platform types [50, 46].
Payload power budgets are most comparable when decomposed into communications payload power $P_{\mathrm{pl}}$, propulsion or station-keeping power $P_{\mathrm{prop}}$, and auxiliary/avionics power $P_{\mathrm{aux}}$. The total platform power consumption can therefore be written as
$P_{\mathrm{tot}}=P_{\mathrm{pl}}+P_{\mathrm{prop}}+P_{\mathrm{aux}}$ (3)
This decomposition enables comparison between propulsion-dominant fixed-wing platforms and station-keeping-dominant aerostatic systems. In particular, the payload power share $\left(P_{\mathrm{pl}}/P_{\mathrm{tot}}\right)$ can be used as a normalized efficiency indicator [29, 46].
To ensure fair interpretation for TN–NTN integration research, these platform-level metrics should be reported alongside altitude, endurance targets, and irradiance assumptions (latitude and season). Furthermore, linking platform power budgets to network-level effects—such as reducing terrestrial densification requirements or supporting traffic spikes—strengthens the relevance of energy comparisons in integrated architectures [121]. Consequently, a transparent and repeatable basis for cross-platform evaluation can be achieved by presenting solar harvesting characteristics, storage capacity and sustainment, and complete power budget breakdowns within a unified comparison table.
III-F HAPS Market Overview
HAPS offer a cost-effective complement to terrestrial and satellite networks in NTN architectures, particularly for rural connectivity and emergency services [182]. This subsection addresses economic viability through cost breakdowns, sustainable business models, and market analysis of key providers. As mentioned above, the CAPEX/OPEX comparison between LEO satellites and HAPS is generally more favorable for HAPS, and the overall environmental burden is typically lower. Meanwhile, sustainable models for HAPS operators emphasize multi-layer integration with satellites and TN infrastructure, industry partnership, and green energy to achieve long-term viability [125]. Reliance on solar power supports UN Sustainable Development Goals 7 and 13, reducing carbon emissions and enabling eco-efficient operations in remote, disaster, and maritime scenarios. Collaborative models, such as those between Airbus, NTT DOCOMO, and SKY Perfect JSAT [166] or SoftBank and Sceye [199], focus on direct-to-device services via HAPS-NTN hybrids, targeting air/sea connectivity and disaster response. These ensure sustainability through spectrum sharing, regulatory alignment, and revenue from government/defense contracts (e.g., surveillance, environmental monitoring).
While the global HAPS market is clearly evolving and expanding across a broad range of applications, it remains nascent and fragmented. The limited availability of long-term historical data makes market projections highly sensitive to assumptions regarding adoption rates and regulatory developments. According to MarkNtel Advisors, the HAPS market is projected to grow from \$99 million in 2024 to \$240 million by 2030, driven by a diverse set of industrial applications [151]. Published in March 2025, this report focuses primarily on core HAPS platforms, excluding ancillary services, ground infrastructure, and broader ecosystem components from its scope. In contrast, Global Insight Services estimates a significantly larger market expansion—from \$1.7 billion in 2025 to \$3.3 billion by 2035—anticipating growth supported by advances in materials science, solar power technologies, and regulatory facilitation [79]. This assessment adopts a broader ecosystem perspective, encompassing not only the platforms themselves but also associated products, subsystems, and supporting equipment. Despite differences in scope and valuation methodology, both reports converge on the conclusion that the communications segment constitutes the dominant share of the HAPS market (approximately 35–45%), followed by surveillance and navigation applications, with the Asia-Pacific region emerging as a leading growth area.
IV HAPS Use Cases
The unique characteristics of HAPS enable a wide range of applications across telecommunications and beyond. In telecommunications, they primarily aim to extend coverage (Section IV-A) and enhance network efficiency (Section IV-B). Their rapid deployment and adaptability make them essential for bridging connectivity gaps in remote, underserved, or disaster-affected regions. Serving as mobile urban BSs or backhaul providers, HAPS improve capacity, reliability, and energy efficiency. Beyond communications, they support advanced applications (Section IV-C), including real-time IoT sensing, immersive augmented reality (AR)/ virtual reality (VR), precision control via URLLC (Section IV-D), and energy optimization (Section IV-E). By bringing processing closer to UEs, HAPS reduce latency and enable innovations across sectors such as agriculture, logistics, and disaster management. This section explores their roles in connectivity enhancement and advanced use cases shaping next-generation networks.
IV-A Connectivity Enhancement
IV-A1 Remote and Rural Area Coverage
The COVID-19 pandemic highlighted global Internet access disparities, especially in remote and underserved regions, emphasizing the need for scalable connectivity solutions [43]. HAPS can address these gaps by extending mobile coverage to areas lacking terrestrial infrastructure. Operating similarly to TNs, they can bridge the digital divide and deliver reliable global connectivity [21].
Collaboration between HAPS and LEO satellites was explored by Jia et al. [109], who formulated a mixed-integer nonlinear programming problem to enhance large-scale access and backhaul. The proposed restricted trilateral matching algorithm improved coordination, connectivity, and efficiency in remote regions. Zhao et al. [249] analyzed HAPS-LAPS integration using stochastic geometry, introducing a framework to assess platform density and altitude effects under backhaul constraints, offering design guidance for robust aerial networks in rural areas. Arum et al. [30] considered HAPS for extending rural coverage with low UE density through integration with TNs, emphasizing coexistence via dynamic resource management, interference mitigation, and techniques such as advanced antennas, radio environment maps, and device-to-device communication. Zhang, Kishk and Alouini [246] optimized resource allocation for heterogeneous services in rural HAPS networks using free space optical (FSO) backhaul. It proposed static and dynamic allocation schemes leveraging DRL, improving age of information, data rates, and overall network efficiency.
IV-A2 Natural Disaster Region Coverage
In the aftermath of natural disasters, when TNs collapse, HAPS can provide real-time aerial surveillance and rapid, reliable communication. Covering vast and remote regions, they ensure effective disaster response and emergency connectivity for first responders [75]. Yu et al. [238] examined UAV-based HAPS for disaster scenarios, highlighting their rapid deployment, flexible coverage, and communication recovery, while identifying challenges such as backhaul limitations, platform constraints, and communication capacity that require targeted solutions. Real-time monitoring using HAPS was demonstrated by Baraniello et al. [37], showing persistent, high-resolution imaging from altitudes around 20 km, enabling stable, continuous coverage of affected areas. In the study provided by Lähdekorpi et al. [129], HAPS restored services in a disrupted universal mobile telecommunications system (UMTS) network, achieving 70% throughput recovery when terrestrial BSs were unavailable, underscoring their effectiveness in disaster recovery operations.
IV-A3 Unexpected Traffic Events
The inherent mobility and on-demand deployment of HAPS make them well suited for managing dynamic communication demands, including large public events, sudden traffic surges, or routine daily traffic variations [116, 61]. By adapting to these fluctuations, HAPS alleviate congestion, enhance quality of service (QoS), and improve energy efficiency. Compared to small cell densification, which requires extensive planning, fixed infrastructure, and offers limited flexibility, HAPS provide a faster and more cost-effective solution that can be deployed in real time without physical constraints. A QoS-driven deployment framework was proposed by Dong et al. [61] to optimize HAPS operation based on coverage, capacity, and latency metrics, balancing efficiency and cost. In integrated networks where ground and aerial BSs share spectrum, interference remains a key issue. As noted by Wei et al. [227], advanced techniques such as dynamic resource allocation, spectrum sharing, and adaptive power control enable efficient HAPS operation and seamless coexistence with TNs, ensuring robust performance under variable traffic conditions. Jeon et al. [108] installed a reconfigurable intelligent surface (RIS) on a HAPS to provide extensive LoS and full-area coverage for UAV-BSs deployed in response to sudden urban traffic spikes. The combined optimization of RIS phases, array-partitioning algorithms, and positioning enhanced the energy efficiency of each UAV-BS.
IV-B Network Infrastructure and Management
IV-B1 Backhauling
Network backhaul is a fundamental component of communication infrastructure, responsible for high-capacity, low-latency data transmission between BSs and the core network. In dense urban areas, fiber optics remain the dominant solution, offering high throughput and minimal latency, whereas copper cables, limited in bandwidth, are largely obsolete for modern traffic demands. In contrast, rural and remote regions often rely on satellite or microwave links, as fiber deployment is economically and logistically prohibitive [56]. However, non-geostationary satellite systems require costly ground gateways with precise placement to maintain reliable connectivity [34]. Although wireless backhauling has been studied extensively [174], HAPS are emerging as a promising enabler for next-generation networks. Positioned in the stratosphere, they provide near-LoS connectivity with lower latency compared to satellites, supporting applications such as small cell backhauling, wireless sensor networks, and Earth observation [158]. In-band mmWave backhaul has been demonstrated by Taori and Sridharan [207], while multi-band approaches for enhanced capacity were discussed by Tezergil and Onur [211].
Recent studies proposed advanced architectures leveraging HAPS. Abbasi and Yanikomeroglu [13], introduced a cell-free scheme with HAPS as an aerial central processing unit (CPU) aggregating UAVs via terahertz (THz) links, where large antenna arrays mitigate path loss. Hybrid radio frequency (RF)/FSO solutions further enhanced flexibility: Dang, Le, Nguyen, and Pham [58] presented a cooperative hybrid ARQ protocol ensuring latency fairness under varying conditions, and Yahia, Erdogan, and Kurt [233] explored hybrid LEO-to-HAPS FSO and HAPS-to-ground RF links. Collectively, these advancements demonstrate the potential of HAPS-enabled backhaul to deliver scalable, low-latency connectivity across heterogeneous environments.
IV-B2 Handover Management
Frequent handovers and associated failures represent a major challenge in satellite communication systems, particularly for fast-moving non-geostationary satellites that require continuous connectivity updates. HAPS mitigate these issues in two key roles: (i) as intermediaries between satellites and UEs, handling handover requests on behalf of satellites, and (ii) as standalone communication managers, directly managing UE connections. In both cases, HAPS improve network stability and reduce signaling overhead. Leveraging HAPS as intermediaries was shown by Cianca et al. [49] to decouple satellite mobility from ground-based handover procedures, simplifying management and enhancing continuity. Dynamic LEO-HAPS handover strategies proposed by Li et al. [131] achieved lower delays, signaling costs, and drop rates than traditional 5G methods. Prior to that, Mohammed, Mehmood, Pavlidou, and Mohorcic [158] explored FSO-based LEO-HAPS links, demonstrating high-speed, low-latency handovers even under high satellite mobility. When operating as quasi-stationary communication nodes, HAPS significantly reduce handover frequency for terrestrial UEs, providing seamless service in high-mobility scenarios (e.g., vehicles, aircraft). Their stable positioning eliminates inter-platform handovers common in LEO constellations. Combining beamforming for HAPS-to-ground and FSO backhaul to LEO/MEO satellites, Srinivasan et al. [202] demonstrated efficient handover management in the airliner system even in dynamic environments. Moreover, optical HAPS-satellite links, as analyzed be Perlot et al. [173], can further enhance reliability by ensuring robust backhaul connectivity, thereby minimizing disruptions in HAPS-to-UE sessions.
IV-C IoT and Emerging Applications
IV-C1 IoT sensing
Regarding this, HAPS can significantly enhance IoT sensing and monitoring by providing wide, stable coverage and integrating edge computing for low-latency data processing close to the source. Their dynamic positioning supports optimized sensor deployment and efficient data acquisition, making them ideal for remote and critical environments.
A reliable IoT architecture leveraging HAPS was proposed by Sibiya and Olugbara [196], integrating narrow band IoT (NB-IoT) and long range (LoRa) to deliver scalable, cost-efficient connectivity across underserved regions. Operating at 17–35 km, HAPS achieved coverage up to 600 km, surpassing terrestrial solutions in scalability and redundancy through overlapping cells.
Andreadis, Giambene, and Zambon [27] introduced a joint HAPS-UAV framework for IoT applications—especially disaster scenarios. Combining umbrella coverage from HAPS and agile UAVs, the system ensured continuous service with sub-millisecond latency (0.07–0.4 ms for HAPS), while NB-IoT/LoRa access schemes minimized collisions under dense deployments. Complementary LoRa-based monitoring was examined by Giambene and Korre [76], which optimized parameters such as altitude, transmission power, and device density under shadowing and collision constraints, confirming the viability of HAPS for large-scale environmental IoT. To enhance HAPS network capacity, Xiao, Li, and Zhao [231] introduced a 5G IoT architecture with a uniform rectangular array antenna and a Long Short-Term Memory (LSTM) model for direction-of-arrival prediction. The approach improved beamforming precision and link stability, indirectly benefiting IoT reliability. Joint LEO-HAPS IoT integration was explored by Ei et al. in [63], where HAPS relayed data to ground centers via LEO satellites, optimizing throughput and power consumption under QoS and storage constraints. Similarly, Monzon and Alvarez [160] presented a smart rural framework combining 5G, cloud, IoT, and HAPS, enabling automated remote data access and processing.
IV-C2 Mobile Edge Computing and Fast UE Movement
MEC enhances network performance by placing computing resources near UEs, reducing latency for real-time services. However, maintaining seamless connectivity for fast-moving UEs requires advanced handover and predictive mechanisms to sustain service continuity. Traspadini et al. [214] analyzed the integration of HAPS and UAVs within Vehicular Edge Computing (VEC) systems using queuing theory to optimize task offloading in dynamic vehicular environments. The authors showed that HAPS, with quasi-stationary operation, broad coverage, and strong LoS links, outperform UAVs by supporting larger computing payloads and offering lower latency, making them ideal for large-scale, high-mobility VEC/MEC scenarios. Later, Seid et al. [188] proposed to use aerial MEC servers to support ground user equipments and emergency communications. The authors introduced a hierarchical deep federated learning resource allocation (RA) scheme for intelligent aerial-enabled smart city cyber-physical system, aiming at optimizing energy consumption while preserving task offloading privacy and maintaining QoS in dynamic environments.Although prior works such as [204, 225] optimized UAV-enabled MEC for task completion and energy efficiency, these approaches can be extended to HAPS, which offer superior payload capacity and operational stability. Yet, research on HAPS-based MEC remains limited, representing a promising avenue for future exploration.
IV-C3 Aerial Data Centers and AR/VR Support
Beyond edge computing, HAPS can function as aerial data centers, delivering centralized cloud services for data storage, processing, and distribution from the stratosphere. With wide coverage, high capacity, and low-latency connectivity, they support resource-intensive applications such as AR and VR [20], enabling immersive experiences even in remote or underserved regions [126]. Traspadini et al. [213] presented a HAPS-enabled VEC system for rural areas, showing performance gains through optimized trade-offs between computational load and latency. A cloud-integrated HAPS system proposed in [153] achieved up to 14% energy savings compared to terrestrial data centers, reducing offloading outages and delays in distributed server networks while supporting services such as IoT, vehicular networks, and social media. Furthermore, [63] introduced a collaborative LEO-HAPS framework in which HAPS acted as relays, forwarding IoT data to ground processing centers with optimized data rates, power efficiency, and QoS compliance.
IV-D Advanced Control and Special Applications
IV-D1 URLLC Support and Remote Driving Control
Operating at high altitudes but closer to the Earth than traditional satellites, HAPS offer a unique combination of wide-area coverage, low-latency communication, and substantial payload capacity. These features enable the integration of advanced antenna systems and onboard processing units, making HAPS well-suited for supporting URLLC services such as remote driving control. Salehi et al. [187] analyzed a heterogeneous network integrating HAPS, terrestrial BSs, and autonomous vehicles, showing that multi-connectivity (MC) is essential to meet stringent URLLC requirements. HAPS acted as auxiliary links sustaining data rates above 700 kbps with high reliability under dynamic conditions. Building on this, the authors in [186] extended the model with LEO satellites, optimizing MC path selection to satisfy QoS constraints with minimal link usage. Results revealed that LoS interference is a key performance-limiting factor in radio access links, which was further addressed by Manzoor, Ozger, Schupke, and Cavdar [150] through refined MC strategies for enhanced reliability and efficiency. Beyond URLLC, HAPS can improve Global Navigation Satellite System (GNSS) performance due to their quasi-stationary position and strong signal power. Zheng et al. [250] demonstrated a HAPS-aided GPS architecture that improves 3D positioning accuracy, enhancing navigation and control functions. This capability is particularly valuable for remote or underserved regions—such as mountainous, coastal, or oceanic areas—where HAPS can act as super macro base stations (HAPS-SMBS) to backhaul isolated nodes and extend coverage [116].
IV-D2 UAV Control and Aerial Highways
The cellular community has made significant efforts to enable reliable UAV connectivity and control by leveraging existing terrestrial networks while minimizing interference with ground UEs [3, 86]. Current approaches, such as time- and frequency-domain resource separation, offer partial improvements but lack scalability to accommodate the projected growth of aerial traffic. To overcome these limitations, research has explored enhanced infrastructure densification, dedicated aerial communication systems [72, 54, 161, 123, 172], and LEO-based augmentation [74].
A promising solution involved establishing aerial highways—structured UAV corridors that enable regulated, safe, and efficient airspace management [44, 40]. To support seamless communication along these corridors, recent works have optimized cellular network configurations and resource allocation strategies tailored for aerial traffic [38, 39, 145, 47, 197], while theoretical frameworks have been proposed to design structured airspace operations [118, 117].
With their quasi-stationary position, large coverage footprint, and low-latency links, HAPS are well-suited to serve as backbone infrastructure for both open-air and corridor-based UAV networks. As discussed in Section IV-B, HAPS significantly reduce handover frequency and ensure continuous connectivity for high-mobility UAVs, while avoiding the inter-platform coordination challenges inherent to LEO systems. This combination of coverage, stability, and operational efficiency positions HAPS as a key enabler for scalable and reliable UAV control.
IV-E Energy and Sustainability
Energy efficiency is a critical consideration for future integrated TNs and NTNs. Traditional TNs rely on continuous operation of macro BSs to provide ubiquitous coverage, leading to substantial energy consumption even during low-traffic periods. This ”zero bit, non-zero watt” phenomenon reflects the inefficiency of maintaining coverage when demand is minimal. To address this issue, a promising solution involves offloading traffic from terrestrial macro BSs to NTN-powered hypercells, such as those supported by LEO satellites, HAPS, or tethered UAVs. During nighttime or periods of low demand, this approach allows power-intensive terrestrial macro BSs to enter low-power states or shut down entirely, significantly reducing energy consumption. In this context, LEO satellites are well-suited for offloading outdoor UEs due to their extensive coverage and ability to serve low-capacity traffic efficiently during nighttime. In contrast, HAPS, with their superior link budget and better signal penetration, are better suited for offloading indoor UEs, especially in urban environments where high spatial reuse is essential. Song et al. [201] developed a heuristic algorithm to optimize this offloading process. By prioritizing the least-utilized BS for offloading, and leveraging HAPS’s self-sufficient energy capabilities, the proposed approach minimized hourly energy consumption across the network. Simulation results indicated that this method could reduce energy consumption by up to 29% over a week, with savings reaching 41% during nighttime hours. These results highlight the complementary roles of HAPS in advancing sustainable and energy-efficient connectivity within integrated TNs and NTNs.
V HAPS Architectures
The integration of TNs and NTNs is critical to addressing the diverse connectivity challenges identified in Section IV, such as extending coverage to remote areas, enabling robust disaster communication, managing unexpected traffic surges, supporting IoT applications, and facilitating advanced services such as mobile edge computing and AR/VR. To this end, NTN platforms must efficiently transmit/receive communication signals between UEs, gateways, BS, and core networks. While the recent progression of NTN standardization is focusing satellites [128], we expect, by serving as dynamic platforms and connectors, HAPS can play a key role in bridging NTN platforms and creating a resilient, high-capacity, global communication infrastructure.
V-A NTN Architecture
NTNs refers to the networks which are using satellites or uncrewed aerial platforms in various constellations as relay nodes or BSs. In general, NTNs (see Fig. 2) include several key components, each playing a specific role in ensuring global, reliable communication. These components can be categorized into three main layers, namely space, aerial, and ground layers [96]. The space layer comprises LEO, MEO, GEO and even highly elliptical orbiting satellites, while the aerial layer consists of HAPS and low-altitude UAVs. Satellites, HAPS, and UAVs in this case are considered as network platforms, which carry the node able to facilitate communications. However, in certain use cases (for example, Earth observation satellites and HAPS or delivery drones), they may serve as network terminals, i.e., connectivity consumers, providing no communication functionality. The ground layer typically includes infrastructure that interacts with the space and aerial layers, such as gateway stations, core networks, and data centers. Considering TN-NTN integration, the different layers can be combined together to provide better coverage, capacity, resilience, and flexibility compared to standalone deployments [219]. The NTN architecture, as defined by the 3GPP in [2], supports two fundamental payload designs: transparent (or bent-pipe) and regenerative. In the transparent approach, the NTN platform carries communication equipment that functions as a simple relay, forwarding signals without modification. In contrast, the regenerative architecture enables onboard signal processing at the NTN platform, allowing the digital processing of the — from packet data convergence protocol (PDCP) to PHY layer — to be performed directly within the platform. These architectural choices impact key factors such as spectral efficiency, rate, latency, and overall system complexity, shaping how NTNs integrate with TNs. Building on this foundation, we can distinguish three main network components within the NTN architecture: (i) platforms, which carry communication equipment and implement either a transparent or regenerative payload to relay or generate communication signals; (ii) terminals, including direct UEs or specialized devices designed to interact with NTN platforms; (iii) gateways, which connect NTNs, e.g., satellites and HAPS, to core networks, facilitating data flow between TNs and NTNs.
In Section V-B, we will describe the different NTN components, including platforms, terminals, gateways, and their interactions, focusing on HAPS. At the same time, the NTN architecture incorporates the following types of links [4]: service links, feeder links, backhaul links, and inter-platform links, which we discuss is subsection V-C, as well as the promising frequency bands for NTNs based on the link types, with considerations related to range, throughput, latency, and interference sensitivity. In Section V-D, we will compare 3GPP bent-pipe and regenerative architectures, focusing on the role of HAPS. Finally, in Section V-E, we will discuss ongoing and already closed HAPS-related projects, challenges they are facing and lessons which must be learned from the progress done so far.
V-B HAPS Platforms and Terminals
Table III: Comparison of HAPS platforms.
Characteristic Aerostatic Aerodynamic Hybrid
Lift mechanism Buoyant force (helium/hot air) Aerodynamic lift (fixed/rotary wings) Combination of buoyant and aerodynamic forces
Mobility Low when operating stationary or drifting slowly; high with favorable wind patterns High, regardless of wind direction; can reposition dynamically Moderate (repositioning capability)
Payload capacity High Low Moderate (payload/mobility trade-off)
Energy efficiency High (minimal energy to stay aloft) Low (continuous energy for flight) Moderate (depends on lift mechanics)
Weather resilience Sensitive to wind and turbulence More resistant to adverse weather, better operational stability Balanced
Deployment flexibility Requires ground infrastructure (hangars, airfields) in close proximity to the serving region Can take off and land outside serving region (deployment flexibility) Moderate (limited by mobility)
Operational duration Long and persistent (weeks/months) Variable (days/months), and may be solar-powered to extend flight life Potentially long, depending on design and hybrid lift use
Development cost Low, due to simpler design and manufacturing High, requiring advanced materials and engineering Highest, due to the complexity of integrating multiple lift systems
Notable example Loon Balloons [229] SoftBank’s “Sunglider” [198] Tactical HAPS by CIRA [104]
Figure 2: General NTN representation.
Now, we shift our focus to the aerial layer, with a particular emphasis on HAPS, the subject of this survey. We distinguish between NTN platforms, terminals, and gateways.
V-B1 Platforms
HAPS can be classified depending on the physical principle providing the lifting force: (i) aerostatic platforms, which rely on buoyant force to stay aloft, and include airships and balloons; (ii) aerodynamic platforms, which generate lift through fixed or rotary wings to maintain flight; (iii) hybrid platforms, which rely on a combination of aerostatic and aerodynamic effects at different stages of flight, such as the tactical HAPS by CIRA [104]. All these types of HAPS have their own benefits and drawbacks, which we discuss in the following and are summarized in Table III.
Aerostatic platforms
Aerostatic platforms are less expensive to manufacture, launch, and operate compared to aerodynamic and hybrid systems. While prototyping and refining their products, the Loon project team was able to fill and launch a balloon into the stratosphere every 30 minutes with their designed and custom-built auto-launchers and keep it in the air for more than 300 days [229]. Thus, by hovering over a specific region, aerostatic platforms can provide consistent coverage and are well-suited for persistent surveillance and localized communication needs. Since this platforms do not require forward speed to maintain flight, they only need power to hold or adjust their position. This results in potentially better power efficiency, and allows for heavier payloads compared to aerodynamic platforms. Furthermore, their large volume capacity enables them to carry more complex communication payloads, making aerostatic platforms particularly suitable for both short-term and long-term missions over specific regions. However, these platforms are more sensitive to atmospheric conditions, which may impact their operational stability and reliability. Additionally, launching and landing aerostatic platforms require specific facilities such as hangars and airfields located near the target region, which may limit their flexibility in deployment. However, the Loon project team demonstrated a method to navigate the balloons using dominant winds at different altitudes. This approach requires advanced predictive models to identify favorable wind patterns and intelligent algorithms to determine the most efficient flight paths through varying wind layers [229]. This allows the launch facilities to be located away from the target region, but makes deployment reliant on wind patterns.
Table IV: Comparison of NTN terminals.
Characteristic Portable consumer terminals Vehicular terminals Fixed UE terminals
Mobility Medium (handheld/wearable devices) High (cars, trains, aircraft, UAVs) Low (stationary gateways/sensors)
Power availability Limited (battery-powered) Moderate to high (power sources) High if fixed power supply available; limited for low-power IoT sensors
Antenna Small, integrated, limited gain Large, more powerful (e.g., VSAT, phased arrays) High-gain dish or directional antennas for long-range communication; limited gain for low-power IoT sensors
Coverage
Handover requirements
Seamless transition between TNs and NTNs for uninterrupted connectivity Multi-band, multi-standard support for high-speed mobility across regions Persistent connectivity with high-capacity backhaul support; limited data rates for low-power IoT sensors
Use cases Wearables, first responder devices, personal communication (phones) Connected vehicles, public transport, autonomous robots, UAVs Broadband Internet, industrial IoT, remote sensors
Network requirements Low latency for applications like gaming and video streaming Low latency for safety-critical applications, high throughput for real-time data High throughput for broadband applications, low power for IoT sensors
Deployment constraints Consumer-driven must be cost-effective and compact Requires vehicle integration and external antennas Fixed installation with higher infrastructure requirements
Example devices Smartphones, tablets, smartwatches, disaster relief radios Connected cars, aircraft communication systems, surveillance drones Satellite broadband dishes, industrial IoT gateways, environmental sensors
Aerodynamic platforms
Aerodynamic platforms offer significant advantages given their high flexibility, allowing them to move easily between regions in response to communication demands or environmental changes. This adaptability reduces the need for multiple airfields and enhances their capability for precise positioning through specific flight patterns. Modern aerodynamic designs are generally more robust against adverse weather conditions than aerostatic platforms, increasing their reliability in various operational scenarios. Additionally, many modern aerodynamic platforms, such as SoftBank’s “Sunglider” [198], can be entirely solar-powered, enabling them to remain operational for extended periods, ranging from weeks to months. However, these platforms come with some limitations. Their payload capacity was generally smaller compared to large aerostatic platforms, which can affect the scale and complexity of their communication systems. Moreover, the development and manufacturing costs of aerodynamic platforms are higher, reflecting their advanced engineering and specialized materials.
Hybrid platforms
Hybrid designs aim to combine the strengths of both aerostatic and aerodynamic platforms, offering a balance between payload capacity, stability, and mobility. They can carry significant payloads like aerostatic platforms, while also benefiting from the mobility and weather resilience typically associated with aerodynamic platforms. This combination makes hybrid solutions suitable for a wide range of applications, including scenarios requiring both high payload capacity and the ability to reposition dynamically. However, these advantages come with notable challenges. The complexity of hybrid designs makes them costly to develop and manufacture. Sophisticated control systems are required to manage the dual mechanisms for lift and propulsion, adding to the operational complexity. Moreover, while hybrid platforms mitigate some disadvantages of pure aerostatic or aerodynamic designs, they often inherit certain drawbacks, such as higher sensitivity to adverse weather compared to fully aerodynamic platforms and increased infrastructure requirements compared to simpler aerostatic systems.
V-B2 Terminals
Platforms ultimately aim at delivering connectivity to network terminals, a.k.a, UEs. Given the broad spectrum of application scenarios outlined in Section IV, it is evident that UEs extends far beyond conventional smartphones, and also includes wearable devices, IoT sensors, robots, and UAVs. Terminals can be classified in terms of features like form factor, operational environment, and communication capabilities. In the following paragraphs, we will describe the main groups of terminals, namely portable consumer terminals, vehicular terminals, and fixed terminals, as well as their features, and specific requirements, as summarized in Table IV.
Portable consumer terminals
Probably the largest group of UEs consists of portable consumer terminals, including smartphones, smartwatches, tablets, and laptops. These devices are designed for everyday use across urban, suburban, and rural areas, and require seamless handover between TNs and NTNs to ensure uninterrupted connectivity. The diverse range of capacity-demanding and latency-sensitive applications, such as gaming and video streaming, necessitates support for multiple frequency bands, compatibility with existing standards (e.g., 5G), and NTN-specific adaptations for satellite and HAPS links. Conventional mobile devices are further complemented by wearable and implantable terminals, such as fitness trackers, health monitors, and medical implants. These devices are typically ultra-portable, low-power, and designed for continuous operation. They often require only short-distance, low-data-rate communication with a UE’s hub device, such as a smartphone, relying on the ubiquitous coverage provided by the hub. Another significant group within portable terminals includes communication devices for first responders, disaster relief teams, and military personnel. These devices are specifically designed to operate in extreme conditions, providing reliable connectivity in areas where terrestrial infrastructure is damaged or nonexistent. They often require enhanced security and encryption features for sensitive communications, as well as high-capacity links to enable rapid data exchange in critical scenarios.
Vehicular terminals
Another distinct category of mobile devices comprises vehicular terminals, installed in cars, buses, trains, ships, and aircraft. These terminals operate under challenging conditions due to high mobility and diverse environments, including remote and maritime regions. Unlike handheld devices, they can accommodate more efficient power sources and advanced antennas, such as VSAT parabolic reflectors or phased array antennas. Typically, very small aperture terminal (VSAT) systems feature parabolic dishes ranging from 0.75 m to 2.4 m in diameter, with some larger models, providing high-gain, narrow-beamwidth links ideal for point-to-point communications. The antenna size depends on required gain, operating frequency, and environmental factors. In contrast, phased arrays, such as those employed in Starlink or BGAN systems, are more compact and rely on electronic beamforming through controlled phase shifts across multiple radiating elements. This enables rapid, precise beam steering without mechanical movement, offering enhanced flexibility and robustness under vibration or harsh conditions. Modern designs integrate features like auto-acquisition, self-alignment, and automatic aiming, simplifying deployment and ensuring reliability during high-speed operations. A related subset of vehicular devices includes unmanned vehicles, such as UAVs, self-driving cars, autonomous robots, and surveillance balloons. These systems combine high mobility with operational versatility across urban and remote areas. They require reliable, low-latency command-and-control links and high-capacity channels for real-time data transmission (e.g., video or telemetry). To maintain continuous connectivity beyond terrestrial coverage, NTN support—via platforms like LEO satellites or HAPS—is essential for ensuring global reach and service continuity.
Fixed UE terminals
Despite the widespread use of mobile UEs, fixed UEs remains an essential category of network terminals. These devices are primarily designed for stationary broadband access in rural and remote areas. Their antenna systems are similar to those used in vehicular VSAT terminals. Unlike mobile devices, however, they are not constrained by strict form-factor limitations, allowing for larger, more efficient antenna designs. Such equipment allows high throughput for broadband applications, the ability to support multiple simultaneous connections, and resilience to adverse weather conditions, particularly for satellite-based links.
A distinctly different subset of fixed terminals comprises IoT devices, including sensors, indoor consumer smart household appliances and industrial equipment. While industrial equipment and household appliances can be connected through industrial gateways or indoor access points, remote sensors rely on their own connectivity solutions. These sensors are typically compact, low-power, and deployed in large numbers. They require massive machine type communication (mMTC) connectivity, often in remote or hard-to-reach locations, and energy-efficient communication protocols to maximize battery life. This diverse range of UE requirements in integrated TN-NTN systems highlights the critical role of platform adaptability and technology integration. Meeting these diverse needs necessitates a collaborative ecosystem of terrestrial, satellite platforms, and HAPS, each uniquely suited to specific terminal categories and operating conditions. Specifically, the TN provides the best capacity and latency for densely populated areas, but is constrained in coverage and mobility support in remote regions, while satellites are indispensable for remote and hard-to-reach areas, but suffer from higher latency and lower rates. At the same time, HAPS serves as a versatile middle ground, offering good coverage, capacity, latency, and mobility, especially in semi-remote areas or regions requiring temporary connectivity.
V-B3 Gateways
Aside from NTN platforms, several essential components constitute the network infrastructure, including gateways and associated network elements that ensure connectivity, reliability, and efficient communication [220]. Gateways, often referred to as ground stations, serve as critical interfaces between TNs and NTNs. Strategically located to optimize coverage and capacity and minimize latency, gateways are typically unconstrained in terms of form factor and power consumption. They are equipped with a large antenna structure designed for higher frequency bands, and are capable of maintaining efficient uplink connections with NTN platforms. Depending on the overall architecture, gateways may perform functions such as signal processing, routing, and resource management. They also aggregate traffic, and facilitate integration of NTN components with the terrestrial core network.
V-C NTN Links and Frequency Bands
Table V: Comparison of frequency bands for NTN, in terms of range, throughput, latency, and interference sensitivity.
Band Typical range Peak throughput Latency profile Interference sensitivity
Sub-7 GHz (410 - 7125 MHz) Hundreds of km Low (<<1 Gbps) Low Low (robust propagation)
cm-Wave (7.125 - 24.25 GHz) Tens–hundreds of km Medium (1–10 Gbps) Medium Moderate (rain fade in Ku-band)
mm-Wave (24.25 - 52.6 GHz), FR2-2 (52,6 - 71 GHz) <<10 km High (10–100 Gbps) Medium High (severe rain attenuation)
Sub-THz (100 - 300 GHz) <<5 km Ultra-high (100+ Gbps) Low Very high (strong atmospheric loss)
THz (0.1 - 10 THz) <<1 km Tbps-class Very low Extreme (highly fragile links)
FSO (3 THz - 30 PHz) Up to hundreds of km Ultra-high (Tbps) Very low Weather-dependent (fog/clouds)
To establish reliable NTN communication, various types of links facilitate data exchange between different network components, namely feeder, service, inter-platform, and backhaul links. These links differ in terms of function, propagation characteristics, and performance requirements, influencing the choice of frequency bands and transmission technologies.
  • Service links provide connectivity between NTN platforms and UEs. These links must accommodate a diverse types of devices, ranging from stationary to highly mobile terminals, as described in Section V-B, posing challenges such as signal power constraints, Doppler effects, and interference management.
  • Feeder links connect NTN platforms directly to gateways, typically operating at high frequencies to enable high-capacity transmission. To ensure reliability, these links incorporate advanced error correction and signal processing to mitigate atmospheric disruptions.
  • Inter-platform links facilitate direct communication between NTN platforms, either within the same layer (e.g., satellite-to-satellite) or across different layers (e.g., satellite-to-HAPS or HAPS-to-UAV). By supporting direct routing and handover, these links reduce dependency on terrestrial infrastructure.
  • Backhaul links interconnect NTN gateways with core network facilities, typically utilizing fiber-optic infrastructure to ensure high-speed, low-latency data transport.
The performance of these links is heavily influenced by the choice of operating frequency bands. Each frequency range offers distinct propagation characteristics, affecting factors such as coverage, capacity, and resilience to atmospheric conditions. The introduction of 5G marked a significant shift in spectrum allocation, expanding the use of high-frequency bands such as in the mmWave spectrum to enhance network capacity. In parallel, NTNs operates in highly heterogeneous environments, requiring careful spectrum planning to address diverse communication needs. As 6G and future TN-NTN integration evolves, the selection of appropriate frequency bands will be critical in optimizing system performance.
Notably, in Release 17 the 3GPP formally integrated NTN into the 5G NR framework, defining use cases within the sub-7 GHz spectrum, particularly for massive access and narrowband IoT applications. Release 18 further expanded NTN support, introducing enhanced 5G NR operations for handheld terminals in the 10 GHz and above range. These updates focused on improved spectrum management and coexistence strategies to ensure seamless integration between TNs and NTNs [4, 2, 1].
The following subsections examine the key frequency bands currently used in NTN communications and those with potential for future adoption depending on the type of link, as summarized in Table V.
V-C1 Sub-7 GHz
Also known as 3GPP frequency range (FR) 1, this well-established frequency range, spanning from 410 MHz to 7125 MHz [8], has been widely used in long term evolution (LTE) and NR networks. FR1 is valued for its excellent propagation characteristics, resilience to atmospheric conditions, and an extensive ecosystem of mature telecommunication equipment. An important advantage of this band is the possibility to establish direct satellite-to-UE links, as it was recently demonstrated by companies like AST SpaceMobile, Starlink and Lynk Global. However, its limited bandwidth and high congestion restrict the achievable data rates, making it less suitable for high-capacity applications.
V-C2 mmWave
Referred to 3GPP FR2-1, mmWave spans frequencies from 24.25 GHz to 52.6 GHz while 3GPP FR2-2 designated for frequencies from 52.6 GHz to 71 GHz [8].Given its large bandwidth and short orthogonal frequency division multiplexing (OFDM) symbols, this spectrum enables ultra-high data rates and extremely low latency, making it ideal for bandwidth-intensive applications such as enhanced mobile broadband (eMBB). mmWave also offers fiber-like speeds in fronthaul and backhaul communications, supporting network densification through small cells and high-directivity antennas.
Despite its advantages, mmWave has seen limited adoption in TNs due to its short range and high infrastructure costs. The severe free-space path loss, combined with high susceptibility to blockage from buildings, foliage, and human movement, requires a dense network of BSs, making large-scale deployment expensive. While small cell and advanced beamforming techniques based on MIMO helps to mitigate these limitations, deployment remains challenging, particularly in rural and remote areas [45]. The use of mmWave in NTNs for broadband connectivity has recently gained significant research attention. The studies [77, 78] explored NTN applications in 6G, evaluating aerial- and space-based mmWave links to provide connectivity for ground terminals under various network configurations. However, expanding NTN into mmWave bands requires effective spectrum-sharing techniques to balance isolation, licensing costs, and coexistence with existing TNs. Moreover, large-scale propagation models must be re-evaluated, considering factors such as path loss, penetration loss (also through the atmosphere), and shadow fading, which significantly impact NTN protocol design [178, 190].
V-C3 cmWave
Also known as 3GPP FR3, cmWave spans 7.125 GHz to 24.25 GHz, bridging the gap between sub-7 GHz and mmWave bands [5]. Unlike mmWave band, which suffer from severe propagation characteristics, FR3 offers improved coverage and efficiency, with channel properties closer to sub-7 GHz in the sub-12 GHz region than to mmWave [52, 115, 154, 193]. Propagation studies [99] indicate that FR3 communication at 10 GHz experiences approximately 9.5 dB higher path loss in LoS, and 9.7 dB in non-LoS, compared to 3.5 GHz; yet these losses remain manageable with advanced beamforming. This band also exhibits lower penetration loss than at mmWaves, making it a strong candidate for outdoor-to-indoor coverage [115]. Additionally, rain attenuation is negligible below 10 GHz, and reduced delay spread improves signal quality, while angular spreads remain consistent with sub-6 GHz deployments. A key challenge in FR3 networks is the coexistence with satellite services, particularly in the 12 GHz band, which is heavily utilized by LEO constellations such as Starlink [52]. Interference mitigation techniques, including advanced beamforming and ephemeris-based coordination, have been proposed to ensure harmonious spectrum sharing between terrestrial and satellite networks [115]. Dynamic spectrum access mechanisms and interference nulling strategies are also being explored to optimize spectral efficiency while maintaining reliable satellite communications [114].
V-C4 Beyond-100 GHz: sub-THz and THz
The beyond-100 GHz region is covering sub-THz typically defined between 100 and 300 GHz [31, 48] and even wider THz range (0.1–10 THz).It offers extreme bandwidth for immersive communications and sensing [184, 110]. Measurements at 140 GHz report 2.3 Gbps over 120 m, highlighting the rate–range trade-off and equalization needs [159]. Simulation studies also show that integrated sensing and communication (ISAC) in these bands can outperform 5G-spectrum baselines [119]. For NTN topologies, sub-THz has been proposed for backhaul between UAV-BSs and CPUs, with HAPS acting as the CPU to preserve LoS [14]. Feasibility analyses indicate that THz feeder/backhaul links around 10 Gbps are possible in light rainfall given careful link budgets and signal processing [206] Reconfigurable intelligent surfaces (RISs) can further alleviate blockage [26]. Despite this potential, beyond-100 GHz faces severe constraints for NTNs: high free-space path loss, strong water-vapor absorption, weak diffraction, and hardware limits (low output power, thermal stability) restrict range and robustness [33, 200]. Consequently, sub-THz/THz should be viewed as complementary, beyond-5G options suited mainly to short-range, LoS inter-platform or backhaul links under controlled conditions—not for handheld service or long-range feeder links. Practical deployment will require multi-link coordination (e.g., fallbacks to FR2/FR3 or FSO), highly directive arrays with null steering, site/spatial diversity, and aggressive modulation/coding adaptation [184, 110].
V-C5 FSO
FSO frequency region covering an ultra-wide range of spectrum, roughly, between 3 THz and 30 PHz, which include infrared, visible light, and ultraviolet categories and basically overlap with THz region. It is worth noting, however, that unlike RF bands, no single bounded range is allocated for FSO by regulations, and discrepancies in the reported values may therefore occur. FSO systems offers a broad range of indoor and outdoor applications including cyber security, surveillance and monitoring, inter-building communication, high-capacity backhauling, etc. [106].FSO technology is a valuable approach for future wireless networks, offering cost-effectiveness, easy deployment, high bandwidth, and strong security. Samir et al. [17] provided an in-depth analysis of FSO systems, highlighting their role in next-generation optical communication. However, FSO propagation is challenged by atmospheric turbulence and alignment issues, necessitating mitigation strategies to maintain link stability. Shang et al. [194] evaluated a ground-to-HAPS-to-RIS-to-UE communication system, where FSO is used for the ground-to-HAPS link, while RF communication is established for HAPS-to-RIS-to-UE transmissions. Key performance metrics, such as ergodic capacity, average bit error rate (BER), and outage probability, are analyzed using the bivariate Fox-H function. Results indicate that pointing errors and shadowing significantly affect system performance, with larger zenith angles and severe shadowing leading to degradation. Heterodyne detection improves efficiency by enhancing capacity and reducing BER and outage probability.
Based on the above discussion, it is important to recognize that power limitations, operational altitude, and environmental factors play a crucial role in HAPS communication. Selecting the appropriate frequency band for each link requires balancing multiple considerations:
  • Power constraints: Higher frequency bands typically require more power for transmission, which can be challenging given the strict power budgets of HAPS.
  • Weather sensitivity: Atmospheric conditions, e.g., rain and turbulence, impact high-frequency bands more significantly than lower ones, affecting link reliability and QoS.
  • Coverage vs. capacity: Lower frequencies (e.g., sub-7 GHz and cmWave) offer wider coverage (with indoor penetration) with lower throughput, while higher frequencies (e.g., mmWaves, THz, FSO) provide higher capacity but require precise beamforming.
The design of efficient HAPS-based NTNs requires balancing trade-offs between coverage, capacity, and resilience. A hybrid multi-band approach can enhance adaptability by leveraging different frequency bands for specific link requirements. For example, (i) service links may operate at lower frequencies (sub-7 GHz, cmWave), to provide broad coverage and reliable signal propagation [4, 2, 1]; (ii) feeder links may operate at cmWave and mmWave bands, to enable high-capacity, long-range transmission [142]; (iii) Inter-platform and backhaul links require ultra-high-speed data transfer, and may operate at sub-THz, THz, and FSO frequencies for seamless network integration. This multi-band, hybrid spectrum strategy ensures a balance between coverage, capacity, and system resilience, supporting the development of robust HAPS-assisted NTN architectures.
Figure 3: Transparent NTN architecture.
V-D 3GPP Specified Architectures
As introduced earlier, in the 3GPP NTN framework [2], two primary platform payload configurations are defined: transparent (or bent-pipe) and regenerative. The choice of payload configuration determines the NTN architecture at a fundamental level, shaping its operational behavior. In the transparent architecture (see Fig. 3), the platform acts as a transparent relay, forwarding signals directly without significant processing. A bent-pipe payload typically implements frequency conversion, analog filtering, and signal amplification. This straightforward design results in reduced payload complexity, lower power consumption on the platform, and generally lower development and deployment costs. However, the absence of onboard signal processing limits the system’s ability to handle dynamic traffic or execute advanced processing tasks. These functions are offloaded to the ground station, leading to a round-trip signal path (from the UE to the NTN platform and then to the ground station) which inherently increases latency.
The regenerative architecture (see Fig. 4) involves onboard signal processing, including tasks such as demodulation, decoding, and routing, in addition to the functions performed in the bent-pipe configuration. This onboard processing capability enables dynamic traffic management, such as load balancing based on QoS requirements. Platforms with sufficient computational resources can execute some UE tasks independently of the ground station, significantly reducing latency for time-sensitive applications. However, this capability comes at the cost of higher payload complexity and power consumption. The advanced technology required for regenerative systems also increases development and operational expenses.
Both transparent and regenerative NTN configurations can be implemented in multiple ways, depending on the functionalities assigned to the platform. In a transparent (bent-pipe) architecture—used in 3GPP Rel-17 and Rel-18—the NTN platform acts as a radio repeater (with or without beam control) while the gNB remains on the ground; this limits payload complexity and enabled early deployment. Notably, the Rel-17 “Satellite Access Node” definition in 3GPP RAN4 still allowed some distribution of gNB functions between Earth and space while keeping RAN4 RF/performance specifications applicable. In contrast, Rel-19 introduces a regenerative (packet-processing) architecture in which a complete gNB is placed on the platform [232]. This enables native use of the $X_{n}$ interface over inter-satellite links (inter-gNB mobility, resource coordination, energy saving), on-payload packet switching and routing/handovers directly between satellites with minimal ground involvement, and reduced $U_{u}$ round-trip time (e.g., for random access and hybrid automatic repeat request). Regenerative payloads also provide the building block for Rel-19 “store-and-forward” (S&F) for IoT UEs, allowing continued operation if the feeder link is severed and even supporting UE-to-UE termination in space. These gains come at the expense of higher payload power and complexity. Earlier alternatives with only a gNB-distributed unit onboard (gNB-central unit on the ground) were found suboptimal due to heavy F1 signaling over the feeder link, lack of a standardized direct inter-satellite interface, and limited support for feeder-link switching and S&F. Consensus therefore converged on the full gNB-on-board approach.
Figure 4: Regenerative NTN architecture.
Regardless of the chosen architecture, NTN platforms can be integrated into a multi-connectivity (MC) framework for uplink, downlink, or bidirectional communication [2]. In this context, a UE can connect simultaneously to both an NTN platform and a terrestrial access point, or to multiple NTN platforms. This approach improves throughput, enhances reliability, and ensures seamless service continuity in heterogeneous service areas.
The role of MC in integrated TN-NTN systems depends on the operational objective. Beyond conventional dual- or multi-node connectivity, secondary MC nodes may be activated on demand to boost throughput. For example, Majamaa, Martikainen, Sormunen, and Puttonen [149] proposed a secondary-node activation algorithm, demonstrating a 9.1% increase in maximum average per-UE throughput compared to systems without MC, and a 5.3% improvement over baseline algorithms.
In addition, MC enables dynamic traffic splitting between terrestrial and non-terrestrial nodes, allowing the network to balance load across layers and offload traffic from congested terrestrial nodes to NTN links. Alam et al. [18] introduced a framework for UE association, BS transmit power and activation, and bandwidth allocation between terrestrial and non-terrestrial tiers. Simulation results showed a 45% reduction in total power consumption and an approximately 250% increase in average throughput compared to standard 3GPP-based network models.
Within integrated NTN systems, HAPS can assume diverse roles: from serving as terrestrial network extensions or satellite signal repeaters to operating as full-scale network centers equipped with baseband units and edge data processing capabilities. These roles directly impact latency, energy efficiency, and coverage performance. For instance, Erdogan et al. [65] proposed a quasi-stationary HAPS to enhance inter-satellite connectivity via FSO links with LEO satellites. Simulations under extreme conditions, such as volcanic activity, demonstrated the feasibility of supporting optical inter-satellite communication. Experimental evaluations further confirm the versatility of HAPS-based systems. Hasegawa, Konishi, Ohta, and Nagate [89] used an automatic gain control system, which enabled a HAPS wireless repeater compatible with 5G NR to maintain stable transmission at maximum power, ensuring wide-area coverage. Similarly, Lou et al. [142] analyzed different HAPS-based link architectures and showed that HAPS can reduce latency in UAV ad-hoc networks, enhance energy efficiency in UAV cell-free architectures, improve coverage probability in satellite cell-free networks, and expand the service area of macro base stations through integrated access and backhaul designs.
Nevertheless, real-world data highlight operational limitations. Empirical analysis of Loon project flight tracking data revealed that while stratospheric balloons offer a cost-effective, delay-tolerant connectivity service, outages can last several hours, limiting the achievable service stability compared to planned terrestrial networks. Such constraints may explain the current trend toward fixed-wing HAPS designs in modern projects, which promise greater reliability and sustained performance [189].
V-E HAPS-Based Projects
Over the past decades, numerous projects have explored various applications of HAPS, particularly in communication, but also in fields like surveillance, disaster management, and environmental monitoring. While some of these projects achieved promising results, many encountered significant challenges that led to their closure, even at advanced stages of development. These challenges included technological difficulties, unsustainable financial models, or regulatory hurdles.
One notable example is the Loon project [229], initiated by X Development LLC, formerly part of Google’s innovation-focused division. The project aimed at extending Internet connectivity to underserved and remote areas using stratospheric balloons. Loon achieved several milestones, showcasing the potential of HAPS-based solutions. For instance, in 2017, the team successfully navigated balloons to disaster-stricken regions, providing emergency connectivity in Peru during severe flooding and in Puerto Rico after a hurricane destroyed critical ground infrastructure. In 2020, Loon partnered with Telkom Kenya to launch a commercial Internet-via-balloon service for Telkom Kenya subscribers, marking a significant step toward scalable deployment. The balloons, following a regenerative architecture, were equipped with an e-NodeB and various antennas to transmit LTE signals to ground UEs and communicate with each other and with gateway stations [189].
However, despite an important role and a number of delivered breakthroughs in both academia and industry, the Loon project was ended in 2021. Its shutdown stemmed from many reasons including economical (high capital expenditure and operating costs), regulatory (spectrum allocation restrictions, international airspace coordination and data sovereignty issues), technological (channel modeling and network planning challenges, complicated station-keeping) and environmental/resource (helium scarcity, power limitations from photo-voltaic panels, occasional environmental pollution caused by polyethylene balloon wreckage) [244].
Another ambitious project, Aquila, initiated by Facebook in 2014, aimed to develop a HAPS system based on solar-powered unmanned aircraft to deliver broadband connectivity to underserved and remote regions. Operating in the stratosphere, Aquila was designed as a quasi-stationary relay providing line-of-sight coverage over areas of several thousand square kilometers, complementing terrestrial networks where infrastructure deployment is economically unviable. The project emphasized long-endurance flights (target: 60–90 days), solar energy harvesting, and mmWave communication payloads for high-capacity air-to-ground links. Despite promising early demonstrations and due to difficulties with in-house aerospace expertise led to insufficient solar energy harvesting and battery efficiency, inadequate flight endurance and, above all, structural integrity issues, their own aircraft development was ended in 2018 in favor of ecosystem partnerships with aerospace firms such as Airbus [87].
Zephyr, a series of HAPS aircraft initially designed by Qinetiq and produced by Airbus, effectively became a successor of Aquila [171]. Now it is an ongoing fixed-wing solar-powered HAPS project by Airbus, is designed as a payload-agnostic platform capable of extended operation in the stratosphere [10]. It provides persistent surveillance and communication services with various types of payloads. Zephyr supports both transparent and regenerative network architectures, utilizing either its default in-house connectivity payload designed to extend 4G and 5G networks, or by integrating other compatible technologies, thereby functioning as platform-as-a-service. Combining a wide range of monitoring, tracking, sensing and detection applications with a communication payload, this platform can offer great opportunities for intelligent network optimization and QoS improvement.
Similarly, HAPSMobile, a collaborative project between AeroVironment and SoftBank, aims at establishing a network of Sungliders [198] to provide connectivity to billions of people lacking basic wireless communication, and to supply additional bandwidth for emerging 5G and IoT networks [16]. SoftBank also announced that, in partnership with the U.S.-based company Sceye Inc., it plans to launch pre-commercial HAPS stratospheric telecommunications services in Japan in 2026. [199]. In these systems, where the HAPS acts as a flying BS, a regenerative payload architecture is the likely choice, although details are hard to find at this stage.
Failures of the Loon and Aquila projects highlighted several major lessons that must be learned to enable more successful future development of HAPS-based solutions. From a technological perspective, it is crucial to ensure platform maturity through iterative prototyping and proper alignment of expertise. Aquila’s structural integrity issues and limited endurance stemmed from insufficient aerospace-domain knowledge. It is also important to consider economic sustainability in addition to technical feasibility. Technological proof-of-concepts (e.g., Loon’s over 300-day flights or Aquila’s mmWave links) are insufficient without a viable business model. Early-stage techno-economic analyses, including sensitivity to subsidies or public–private partnerships, are required to avoid over-optimistic assumptions. Similarly, environmental and resource constraints must be accounted for in the design phase. Helium scarcity limited Loon’s scalability, while Aquila’s solar–battery efficiency proved inadequate under variable conditions. Consequently, it is essential to adopt sustainable alternatives aligned with 6G sustainability goals, emphasizing energy-efficient NTN solutions to reduce ecological footprints. Regulatory foresight and spectrum allocation strategies should also be embedded into project roadmaps and aligned with International Telecommunication Union (ITU) and 3GPP frameworks in order to avoid conflicts with local regulations and international coordination challenges. These issues may be mitigated through closer academia–industry collaboration and effective partnerships across multiple industrial sectors, including aerospace, communications, energy harvesting, and materials engineering. For this purpose, future HAPS research should prioritize system-level prototyping using open-source tools, demand-driven deployment strategies, and technological cooperation. As an example of such collaboration, the HAPS Alliance, founded in 2020, brings together globally recognized companies from telecommunications, technology, aviation, and aerospace, as well as academic research groups. Moreover, shifting from competition for market share among terrestrial providers, SatCom operators, and emerging HAPS-based projects toward mutually beneficial collaboration may represent a key step toward a fully integrated TN-NTN architecture incorporating multiple layers to achieve truly ubiquitous and seamless connectivity.
VI HAPS Technologies
Technology achievements enable HAPS to enhance connectivity, optimize performance, and ensure sustainable operations. As an example, Karabulut Kurt et al. [116] explored the application of advanced techniques such as massive MIMO and mmWave communication into HAPS systems, identifying these as promising directions for future research. Similarly, d’Oliveira, Melo, and Devezas [55] discussed key technological developments, trends, and challenges associated with HAPS. Nonetheless, this work primarily adopts an aeronautical engineering perspective, and lacks a thorough analysis from a telecommunications point of view. In addition, the study was published in April 2016, and does not reflect the most recent developments in the field.
In this section, we address these limitations, and present an updated and comprehensive overview of recent technological trends that support the large-scale adoption of HAPS in the context of 6G networks. We focus on key enablers such as channel modeling (Sec. VI-A), resource management (Sec. VI-B), user assignment and cell association (Sec. VI-C), multi-link connectivity (Sec. VI-D), mobility management and handover (Sec. VI-E), antenna (Sec. VI-F), channel estimation (Sec. VI-G), interference management and beamforming (Sec. VI-H), energy efficiency (Sec. VI-I), ML (Sec. VI-J) and Cybersecurity (Sec. VI-K) techniques for HAPS (See Table LABEL:tab:haps_technologies_summary in Appendix). In order to review technology achievements systematically we classify studies using a multi-dimensional taxonomy based on following dimensions: research methodology (analytical, simulation-based, experimental); network integration level (standalone HAPS, HAPS-TN, HAPS-NTN, HAPS-TN-NTN, no HAPS); and related frequency bands (Sub-7 GHz, mmWave/FR2, Sub-THz/THz, and FSO). From Table LABEL:tab:haps_technologies_summary, it can be observed that analytical research methodology dominates the body of HAPS-related studies. This prevalence reflects the current maturity level of HAPS technology, where theoretical frameworks enable rapid feasibility assessments, while simulation-based investigations are limited by the availability of validated channel and system models. Experimental studies remain scarce due to regulatory, financial, and technical barriers. Consequently, researchers often rely on self-developed simulation models with limited capabilities in order to validate their analytical findings. Nevertheless, collaborative initiatives—such as those led by the HAPS Alliance—may facilitate the collection of empirical data and contribute to the standardization of datasets for more reliable and reproducible simulation studies. It is also worth noting that the majority of existing studies focus on the sub-7 GHz and mmWave/FR2 frequency ranges, while the emerging cmWave and THz bands remain comparatively underexplored.
VI-A Channel Modeling
Channel modeling for HAPS is evolving with new technologies, improving accuracy in predicting signal propagation, especially based on advanced simulation techniques and real-time data. However, existing HAPS-related channel modeling studies predominantly focus on HAPS-to-ground links, whereas inter-platform channel characteristics remain comparatively underrepresented.
Network simulators often lack realistic path loss and fading modules. To address this, models for rain, cloud, gaseous absorption, and Ricean fading have been developed. Using QualNet, Zheng, Wang, and Meng [251] showed how different attenuation types affect the packet delivery rate of HAPS-to-ground links on 10, 15, 20 and 25 GHz communication frequency, improving simulation realism and accuracy. Also, Zain et al. [241] showed Ka-band rain attenuation is about 10× higher than in C-band and strongly location-dependent, and presented an analytical model based on the ITU-R empirical model to predict Earth–space attenuation for fully and partially rain-affected paths. For a 1-km cell, attenuation reaches 43 dB at 0.01% availability and roughly doubles as cell size increases.
To address the complexity of propagation models in suburban and urban environments, Hsieh et al. [97] applied ray tracing simulations to study HAPS-to-ground signal propagation in Chicago downtown. They developed a path loss model in order to estimate total propagation loss by considering terrestrial clutter and free-space propagation. The model adapts to different aerial vehicle heights and elevation angles using two environment-specific factors, and the study also analyzes multipath characteristics to support system design and fast fading modeling. To examine the inter-platform propagation, Yılmaz et al. [237] analyzed a channel model of HAPS-to-HAPS links in order to determine its capacity while considering polarization mismatches, atmospheric gases, rain, and clouds/fog. Their simulations revealed the path gain characteristics and channel capacity across different antenna types, platform heights, and carrier frequency scenarios. Meanwhile Palanci et al. [170] modeled HAPS-to-satellite communication channels with a focus on ionospheric effects. It showed that plasma density variations influence the channel while magnetic field effects on the refractive index can be neglected. Using the Appleton-Hartree model and an adapted Chapman model, the authors analyzed how Total Electron Content impacts path loss, supported by extensive numerical results.
A multi-state statistical channel model for HAPS communication in the Ka-band was proposed by Zhao et al. [248] to account for tropospheric weather effects and large-area user mobility. The model uses three fading states—LoS, moderate shadowing, and deep shadowing—represented by a Markov chain. It also provides the probability density function of the received signal power, a method for generating time-series, and evaluates BER performance using the proposed model.
Guan et al. [83] show that HAPS mobile channels invalidate the terrestrial plane-wave assumption, requiring consideration of both elevation and azimuth spreads. Large delay differences, Doppler shifts, and strong spatial correlation further complicate analysis. By characterizing dispersion across time, frequency, and space, the authors develop a wideband tapped-delay-line HAPS channel model with environment-specific parameters to support simulation and verification of broadband HAPS systems.
VI-B Resource Management
Given the recent relevance of HAPS in wireless communication, RA has become one of the key areas of investigation, including the positioning of the aerial tier units within the network. While multiple parameters must be optimized to address the RA problem comprehensively, more practical solutions operate under specific circumstances and constraints, defining optimization objectives accordingly. In this subsection, we will consider various system models, from single-tier (based on HAPS) to multi-tier (also known as space-air-ground integrated networks) architectures.
VI-B1 Single-tier architecture
Ibrahim and Alfa [100] maximize the number of UEs receiving multicast session transmissions in an orthogonal frequency division multiple access (OFDMA) frame under QoS constraints by decomposing the problem into two linear programs solved iteratively. Sub-problem 1 selects UEs and assigns frequency–time slots via three Lagrangian-relaxation methods. Sub-problem 2 allocates power, modeled as a continuous knapsack, solved optimally with a greedy algorithm. The scenario assumes a single HAPS serving GUEs in a defined area.
In a multi-platform setup with several HAPS serving an area, Grace et al. [80] exploits fixed user dishes with narrow beamwidths steered to their main HAPS to boost bandwidth efficiency. Two cases are studied—single-beam per HAPS and multi-beam per platform—and results show higher capacity when multiple HAPS share the same spectrum, with gains that scale as the number of users grows.
In multi-cell settings, both inter- and intra-beam interference are significant. In integrated networks, overlapping coverage between terrestrial base stations and HAPS-based BSs—especially with co- or adjacent-channel operation—induces inter-system interference, making effective interference coordination essential for future integration (see Sec. VI-H) In this context, Liu et al. [140] proposed a dynamic spatial–time–frequency coordination method between HAPS-mounted BSs and TBSs. Two service modes are considered: direct HAPSGUE links and indirect access where HAPS serve TBSs that, in turn, serve GUEs. UEs associate to HAPS (direct or indirect) or TBSs based on maximum received power. The scheme distributes traffic load and employs an ergodic search over allocation patterns across HAPS and terrestrial networks. Simulations show the best average per-cell throughput for user groups at any distance from HAPS. Unlike the single-HAPS model in [140], Wei et al. [227] studies a HAPS network via stochastic geometry, modeling TBSs and HAPS as Poisson point processes and deriving a closed-form coverage probability for spectrum sharing. Results show HAPS cause minimal interference to TBSs, supporting spectrum sharing in integrated networks.
VI-B2 Multi-tier architecture
In the space domain, LEO satellite systems like Starlink by SpaceX are already extending communication services to areas beyond the reach of terrestrial networks. In January 2024, Starlink launched its first six satellites equipped with direct-to-cell capabilities, enabling device-to-device connections with unmodified LTE phones, and providing text messaging services integrated with mobile operators’ core networks [203]. To achieve this, Starlink developed large, advanced phased arrays, measuring 2.7 m x 2.3 m, which feature highly sensitive receivers and powerful transmitters. These arrays allow the satellites to function as cell towers in space, capable of communicating with cellphones equipped with low-gain antennas and limited transmit power. The goal faces significant technological challenges, particularly in densely populated areas with many users. Addressing these requires advancements in both network architecture and RA. An integrated multi-tier SAGIN can provide major performance benefits through deployment across diverse nodes and links.
Li et al. [134] focused on RA optimization within a multi-tier architecture where LEO satellites provide wide-area coverage, while HAPS and full-duplex terrestrial relays cover hot-spot regions to enhance capacity. Additionally, GEO satellites were considered as routing nodes. The authors designed a hierarchical resource management framework for this system, and proposed an optimization method to maximize LEO-to-ground and HAPS-to-ground downlink throughput based on predictions derived from the regular orbit information of LEO satellites. In the proposed architecture, mobile UEs in hot-spot regions connect via HAPS or terrestrial relays, while other areas use direct LEO links. This approach boosts system capacity in hot-spots and lowers handover rates. Different frequency bands are allocated for space/air-to-ground, terrestrial, space-to-air, and inter-satellite links. Capacity analysis and simulations show that the multi-tier design greatly outperforms satellite-only systems. Additionally, dynamic radio resource optimization improves the average throughput of HAPS and LEO beams by over 10%. In this scenario, spectrum sharing is required if HAPS and satellites are designed to transmit on the same or adjacent frequency bands, or when deploying HAPS in a region already covered by a separate satellite system.
Meanwhile, a LEO-HAPS spectrum sharing technique under non-ideal spectrum sensing was studied by Wang et al. in [222], where a multi-beam LEO served as the primary system and a single-beam HAPS supported GUEs without direct LEO access. A dynamic spectrum and power allocation strategy was proposed to maximize HAPS-GUE throughput with minimal QoS degradation for LEO-associated GUEs. Results showed that HAPS interference to LEO-associated GUEs is negligible and affects only a small fraction.
Later, RA for a multi-tier cognitive satellite-HAPS network with non-orthogonal multiple access (NOMA) was studied by Liu et al. [138]. The model included a primary satellite and a secondary HAPS network sharing spectrum via underlay cognitive radio to serve terrestrial users. The sum-rate maximization problem for the secondary network was addressed under NOMA group connectivity limits, power constraints, and QoS requirements. The proposed joint RA algorithm (sub-channel and power allocation) outperformed orthogonal multiple access, especially as user or sub-channel numbers increased.
Introducing smaller aerial components, such as LAPSs like drones, into the network architecture adds an additional degree of freedom in deployment planning. While this approach can further improve coverage and system capacity, it also complicates RA due to the need for dynamic load balancing and 3D placement of mobile network nodes. An aerial heterogeneous cellular network with HAPS-based macrocells and LAPS-based femtocells was studied by Helmy and Arslan [93]. Using stochastic geometry, the authors analyzed cell loads, outage probability, optimal LAPS placement, and downlink signal to interference plus noise ratio (SINR). Results showed that users associated with an LAPS BS experienced much lower outage probability, and the proposed LAPS positioning algorithm further reduced outages.
Later, Qin et al. [175] investigated joint RA and UAV 3D-location optimization in a cluster-NOMA network with a HAPS macro BS and multiple UAV small BSs. The goal was to maximize IoT uplink sum rate by decomposing the problem into spectrum/power allocation and UAV positioning (horizontal and altitude), solved iteratively. Results showed improved uplink sum rate and spectrum efficiency over benchmark RA schemes.
As previously mentioned, the SAGIN architecture is likely the most promising option for future communication networks, including 6G. Therefore, it is crucial to clarify the features that enable the coexistence and collaboration of HAPS and terrestrial systems within this integrated framework. In [24], a RA framework for an integrated satellite-HAPS-TBS network was proposed to improve GUE downlink throughput under power, access, and backhauling constraints. Two problems were formulated: a short-term problem optimizing access, backhauling, and power with a fixed HAPS location, and a long-term problem optimizing HAPS placement based on average user distribution. The studied system model comprised three tiers: a satellite in the space tier, multiple HAPS in the air tier, and several TBSs in the ground tier, with gateway feeder stations for backhauling. Based on OFDMA, spectrum was divided into resource blocks, with each user served by one BS using one RB, ensuring no intra-cell interference but allowing inter-cell interference among TBSs. No inter-cell interference was assumed for satellites and HAPS due to distinct frequency sets. Simulations showed a notable average data rate gain compared to two baselines: (i) joint optimization with uniform power, and (ii) optimized HAPS placement with random access/backhaul under uniform power.
From the literature, it is clear that aerial network mobility boosts capacity and flexibility but complicates optimization due to 3D positioning, dynamic load balancing, and spectrum diversity. Effective interference management is vital for stability, while a well-designed RA strategy can mitigate inter-tier interference and enable smooth integration with TN.
VI-C User Assignment and Cell Association
Unlike stationary TBSs, HAPS platforms enable flexible deployment that can be dynamically optimized according to user distribution. Consequently, the HAPS deployment problem has been investigated using various analytical models across different application scenarios and architectural configurations.For example, Zong et al. [254] examined HAPS deployment for maximizing network transmission rate while meeting service requirements, where each GUE connects to the HAPS offering the best rate. A game-theoretic, self-organized optimization using a restricted spatial adaptive play algorithm was proposed and shown to converge in simulations. Although simplified compared to full SAGIN architectures, the results highlighted the potential of HAPS for self-organized deployment optimization.
To meet 6G data rate demands, Cumali et al. [53] studied user selection in a downlink NOMA-based multi-user HAPS system. A joint user selection and correlation-based pairing algorithm was proposed to maximize beamforming gain, along with a polar-cap codebook design for limited-feedback HAPS channels under Rician fading. Simulations showed superior data rates over existing NOMA schemes. However, the model did not consider integration with terrestrial/satellite networks or support for aerial users, which are essential for 6G.
While satellite deployment is less flexible than HAPS, its inherently wide coverage area provides opportunities for optimizing user association strategies, especially within multi-tier integrated architecture.Accordingly, the user association problem in the context of computational task offloading has been investigated by Zhang et al. [245]. The authors considered a disaster scenario where TBSs have been destroyed, and two-tier UEs, including both GUEs and UAVs, with limited computational capabilities offload disaster information to HAPS and/or LEO satellite servers for further processing. The study aimed to jointly optimize uplink fronthaul capacity and computation performance by modeling the problem as a weighted 3-uniform hypergraph, where each hyperedge weight represents the combined fronthaul uplink and computing rates. A local search-based algorithm was developed to find a near-optimal 3D hypergraph matching with maximum total weight. The model considered a LEO satellite with an enhanced MEC server providing backhaul to multiple HAPS, which cover disaster areas and offload computation, while UEs use orthogonal sub-channels within shared Ka-band RBs, avoiding uplink interference. In simulations with 6 HAPS and 20 UAVs over a 0.25 km² disaster area, the proposed hypergraph search algorithm achieved significant system sum rate gains over greedy search, showing strong practical potential.
Later, a vHetNet comprising one GEO satellite, one HAPS, and several TBSs was examined by Liu, Dahrouj and Alouini [139]. In this setup, both the HAPS and TBSs are equipped with multiple antennas, enabling them to simultaneously serve multiple GUEs through RF links. The connection between the satellite and the HAPS is established via an FSO link. In this scenario, when a GUE is served by the HAPS, the required data is first transmitted from the satellite to the HAPS, and then relayed to the GUE.The system model employs a space division multiplexing scheme, where all RF links share the same central frequency. Consequently, the performance of the proposed architecture is highly dependent on the intra-cell and inter-cell interference among TBSs, intra-cell interference within the HAPS, and the interference between HAPS and TBSs. The study formulated a joint optimization problem for associating GUEs with either TBSs or the HAPS and designing corresponding beamforming vectors, aiming to maximize sum-rate under constraints of HAPS payload connectivity, FSO backhaul capacity, and power limits of HAPS/TBSs. An iterative algorithm optimized parameters step by step, and simulations demonstrated that integrated satellite-HAPS-terrestrial networks can significantly enhance the sum-rate.
As heterogeneous integrated networks face load imbalance from uneven user distribution and mobility, Tu et al. [215] proposed a priority-based load-balancing scheme for a two-layer downlink SAGIN using network slicing to mitigate underutilization and overload across cells. The proposed two-layer radio access network (RAN) includes TBSs in the terrestrial layer (priority for user association) and UAVs/LEO satellites in the non-terrestrial layer. Overloaded TBSs offload traffic to the non-terrestrial layer via RAN slicing into high-throughput, low-delay, and wide-coverage slices, with lower-priority slices preferentially offloaded. The optimization problem jointly considered throughput, delay, coverage, user association, resource allocation, and UAV deployment. Results showed improved slice throughput and average delay. While HAPS were not included, the study tackled the critical load-balancing challenge in SAGIN. Prior to that, Shahid et al. [192] explored load balancing for 5G integrated satellite-terrestrial networks using multiple radio access technologys. The approach handled both intra- and inter-RAT UE offloading from overloaded cells by introducing the radio resource utilization ratio as a common load metric, enabling efficient traffic redistribution without overloading target cells. Simulations showed improved throughput and QoS. While limited to a GEO satellite, the metric can be generalized to other NTN components, including HAPS. Alam et al. [19] studied load balancing as a joint radio resource management problem—covering bandwidth allocation, UE association, and power control—in an integrated TN-NTN rural downlink network with multiple macro TBSs and LEO satellites. The proposed framework maximized the sum of log throughput across all UEs, improving overall network capacity. The framework tackled the optimization sequentially—first user association and bandwidth allocation under fixed power, then transmit power with the first two fixed. Simulations showed notable throughput gains over the maximum reference signal received power (RSRP) baseline and confirmed that integrating TNs and NTNs in rural areas can greatly reduce out-of-coverage UEs. In their subsequent study [18], the authors proposed a framework for bandwidth allocation, user association, and power control that can balance network capacity and energy consumption based on the current traffic conditions. By adapting the network behavior according to traffic demands, this approach significantly enhances mean network throughput during high-traffic periods and reduces power consumption during low-traffic periods.
Based on the reviewed studies, we can conclude that user association in SAGINs that incorporate HAPS-mounted BSs and are fully integrated with TNs, still requires further research. It is particularly important to consider scenarios involving a realistically sufficient number of mobile ground and aerial UEs. Additionally, there is a need to investigate the user association challenge under the practical constraints imposed by interference management solutions and load balancing.
VI-D Multi-Link Connectivity
As mentioned earlier, some emerging 6G applications, such as aerial vehicle control and autonomous driving, impose stringent requirements in terms of both latency and connection reliability, a paradigm referred to as URLLC. Meanwhile, the Stage-2 architectural framework for MC across E-UTRA and NR is defined by 3GPP TS 37.340 [7]. It outlines the functions of the master node and secondary node, bearer splitting and duplication mechanisms, and the related signaling architecture that permits simultaneous connectivity to multiple radio nodes. In order to make clear how these processes might be used to enhance service continuity and resilience in HAPS-enabled NTN deployments, we have broadened the conversation.
In this context, multi-link connectivity (or MC) has shown the potential to meet the URLLC service requirements [187]. Specifically, MC architectures can be categorized at the BS level (i.e., the user connects to multiple BSs within the same network, usually managed by the same operator), at the network level (i.e., the user is connected to different networks, possibly from different operators), or at the technology level (i.e., the user connects via different RATs) [186]. In this paper, we focus on the first category, which also includes techniques such as coordinated multi-point (CoMP), which gained significant importance in LTE-advanced and 5G, where multiple BSs within the same network simultaneously serve the same user.
Li et al. [133] studied air-to-air communication using a 3D model integrating aerial BSs, aerial UEs, TNs, and GUEs. Two UAV types were considered: aerial BSs connecting via HAPS/TBS, and aerial users relaying data to GUEs. A joint-transmission CoMP scheme was proposed where each aerial user is served by four aerial BSs forming a tetrahedral cell. The study derived data rates and coverage probabilities, and introduced a frequency allocation scheme to reduce inter-cell interference. Compared to the binomial-Voronoi dynamic cooperation set, the approach achieved similar performance with much lower complexity and signaling overhead. Meanwhile, Amer, Saad, and Marchetti [25] studied aerial UE connectivity using CoMP transmissions for UAVs served by TBSs. Two scenarios were analyzed—static hovering and mobile UAVs—with coverage probability bounds derived and handover rates evaluated via a custom mobility model. Results showed that CoMP transmission greatly improves achievable coverage probability. Prior to that, Liu, Zhang, and Zhang [136] considered an uplink CoMP scenario where UAVs collect and forward signals from GUEs. A UAV placement and movement scheme was proposed to maximize average throughput for mobile GUEs, though UAV energy constraints remain a critical challenge. Recently, Salehi, Ozger, and Cavdar [186] proposed an MC SAGIN algorithm for remote aerial vehicle (AV) piloting beyond LoS. The system model included LEOs, HAPS, and TBSs for AV communication, along with direct inter-AV links. However, it excluded LEO-HAPS and LEO/HAPS-TBS links, assuming both LEO and HAPS connect to the core network via their ground stations. The transmission strategy in [186] used message cloning, sending multiple copies via all available links for independent decoding at the AV. The model incorporated practical antenna setups, unreliable backhaul, buffer queues, and fading effects. Reliability and availability of different MC configurations were evaluated through simulations, with brute-force optimization used to find the optimal MC path ensuring QoS. Results showed that LoS interference strongly impacts link performance, and MC is vital for safe operation of certain AVs. At thesame time, latency-sensitive scenario was explored by Sadovaya et al. [183], which addressed the issue of MC computational task offloading in NTN. The system model featured a HAPS and several UAVs providing on-demand MEC services to randomly distributed GUEs. The simulation results showed that MC offloading strategies can significantly reduce task computation delay.
VI-E Mobility Management and Handover
Mobility management faces major challenges in seamless handovers, especially for LEO and HAPS systems. The architectural and functional principles of connected-mode mobility and handover (HO) in NR and LTE are specifically described in 3GPP TS 38.300 and TS 36.300 [9, 6], respectively. Referencing the specific handover signaling and timing protocols outlined in the related Radio Resource Control Radio Resource Control (RRC) standards, these specifications specify the functions of radio access network (RAN) nodes, mobility statuses, and high-level handover concepts. In the context of HAPS-assisted mobility, where wide coverage regions and quasi-stationary platforms produce different handover dynamics than terrestrial networks, its applicability to NTN scenarios is examined.
Juan et al. [112] evaluated the standard 5G NR handover procedure for a high-speed train in a LEO-based deployment versus a 3GPP urban macro TN. Simulations showed radio connection failures during handovers were 10× higher and outages 5× longer in the NTN scenario. Further, an intra-satellite mobility solution was proposed [113] that avoids UE radio measurements by exploiting predictable satellite trajectories and antenna gains. Simulations in rural and urban scenarios showed reduced service failures and unnecessary handovers, with improved downlink SINR. Zhang et al. [243] introduced a convolutional neural networks-based handover approach for LEO satellite network. The authors modeled the handover process as a directed graph, where a user with (partial) knowledge of its future signal strength can identify the optimal handover sequence. The CNN was trained to learn near-optimal handover strategies from historical signal strength data. Voicu, Bhattacharya, and Petrova [218] compared three inter-satellite handover methods for space-to-ground performance: Closest Satellite, Max Visibility, and SINR-based. The first two improved spectral efficiency and reduced delays, while the SINR-based approach minimized the number of handovers. Warrier et al. [226] introduced a graph theory-based algorithm for UAV handovers between terrestrial and satellite networks. By adapting in real time to UAV mobility and network changes, it minimizes latency and disruptions. The algorithm considers RSRP, signal to noise ratio (SNR), and elevation angle to optimize handover timing and ensure smooth connectivity.
The quasi-stationary nature of HAPS in the stratosphere introduces additional challenges for conventional handover algorithms, which rely on fixed thresholds. He, Cheng and Ni [91] introduced an adaptive handover algorithm, addressing the limitations of fixed-threshold methods, in order to handle possible HAPS disturbance. Using a least mean square prediction-based power model, the scheme adjusts handover parameters according to terminal speed and the confidence levels of expected signal strength, improving reliability under HAPS mobility conditions. Further [90] a cooperative transmissions scheme was proposed for HAPS systems to mitigate outage probability and long service interruptions. The algorithm adaptively selects the platform with the highest channel gain based on user mobility and real-time channel conditions, improving reliability and service continuity. Later [224] Wang, Li, and Zhou analyzed the impact of HAPS vertical and swing movements on handover performance. A refined ground coverage model and analytical expressions were developed, showing that path loss strongly affects coverage, with swing movement exerting a greater influence on handover probability than vertical movement. Menwhile, Li et al. [131] developed various handover techniques for an integrated LEO-HAPS network, and evaluated the handover delay and signaling overhead. Specifically, they formulated the problem of determining the optimal handover triggering time, which was analytically solved using Lagrange duality. Later, Almuallim et al. [23] studied a mobility management solution for integrated HAPS-ground networks, allowing aerial users to connect to either terrestrial BSs or a HAPS. The approach focused on minimizing unnecessary handovers while ensuring QoS.
VI-F Antenna Technologies
Integrating advanced antenna technologies, such as massive multiple-input multiple-output (mMIMO), spatial multiplexing, digital beamforming, and RISs, into HAPS systems can improve network capacity and efficiency across a range of diverse and challenging environments. For example, Dicandia and Genovesi [60] evaluated the advantages of using a triangular (rather than square) lattice for phased arrays in HAPS systems operating in the 5G NR n257 and n258 frequency bands. By considering a 64-element planar array, the results demonstrated that the triangular lattice array offers superior performance in terms of array gain, average sidelobe level, and mutual coupling, making it a more effective choice for 5G mMIMO systems. Meanwhile, Tashiro, Hoshino, and Nagate [208] proposed a cylindrical mMIMO system for HAPS. They showed that the cylindrical mMIMO shape provides up to 2.1 times higher capacity compared to traditional planar arrays, with an extended coverage of up to 100 km in radius, and improves the SINR by approximately 10 dB at the cell edge. Further, in order to consider the coexistence with terrestrial systems, the authors introduced a two-stage precoding system for TN-HAPS spectrum sharing using phased array antennas [209]. The first stage forms nulls toward TN cells, and the second directs beams to HAPS users. With a 196-element cylindrical array, the approach improved SINR for TN users by up to 10 dB over a conventional minimum mean square error (MMSE) precoder. Cylindrical antenna configuration was later considerd by Shafie et al. [191]. The authors proposed a vertical uniform linear array sector-based cylindrical antenna for HAPS, combined with a NOMA clustering method to mitigate strong geographic user correlation. An optimization algorithm for power distribution was introduced to maximize energy efficiency, spectral efficiency, and QoS using successive interference cancellation. Simulations showed that spatial correlation strongly affects spectrum and energy efficiency in multi-antenna HAPS systems.
Alternative antenna configuration was studied by Abbasi, Yanikomeroglu, and Kaddoum [12] to address the suboptimal gain for users located beneath the HAPS. The authors introduced a hemispherical antenna array for HAPS to maximize gain by aligning users with specific antenna elements. Using analog beamforming, user data rates were estimated from steering vectors, while an antenna selection algorithm and Bisection-based power allocation optimized performance. Results showed cumulative data rates up to 14 Gbps. The authors also proposed a cell-free scheme to mitigate inter-cell interference between GUEs and neighboring UAV-BSs [14]. The scheme operates in two stages: GUEs send messages to UAVs in the sub-6 GHz band, and UAVs then allocate resources and beamform signals to a HAPS in the sub-THz band acting as a CPU. An optimization problem was formulated to maximize the minimum SINR by optimizing power allocation and UAV positions.
Regardin the UE antenna systems, Kim et al. [122], proposed a dual-polarized shared aperture antenna for Ku- and Ka-bands, developed through a four-step derivation process using slot aperture coupling to achieve broadside radiation in both bands. Based on array factor theory, the optimal inter-element spacing was determined, resolving the trade-off between beam steering angles and array spacing in dual-band systems. The use of RISs is another approach to improve signal quality and coverage. In fact, low-cost RIS components enable customizable communication channels, LoS connectivity, and obstruction avoidance [235]. When combined with traditional mMIMO techniques, RISs can also boost data rates and reliability. In a system where HAPS relay data via RISs due to transmission blockages, the RIS acts as an amplify-and-forward relay between source and ground users. This approach was evaluated by Odeyemi, Owolawi, and Olakanmi [167] through metrics like BER, capacity, and outage probability, with analytical and asymptotic results showing significant performance improvements for HAPS and RISs integration.
VI-G Channel Estimation
6G is expected to use high-frequency bands such as cmWave, mmWave or even sub- THz, to support new scenarios for seamless ubiquitous 3D connectivity (see Sec. V-C). However, the channel conditions in these bands will become more variable, suffering from significant path loss, fading, and interference. Thus, accurate channel estimation is necessary to adapt transmission parameters dynamically, ensure reliable connectivity, and maximize data throughput while minimizing latency. Advanced techniques like beamforming, mMIMO, and ISAC, also require precise channel state information (CSI) via channel estimation, to provide reliable communication services. In this context, Iskandar and Aziz [103] evaluated LTE downlink performance over HAPS, comparing least square and MMSE channel estimation under a Ricean fading model, and found that GUEs at low elevation angles suffer degradation from multipath and Doppler. Later Hidayat and Iskandar [94] analyzed a single-carrier LTE HAPS channel, showing that elevation angles above 40° yield optimal conditions, with performance influenced by bandwidth, modulation, and Doppler shifts. More recently, Rahayu and Iskandar [176] compared least square, MMSE, and MMSE+ discrete Fourier transform (DFT) estimation methods for HAPS-based OFDM systems. Also, Nawaz et al. [164] estimated sparse HAPS channels using a superimposed pilot sequence method that leverages prior knowledge of UEs’ locations. The model featured a HAPS-based beamforming multi-antenna BS serving mobile GUEs, with backhaul via ground stations and satellites. Numerical results examined the influence of antenna configuration, channel sparsity, SNR, and interferer density. Later, Tekbıyık et al. [210] studied a vHetNet combining GEO/LEO satellites, HAPS, and TN with full-duplex RIS-assisted backhauling over HAPS channels. A graph attention network-based channel estimator was proposed to address channel variability and hardware impairments, requiring estimation at only one transmitting node without on-off control or half-duplex switching. Simulations confirmed strong performance across diverse channel and hardware conditions. Meanwhile, Güven and Karabulut Kurt [85] studied a HAPS-LEO architecture to address channel estimation and synchronization challenges. A CNN-based estimator was proposed to mitigate carrier frequency offset from residual Doppler effects, with results showing improved mean square error and BER performance over benchmark methods. And recently, Guo et al. [84] analyzed the effect of channel estimation errors in an integrated satellite-HAPS-terrestrial communication system, and proposed a novel NOMA-based power allocation scheme to satisfy QoS requirements.
VI-H Interference Management and Beamforming
Interference management and beamforming technologies enhance HAPS communication by jointly improving signal strength and reducing interference in remote and demanding areas. A comprehensive analysis of stratosphere-to-Earth co-channel interference caused by HAPS was presented by Milas and Constantinou [156], who introduced a methodology for assessing the impact of this interference component on TNs by evaluating performance degradation. The methodology considered various factors, including the mobility of HAPS, the realistic distribution of azimuth and elevation angles of terrestrial microwave links, and the gradual loading of HAPS structures. A 3D mMIMO system was examined by Hu, Hong and Evans [98], who considered an uplink scenario with pure LoS propagation from multiple GUEs to a HAPS mounting a horizontal planar antenna array. The study modeled intra-cell interference using Beta distributions, and demonstrated that adjusting the altitude of the array can significantly reduce interference, thereby improving key performance metrics such as signal to interference ratio, coverage, and throughput.
Later, Guan et al. [82] introduced a space-division multiple-access scheme with a multi-beam antenna and beamforming algorithm that adapts array weight vectors to improve performance and reduce interference. An improved elliptical cell design balanced SNR by broadening inner beams and narrowing outer ones, expanding inner cell coverage while shrinking outer areas. This reduced the number of required cells while maintaining coverage quality. The integration of space-time block coding with multi-beamforming further enhanced coding and diversity gains without extra bandwidth.
Fujii, Ohta and Fujii [70] analyzed a multi-cell HAPS system, where capacity is limited by feeder link bandwidth. To enhance spectrum efficiency, a multi-gateway approach was proposed, allowing gateways to share feeder link frequencies to reduce interference and increase capacity. Building on this, the authors [69] introduced interference cancellation techniques for the multi-gateway feeder link, further improving transmission quality and overall communication performance. To further reduce interference, Ishikawa et al. [102] introduced a null-sweeping technique where the HAPS dynamically adjusts null points, while terrestrial BSs schedule GUEs accordingly. This coordinated approach improved the SINR for GUEs without degrading the HAPS link quality, resulting in up to a 15% increase in ground system capacity.
Considering HAPS-aided SAGIN architecture, Kawamoto et al. [120] aimed to mitigate the interference between HAPS-to-GUEs links and back lobes from HAPS-to-satellite links. Unlike traditional zero forcing, which allocates random nulls, the proposed method generates targeted nulls based on GUE locations using a codebook of antenna patterns tied to weight matrices, enabling dynamic adaptation to reduce interference and improve communication quality.
VI-I Energy Efficiency
A key challenge for emergency communication is ensuring energy-efficient transmissions due to the limited power resources available during large-scale disasters. Integrated HAPS/satellite networks can help overcome this issue by optimizing power consumption while maintaining robust connectivity. In this context, Dong et al. [62] proposed an adaptive link-state advertisement routing algorithm designed for a slow-flat Rician fading channel to estimate link information. This approach can reduce energy consumption at the terminal by selecting the most energy-efficient path.
In order to improve energy efficiency, Rallage and Kandeepan [177] proposed a low-power-consumption decoding and forward relay system via HAPS.The results demonstrated that, in certain channel conditions, a relay-based system can outperform a non-relay-based system in terms of energy efficiency. Prior to that, Xing et al. [232] studied HAPS as repeaters and aerial BSs, with energy efficiency assessed via consumption factor theory in single- and multi-cell scenarios using Monte Carlo simulations. Results showed both architectures achieve good downlink spectral efficiency, though coverage is constrained by GUE transmit power and antenna gain. Performance in uplink and downlink can be further improved with directed GUE antennas on LTE band 1. Also, Alfattani et al. [22] proposed an energy-optimization algorithm for HAPS, enabling mode switching to reduce consumption and extend loitering duration. The modes include a low-power HAPS-RIS mode for passive communication and a HAPS-SMBS mode for enhanced computing, caching, and communication. The system prioritizes the passive mode, activating higher-energy modes only when additional capacity is required. Finally, Salamatmoghadasi, Mehrabian, and Yanikomeroglu [185] studied a vertical HetNet with a HAPS as an SMBS, alongside an macrocell base station (MBS) and multiple small base stations, to improve energy efficiency. An algorithm was proposed to minimize consumption by leveraging the HAPS and dynamically managing SBS sleep modes based on real-time traffic load, reducing power use by 30% while maintaining QoS.
Li et al. [135] studied energy-efficient RA in SAGINs for remote IoT using UAV relays. The goal was to improve energy efficiency by optimizing sub-channels selection and the uplink transmission power using Lagrangian dual decomposition, and UAVs deployment via sequential convex approximation. The proposed approach achieved more than 21.9% improvement in terms of energy efficiency. Later, Zhu et al. [253] explored the use of HAPS in SAGINs to support data transmission between LEO satellites and smart devices for remote IoT applications. The proposed approach jointly optimized RA and HAPS deployment to minimize the power consumption. By breaking down the complex problem into two sub-problems, the authors found a near-optimal solution. Simulation results showed a significant reduction in power consumption compared to conventional RA strategies, while still satisfying smart device rate requirements. Finally, Zhao and Chang [247] optimized energy efficiency for data collection in a UAV network supported by a HAPS aerial BS. The approach dynamically clustered the target area, selected sink sensing devices, adjusted transmit power, and managed UAV operations. A centralized reinforcement learning (RL) framework based on proximal policy optimization (PPO) jointly optimized UAV trajectories and power, while an Affinity Propagation clustering algorithm identified optimal sink devices for device-to-UAV communication.
Recently, Song et al. [201] evaluated the energy efficiency of a HAPS-based RAN against a terrestrial RAN using a numerical model based on real traffic trends and showed that leveraging HAPS features—self-sustainability, high altitude, and wide coverage—cuts energy consumption by up to 30%.
VI-J Machine Learning
AI has been utilized to tackle several critical challenges facing current and future wireless networks [35]. Notably, ML-based algorithms can be applied to HAPS to optimize key functions such as topology management, RA, handover management, and improve interference mitigation, ensuring more efficient, reliable, and adaptive communication networks [116]. These algorithms can dynamically adjust to changing conditions, reducing disruptions and maximizing service quality.
[…]
RL is increasingly applied to NTNs to optimize resource allocation, platform positioning, and trajectory control. By leveraging Markov decision process (MDP) frameworks, RL enables aerial platforms like HAPS and UAVs to autonomously adapt and enhance network performance under real-world constraints. In their survey, Naous et al. [163] reviewed RL applications for NTNs—including LEO satellites, HAPS, and UAVs—focusing on resource allocation, platform positioning, and trajectory control under NTN-specific constraints. Using MDP frameworks, it provided a taxonomy of current research, assessed study realism, and identified open challenges to narrow the simulation-to-reality gap.
In this context, Nguyen et al. [165] developed a deep RL-based offloading strategy for HAPS-assisted vehicular networks. The partial task offloading problem was formulated as a MDP and solved using the deep deterministic policy gradient (DDPG) algorithm. Among all evaluated schemes, the DDPG-based approach achieves the highest task success rate, with only a 0.7% failure rate when the task size is 1.2 Mbits. Dahrouj, Liu, and Alouini [57] highlighted the advantages of optimizing user scheduling in space-HAPS-ground networks by ensembling deep neural networks. In particular, it was demonstrated that, when the ensemble size increases, the suggested approach performs better than all existing offline optimization strategies, with a sum-rate improvement of up to 26%. Also, Sharifi et al. [195] proposed a deep SARSA-based RL algorithm for user scheduling in a wireless network with a HAPS backup. Modeled as an MDP, it operated with outdated CSI to maximize sum-rate while minimizing active antennas at the HAPS for energy savings. Results showed near-optimal sum-rate performance, and a heuristic method was also introduced for solving the scheduling problem. Jo et al. [111] introduced a multi-agent deep Q-learning algorithm for HAPS power control to reduce outage probability while satisfying interference constraints with TNs. A double deep Q-learning variant was also introduced to avoid action-value overestimation. Simulations showed that agents collaboratively learned a near-optimal power control policy through effective state, reward, and training design. Later, a dynamic coverage path planning algorithm was proposed by Liu et al. [141] for multi-airship formations to overcome the limitations of static approaches for stratospheric airship formations in dynamic coverage scenarios. Using RL, the method stored experiences in a pool for rapid updates via semi-centralized exploration and centralized playback, with task-specific rewards improving network coverage. Meanwhile, Arani, Hu, and Zhu [28] considered a heterogeneous network with UAVs and HAPS as aerial BSs, aiming at optimizing fairness and QoS for GUEs. The proposed approach combines DRL with fixed-point iteration to determine the optimal UAVs positions, 3D trajectories, and channel allocation strategies based on real-time traffic load. Minani et al. [157] introduced an RA method for uplink HAPS-LEO communication using multi-carrier NOMA. The problem was formulated as a mixed-integer nonlinear program under power and channel constraints, and solved with a multi-agent double deep Q-network employing prioritized experience replay and priority sampling to handle dynamic resources and short satellite visibility. Later, Le et al. [130] studied HAPS as flying BSs serving GUEs with NOMA in a downlink THz system. The method optimized power allocation, bandwidth, antenna beamwidth, and HAPS altitude to improve transmission rates while meeting data requirements. A DRL-based DDPG algorithm was applied to manage network dynamics and complexity, maximizing long-term rewards with continuous decision variables.
Recent work shows that federated learning (FL) can strengthen HAPS-enabled networks by enhancing privacy and lowering communication overhead. Approaches like support vector machine (SVM)-based FL for task offloading and hierarchical FL with HAPS as distributed servers improve model accuracy, energy efficiency, and reduce training time for large-scale applications. For example, Wang et al.[223] addressed the minimization of energy and time consumption in mobile-edge computing-enabled HAPS networks. It formulates an optimization problem for user association, service sequencing, and task allocation, and proposes a SVM-FL algorithm to predict user associations without sharing private data. Simulations demonstrate up to 15.4% improvement in energy and time efficiency compared to terrestrial centralized methods. Addressing the limitations of traditional ML in satellite communications, Elmahallawy and Luo [64] introduced a FL framework that employs HAPS as distributed parameter servers for LEO satellite constellations. The proposed framework integrates a hierarchical communication architecture along with model dissemination and aggregation algorithms, significantly reducing training time from days to hours while improving model accuracy. To enable global-scale FL in NTNs, a distributed hierarchical FL framework was proposed by Farajzadeh, Yadav and Yanikomeroglu [66], with HAPS as intermediate servers. Integrating LEO satellites and ground clients while using GEO and MEO satellites for model relaying, the framework enhances privacy, improves model accuracy, manages latency, and boosts scalability. Numerical results confirm its efficiency in training loss reduction and network optimization. Güven and Karabulut Kurt [85] proposed a CNN-aided receiver for HAPSLEO links that jointly addresses channel estimation and carrier frequency offset (CFO) estimation under high Doppler and rapid channel variations. Supervised CNN models are trained to learn the nonlinear mapping from received pilot observations to channel and CFO estimates, reducing estimation errors compared to conventional pilot-based methods. Simulation results demonstrated improved synchronization accuracy and detection performance in representative HAPS–LEO scenarios. Mahyastuty et al. [147] investigated CNN-assisted and CNN-based Successive Interference Cancellation (SIC) replacement for PD-NOMA in HAPS-enabled mMTC and WSN scenarios, where the HAPS acts as an aggregation node. The proposed approach targets low-complexity massive access by leveraging CNNs for multiuser detection, and simulation results showed improved BER performance compared with conventional PD-NOMA receivers under the considered system configurations. Later, Mahyastuty, Pamukti, and Oktaviana [148] introduced a CNN-based decoding framework for power-domain NOMA in HAPS systems to overcome the limitations of successive interference cancellation (SIC), particularly error propagation and sensitivity to user dynamics. The study analyzed the impact of user location and mobility on decoding performance, showing that users closer to the HAPS achieve target BER at lower SNR, while increasing user density raises the required SNR for neighboring users.
Table VI: Mapping between learning paradigms, HAPS-specific constraints, representative applications, and open challenges in HAPS-enabled NTN.
Learning Paradigm HAPS-Specific Constraints Addressed Representative Applications in HAPS/NTN Pending Challenges and Open Issues
DRL / RL [163, 28, 165, 195, 141, 111, 157, 130] Mobility and non-stationarity (platform/user dynamics, intermittent visibility). Real-time decision latency (online control and scheduling). Energy budget for onboard inference/training. Platform positioning and 3D trajectory control (HAPS/UAV). Dynamic RA, power control, and interference mitigation. User scheduling under outdated/partial CSI. Joint beamforming and RIS configuration. Sim-to-real gap; robustness to channel/traffic model mismatch. Sample inefficiency and training instability in non-stationary NTN. Multi-agent scalability and coordination overhead. Safety/constraint satisfaction (hard QoS, interference, energy).
FL [223, 66, 64] Privacy constraints (no raw-data sharing). Limited backhaul / feeder link capacity. Intermittent connectivity and long RTTs in NTN. User association and task offloading prediction at the edge. HAPS-assisted distributed training (HAPS as aggregator/parameter server). Global-scale hierarchical learning across ground–HAPS–satellite tiers. Communication overhead of model updates; compression/quantization needs. Stragglers and synchronization under intermittent links. Statistical heterogeneity (non-IID data) and personalization requirements. Secure aggregation and adversarial/poisoning robustness.
CNN [85, 147, 148] High Doppler, residual CFO, and fast channel variations in HAPS/LEO links. Strong multiuser interference and imperfect SIC in NOMA-based HAPS systems. Non-stationary propagation, traffic dynamics, and domain shift across scenarios. CNN-based channel and CFO estimation for HAPS-assisted and HAPS–LEO links. CNN-based multiuser detection/decoding for PD-NOMA in HAPS and mMTC scenarios. Improved BER, throughput, and robustness compared to conventional receivers. Dependence on representative training datasets; sensitivity to domain mismatch. Increased computational and energy cost at HAPS/UE receivers. Limited interoperability and deployment trade-offs (accuracy vs. latency).
VI-K Cybersecurity and Adversarial Threats in HAPS-Enabled NTN
Cybersecurity threats can put NTN systems at significant risk. Due to exposed wireless links and strong reliance on Positioning, Navigation, and Timing (PNT), platforms such as HAPS are particularly vulnerable, making robust security mechanisms essential. By mapping threats to HAPS subsystems (such as communication payload, power management, telemetry, tracking, and flight control, as well as ground station interfaces), Hjaiji, Ouni and Alouini [95] provided a HAPS-focused cybersecurity analysis. The authors specifically highlighted jamming, replay, and system intrusion attacks, data manipulation and spoofing, DoS/DDoS, adversarial ML attacks, and supply-chain risks; and extensively reviewed the resent advancements in defense mechanisms applicable to HAPS subsystems. A simulation-based distributed-denial-of-service case study that quantifies service/control-plane availability deterioration under assault was also provided as a significant technical component of the study.
Yu et al. [239] examined the communication between several transmitters and a multi-antenna UAV when jamming is present. They suggested cooperative beamforming techniques that reduce deliberate jamming and inter-user interference. The analysis showed how anti-jamming performance is affected by geometry, UAV motion, and beam design. In addition to HAPS-assisted NTN systems, it offers practical techniques for aerial-layer resiliency. Using the minimum variance distortionless response (MVDR) beamforming approach, Xi, Liu and Zhao [230] investigated anti-jamming for satellite ground-station arrays. After creating a signal model that takes direction-dependent array/reflector effects into consideration, proposed approach calculates MVDR weights to reduce interference. Simulations verified the improvements in output SINR in relation to jammer angle, power, and snapshot availability. Aiming array-level interference reduction in SatCom/NTN feeder lines, the study may be potentially generalized to HAPS-related application. Otay, Humadi, and Karabulut Kurt [169] investigates a security-relevant scenario in which a HAPS prevents LEO-to-ground communication by acting as a downlink jammer. A direct satellite-ground link with a jamming HAPS and a cooperative two-satellite path (direct+relay) under the same HAPS jammer are the two scenarios that are examined. The authors derive analytical formulas based on outages that quantify the degradation of link dependability caused by jammer power and shape. One important lesson is that satellite cooperation and relaying can lower the likelihood of a serious disruption because, in the worst scenario, both pathways must be jammed at the same time. Cao et al. [42] adapts spectrum-access choices for LEO satellite downlinks under jamming by using multi-agent deep reinforcement learning. Agents choose channels and actions that minimize the impact of jamming while maintaining throughput by learning policies from observation histories. Performance was compared to less adaptable techniques using a jammer model with channel-switching dynamics. For NTN resource management, including HAPS-based architectures, it offers a practical illustration of autonomy-driven anti-jamming. Gu, Xing, and You [81] proposed a global navigation satellite systems spoofing detection approach based on coupled visual/inertial/GNSS positioning algorithm. To lower false alarms, it incorporates confidence handling and residual/consistency testing in the fused navigation solution. Datasets and experimental settings that mimic spoofing effects are used to validate the method. It is pertinent to aerial platforms where secure control and safe flying depend on strong PNT. A lightweight 1D-CNN detector targeted at small UAVs with an emphasis on early detection of global positioning system (GPS) spoofing was proposed by Sung et al. [205]. It reports practical runtime/energy feasibility on embedded hardware and assesses detection performance against classical ML baselines. The technique supports defensive actions (such as safe landing logic) and was shown under practical UAV operational restrictions. In layered NTN ecosystems, it is a powerful applicable reference for UAVs security.
The issue of huge, nearly simultaneous handovers in NTN (e.g., LEO footprints moving quickly), which might overwhelm control signaling and create security gaps, was addressed by Zhang et al. [242]. It suggests a safe group-handover protocol that combines steps to lower signaling costs while maintaining authorization and authentication objectives. The study compares cost and delay to baseline methods and assesses performance using discrete-event simulation. For LEO-to-device links, Coppola et al. [51] develop a physical-layer ranging technique that can withstand spoofing and relay-style manipulation. It examines attack surfaces and presents a safe range waveform/processing chain that works with realistic OFDM-style systems. The authors assess ranging accuracy and resilience under representative satellite channel effects and offer security justifications. The end product is a tangible foundation for NTN systems’ secure timing/ranging (PNT-adjacent) services. Wu et al. [228] examine covert communication in an integrated HAPSTN, approaching security from a low likelihood of detection perspective. To conceal the signal from a warden or eavesdropper, it employs a constant-power auxiliary node that sends artificial noise. The authors present closed-form formulas for outage probability and effective covert rate after deriving the covert constraint connecting transmitter and auxiliary node powers. These findings apply to HAPS deployments where a design objective under surveillance is to conceal transmission behavior (not just encryption).
VII Open Research Challenges
A key direction for future 6G systems is the integration of TNs and NTNs through HAPS-enabled connectivity. Although several related challenges have been partially studied in terrestrial, satellite, and standalone HAPS networks, many issues remain unresolved—or become more critical—under large-scale, multi-tier TN/NTN integration. In particular, heterogeneous propagation conditions, wide coverage areas, platform mobility, and stringent latency and reliability constraints introduce new cross-layer coupling effects spanning the physical, MAC, and network layers. The following summarizes the key open research challenges (see Table VII), clearly distinguishing HAPS-specific open issues from those already identified in prior work.
Table VII: Summary of Open Research Challenges in HAPS-Enabled TN/NTN Integration
Research Challenge (Time Horizon, Type) Key Open Issues
Channel Modeling (Short-term, Technical) Especially at mmWave and higher frequencies, joint consistent modeling spanning terrestrial–HAPS–satellite links under non-stationarity, large coverage, and weather-/terrain-dependent effects.
Spectrum Management (Short- to Long-term, Technical/Regulatory) Interference cooperation and scalable spectrum sharing across diverse tiers with dense deployments and dynamic beams.
Mobility Management and Handover (Short-term, Technical) Robust signaling and precise handover time in the face of diverse propagation delays, quasi-stationary platforms, and large coverage areas.
Multi-Connectivity (Short-term, Technical) Bearer splitting/duplication across heterogeneous networks under dynamic mobility: control-plane overhead, synchronization, and scalability.
Interoperability and Standardization (Long-term, Technical/Standardization) Protocol development beyond existing norms to facilitate real-time multi-tier coordination, huge connection, and large delays.
Energy Efficiency (Short- to Long-term, Technical) Energy harvesting, storage, and energy-conscious scheduling are all components of sustainable HAPS operation under stringent energy limitations.
Legal and Regulatory Issues (Long-term, Regulatory) Large-scale HAPS deployment requires international cooperation, safety, spectrum control, and airspace management.
Deployment Costs (Long-term, Economic) For widespread adoption, public-private collaborations, sustainable business models, and economic viability are required.
Security, Jamming, and Adversarial ML Threats (Short- to Long-term, Technical) Protection in wide-area LoS situations from hostile ML attacks, spoofing, cyberthreats, and jamming.
Channel Modeling: As previously examined in the literature (e.g., [251, 97, 237, 170, 241]), Doppler effects, delay spread, and atmospheric attenuation are among the important propagation characteristics of HAPS-enabled NTNs that have been examined under a variety of assumptions. However, the requirement for jointly consistent representations across diverse links and large geographic areas keeps channel modeling open in HAPS-enabled TN/NTN integration. Channel coherence is affected by non-stationarity brought on by platform motion, altitude changes, and atmospheric dynamics, while spatial heterogeneity is introduced by effects that rely on the topography and the weather. At mmWave and higher frequencies, when attenuation and beam misalignment can significantly impair link reliability, these difficulties are made worse.
Spectrum: As reported in earlier studies on spectrum sharing and interference management (e.g., [80, 140, 227, 134]), spectrum coexistence between HAPS, satellite, and terrestrial systems is a well-known challenge. This problem is exacerbated by diverse coverage scales, dynamic beamforming, and dense node distributions in HAPS-enabled multi-tier deployments. Scalable coordination methods, scheduling that considers interference, and adaptive spectrum access techniques that function well in the face of mobility and quickly shifting interference situations are among the unresolved difficulties.
Multi-connectivity: It has been demonstrated that multi-connectivity can support strict URLLC requirements, including in circumstances related to NTN [187, 186]. But even with the Stage-2 architecture outlined in 3GPP TS 37.340 [7], there is still room for multi-connectivity advantages to be practically realized in HAPS-enabled NTN deployments. Syncing across heterogeneous links, scalability under dynamic mobility situations, and higher signaling overhead are among the main obstacles [183].
Mobility Management: Numerous studies have been conducted on mobility and handover in HAPS-assisted NTNs [112, 113, 131]. However, mobility management is still a challenge in HAPS-enabled TN/NTN integration because of diverse propagation delays, quasi-stationary aerial platforms, and vast coverage regions. The terrestrial assumptions that underlie current LTE/NR mobility procedures [6, 9] may result in inefficient handover timing and signaling, especially for services that are sensitive to latency and reliability under high mobility and limited signal strength gradients [240].
Interoperability: Interoperability is still a major obstacle, as stated in current TN/NTN standardization initiatives (e.g., [2, 4]). Although the present standards offer a starting point, more protocol development is needed to accommodate huge connectivity, substantial propagation delays, and real-time coordination between satellites, HAPS, and terrestrial nodes.
Energy: Energy efficiency is still a significant challenge, as was previously mentioned in the context of energy-efficient HAPS operation (e.g., [177, 41]). In order to guarantee continued operation under changing lighting and traffic situations, there are still unresolved challenges with effective energy harvesting, long-term storage, dynamic power allocation, and energy-aware scheduling.
Legal Regulations: Large-scale HAPS deployment presents unresolved issues with airspace management, safety, spectrum regulation, and international cooperation, despite the fact that regulatory features of aerial and non-terrestrial platforms have been taken into account in previous work.
Deployment Costs: The economic viability of large-scale HAPS deployment is still up for debate, despite the fact that techno-economic elements have been examined in previous research. To justify long-term investment, viable revenue models and public-private partnerships are necessary due to high deployment and maintenance costs.
Security, Jamming, and Adversarial Machine Learning Threats: As was previously mentioned in the context of HAPS-enabled NTN security [51, 242, 179, 230, 234, 239, 81, 205, 42, 169, 95, 228], HAPS-enabled systems are susceptible to cyberattacks, spoofing, and jamming due to their exposed wireless configuration. Additionally, the growing dependence on machine learning for optimization and control raises the possibility of adversarial attacks, underscoring the necessity of strong and reliable security measures.
VIII Discussion and conclusion
The evolution toward 6G requires tight TN–NTN integration. Operating in the stratosphere, HAPS can bridge terrestrial cells, LEO satellites, and UAV networks to deliver wide-area, reliable, low-latency connectivity. This survey synthesized HAPS use cases, deployment architectures (transparent vs. regenerative), spectrum options from sub-7 GHz to FSO, and key enablers (channel modeling, UE association, resource allocation, multi-link connectivity, mobility management, antenna technology, channel estimation, beam and interference control, AI/ML, and energy efficiency), and outlined open challenges. Unlike prior surveys, we positioned HAPS as intelligent network nodes—not mere relays—within standardized NTN frameworks and multi-band designs. Looking ahead, priorities include: (i) efficient multi-band strategies (FR1/FR3/FR2 + optical) with robust coexistence; (ii) cross-layer TN–NTN interoperability aligned with recent 3GPP NTN releases; (iii) AI-based control with reliability/energy guarantees; (iv) sustainability-driven platform and gateway design; and (v) progress on regulation and cost models to enable scale. With continued innovation, HAPS can become a practical cornerstone of 6G and a lever to narrow the digital divide.